### v1

# Deep Reinforcement Learning for Bipedal Locomotion: A Brief Survey

Abstract
Bipedal robots are garnering increasing global attention due to their potential applications and advancements in artificial intelligence, particularly in Deep Reinforcement Learning (DRL). While DRL has driven significant progress in bipedal locomotion, developing a comprehensive and unified framework capable of adeptly performing a wide range of tasks remains a challenge. This survey systematically categorizes, compares, and summarizes existing DRL frameworks for bipedal locomotion, organizing them into end-to-end and hierarchical control schemes. End-to-end frameworks are assessed based on their learning approaches, whereas hierarchical frameworks are dissected into layers that utilize either learning-based methods or traditional model-based approaches. This survey provides a detailed analysis of the composition, capabilities, strengths, and limitations of each framework type. Furthermore, we identify critical research gaps and propose future directions aimed at achieving a more integrated and efficient framework for bipedal locomotion, with potential broad applications in everyday life.
Index Terms:
Deep Reinforcement Learning, Humanoid Robots, Bipedal Locomotion, Legged Robots
I Introduction
Humans navigate complex and varied environments, performing diverse locomotion tasks with only two legs. To facilitate dynamic bipedal locomotion, model-based methods were introduced in the 1980s and have since evolved significantly [1, 2, 3]. These methods, characterized by rapid convergence, furnish a predictive framework for understanding environmental structures. However, they struggle to adapt in dynamically challenging environments that are difficult to model precisely. More recently, advancements in machine learning have provided novel approaches. Reinforcement learning (RL)-based methods, in particular, are adept at navigating the full dynamics of robot-environment interactions [4]. Additionally, hybrid approaches that combine model-based and learning-based methods have been developed to leverage the advantages of both. Yet, the question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
Figure 1: Common bipedal and humanoid robots used as platforms for testing DRL frameworks. (a) NAO, a toy-like 3D humanoid robot, actuated by servo motors [5]. (b) Rabbit, a 2D bipedal robot, actuated by torque control [6]. (c) Cassie, a 3D bipedal robot, also actuated by torque control [7]. (d) ATLAS, a 3D humanoid robot, driven by hydraulics [8]. (e) Digit, a full human-sized 3D humanoid robot, an upgrade based on Cassie and actuated by torque control [9].
To address this, we explore recent advancements in deep reinforcement learning (DRL)-based frameworks, categorizing control schemes into two primary types: (i) end-to-end and (ii) hierarchical. End-to-end frameworks map robot states directly to control outputs at the joint level, while hierarchical frameworks adopt a structured approach, decomposing decision-making into multiple layers. Here, a High-Level (HL) planner addresses navigation and path planning, while a Low-Level (LL) controller focuses on fundamental locomotion skills. The highest decision-making tier, the task level, receives direct input from task or user commands.
Figure 2: Classification of DRL-based control schemes.
The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications. Although the application of RL to simple 2D bipedal robots began in 2004 [10, 11], it took several years before deep reinforcement learning (DRL) algorithms emerged. These DRL-based methods have since shown promising results in physical simulators [12, 13, 14]. Agility Robotics introduced the first sim-to-real end-to-end learning framework in 2019, which was applied on the 3D torque-controlled bipedal robot Cassie, as shown in Fig. 1(c) [7]. Besides model-based reference learning, the policy can also incorporate motion capture data [15, 16, 17, 18], or start from scratch [19] to explore solutions freely. Recent studies demonstrate that end-to-end frameworks robustly handle complex and diverse tasks [20, 21, 22].
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. A notable framework employs a learned High-Level (HL) planner coupled with a Low-Level (LL) model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [23, 24, 25]. Another innovative construction integrates a learned feedback controller with an HL planner, categorized under the DRL Feedback Control Hybrid Scheme [26, 27]. Additionally, a learned hierarchical control scheme [28] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [12, 13, 29].
While several review papers discuss RL for general robotics [4] and model-based methods for bipeds [1, 2, 3], none specifically focus on DRL-based frameworks for bipeds. This survey aims to address this gap by summarizing current research progress, highlighting the structure and capabilities of bipedal locomotion frameworks, and exploring future directions. We also catalogue DRL-based frameworks, as depicted in Fig. 2. The primary contributions of this survey are:
  • A comprehensive summary and cataloguing of DRL-based frameworks for bipedal locomotion.
  • A detailed comparison of each control scheme, highlighting their strengths, limitations, and characteristics.
  • The identification of current challenges and the provision of insightful future research directions.
The paper is organized as follows: Section II focuses on end-to-end frameworks, categorized by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV addresses existing gaps, ongoing challenges, and potential future research directions. Finally, Section V concludes the paper.
II End-to-end framework
The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly translates sensory inputs—such as images, lidar data, or proprioceptive feedback [30]—along with user commands [19] or pre-defined references [31], into joint-level control actions. These actions encompass motor torques [32], positions, and velocities [15]. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
End-to-end strategies primarily simplify the design of low-level tracking to basic elements, such as a Proportional-Derivative (PD) controller. These methods can be broadly categorized based on their reliance on prior knowledge into two types: reference-based and reference-free. The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
In the following sections, we will delve into various representation frameworks, exploring their characteristics, limitations, and strengths in comprehensive detail. To facilitate an understanding of these distinctions, Table I provides a succinct overview of the frameworks discussed.
TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework. The dashed line in the implementation flow chat refers to optional.
Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
Residual learning [33] [34] [35] Forward walk undirectional walk Omni-walk Add residential term to the known motor positions at the current time step. A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains. Refer to caption
Guided learning [36] [31] [37] [22] Forward walk Versatile walk Versatile jump Versatile motions Mimic the predefined reference and directly specifies joint-level command. A: Accelerate the learning process and robust to the terrains. D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment. Refer to caption
Reference-free learning [19] [38] [20] Periodic motions Stepping stones Vision-based Learn locomotion skills from scratch without any prior knowledge. A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources. Refer to caption
00footnotetext: Forward Walk involves the bipeds walking straight ahead. Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities. Omni-Walk grants the bipeds the ability to walk in any direction. Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities. Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping. Versatile Jump refers to jump towards different desired targets. Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
II-A Reference-based learning
Reference-based learning utilizes prior knowledge, allowing the policy to develop locomotion skills by adhering to predefined references, which may be derived from trajectory optimization (TO) techniques or captured through motion capture systems. This method facilitates the acquisition of locomotion skills compared to alternative approaches, though it typically results in locomotion patterns that closely resemble the predefined references or motion clips, thus limiting the variety of gait patterns. Generally, this approach can be divided into two primary methods: (i) residual learning and (ii) guided learning.
II-A1 Residual learning
This method involves a framework that is aware of the current reference joint positions and applies offsets determined by the policy to modify motor commands at the current timestep. By utilizing predefined motion trajectories, the residual term acts as feedback control, compensating for errors and enabling the biped to achieve dynamic locomotion skills.
Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [33]. This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimization (PPO) algorithms, as detailed in Appendix A. The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep. These modified references are then processed by a Proportional-Derivative (PD) controller to set the desired joint positions. While this framework enhanced the robot’s ability to perform tasks beyond standing [39], its physical deployment on a bipedal robot has not yet occurred, potentially rendering it impractical for managing walking at varying speeds and limiting movement to a single direction.
To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied on a physical bipedal robot [34]. This process and its key points are further explored in Appendix B. Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks. However, the robot’s movements remain constrained to merely walking forward or backward. A novel approach in residual learning was introduced to enable unidirectional walking, where the policy outputs a residual term added to the current positional states, facilitating gradual omnidirectional walking [35].
II-A2 Guided learning
Guided learning trains policies to directly output the desired joint-level commands, eschewing the addition of a residual term. The reward structure in this approach is focused on closely imitating predefined references.
A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [36]. In this framework, the action space directly maps to the joint angles, and desired joint positions are managed by joint PD controllers. The framework also incorporates a Long Short-Term Memory (LSTM) network, as detailed in Appendix A, which is synchronised with periodic time inputs. However, this model is limited to a single locomotion goal: forward walking. A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [31], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
Despite these advancements, the parameterization of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances. To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [37]. This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping maneuvers. An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviors [18]. This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [22] based on the foundational work in [37]. This framework enhances the learning process by incorporating a wide range of locomotion skills and introducing a dual I/O history approach, marking a significant breakthrough in creating a robust, versatile, and dynamic end-to-end framework. However, experimental results indicate that the precision of locomotion features, such as velocity tracking, remains suboptimal.
Guided learning methods expedite the learning process by leveraging expert knowledge and demonstrating the capacity to achieve versatile and robust locomotion skills. Through the comprehensive evaluation [22], it is demonstrated that guided learning employs references without complete dependence on them. Conversely, residual learning exhibits failures or severe deviations when predicated on references of inferior quality. This shortfall stems from the framework’s dependency on adhering closely to the provided references, which narrows its learning capabilities.
Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities. Additionally, this approach exhibits limited adaptability in responding effectively to unforeseen environmental changes or novel challenges.
II-B Reference-free learning
In reference-free learning, the policy is trained using a carefully crafted reward function rather than relying on predefined trajectories. This approach allows the policy to explore a wider range of gait patterns and adapt to unforeseen terrains, thereby enhancing innovation and flexibility within the learning process.
The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models. A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [14]. This framework introduced a novel term into the loss function and utilized a curriculum learning strategy to effectively shape gait patterns. Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model, Cassie, though this has yet to be validated outside of simulation [40].
Considering the practical implementation of theoretical models, significant efforts have been directed towards developing sim-to-real frameworks in robotics studies. A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [19]. This framework employs periodic rewards to facilitate initial training within simulations before successfully transitioning to a physical robot. It has been further refined to adapt to diverse terrains and scenarios. For instance, robust blind walking on stairs was demonstrated through terrain randomization techniques in [38]. Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [41], thus enabling the robot to effectively navigate stepping stones [20]. Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalizes terrain information [42].
This approach to learning enables the exploration of novel solutions and strategies that might not be achievable through mere imitation of existing behaviours. However, the absence of reference guidance can render the learning process costly, time-consuming, and potentially infeasible for certain tasks. Moreover, the success of this method hinges critically on the design of the reward function, which presents significant challenges in specifying tasks such as jumping.
III Hierarchy framework
Unlike end-to-end policies that directly map sensor inputs to motor outputs, hierarchical control schemes deconstruct locomotion challenges into discrete, manageable layers or stages of decision-making. Each layer within this structure is tasked with specific objectives, ranging from high-level navigation to fundamental locomotion skills. This division not only enhances the framework’s flexibility but also simplifies the problem-solving process for each policy.
The architecture of a hierarchical framework typically comprises two principal modules: an HL planner and an LL controller. This modular approach allows for the substitution of each component with either a model-based method or a learning-based policy, further enhancing adaptability and customisation to specific needs.
Hierarchical frameworks can be classified into three distinct types based on the integration and function of their components:
  1. 1.
    Deep planning hybrid scheme: This approach combines strategic, high-level planning with dynamic low-level execution, leveraging the strengths of both learning-based and traditional model-based methods.
  2. 2.
    Feedback DRL control hybrid scheme: It focuses on integrating direct feedback control mechanisms with deep reinforcement learning, allowing for real-time adjustments and enhanced responsiveness.
  3. 3.
    Learned hierarchy scheme: Entirely learning-driven, this scheme develops a layered decision-making hierarchy where each level is trained to optimise specific aspects of locomotion.
These frameworks are illustrated in Fig. 3. Each type offers unique capabilities and exhibits distinct characteristics, albeit with limitations primarily due to the complexities involved in integrating diverse modules and their interactions.
For a concise overview, Table 3 summarises the various frameworks, detailing their respective strengths, limitations, and primary characteristics. The subsequent sections will delve deeper into each of these frameworks, providing a thorough analysis of their operational mechanics and their application in real-world scenarios.
TABLE II: Summary and comparison of Hierarchy framework
Control Scheme Works Module characteristic Advantages and Disadvantages
Deep Planning Hybrid Scheme [24] [43] [44] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks. A: Enhanced command tracking capabilities, generalized across different platforms, sampling efficiency, and robust. D: Complicate system and communication between layers, require precise model, lack generalization regarding different tasks.
Feedback DRL Control Hybrid Scheme [45] [26] [27] Gait library + feedback policy Footstep planner + feedback policy Model-based planner + feedback policy LL feedback policy receives non-learned HL planner as input to achieve locomotion skills. A: Short inference times, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, reducing sampling efficiency.
Learned Hierarchy Framework [12] [13] [29] HL policy + LL policy HL policy + LL policy HL policy + LL policy Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills. A: provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy. D: inefficienty sim-to-real, complicated interface between layers, training expensively.
III-A Deep planning hybrid scheme
In this scheme, robots are pre-equipped with the ability to execute basic locomotion skills such as walking, typically managed through model-based feedback controllers or interpretable methods. The addition of an HL learned layer focuses on strategic goals or the task space, enhancing locomotion capabilities and equipping the robot with advanced navigation abilities to effectively explore its environment.
Several studies have demonstrated the integration of an HL planner policy with a model-based controller to achieve tasks in world space. A notable framework optimises task space level performance, eschewing direct joint level and balancing considerations [24]. This system combines a residual learning planner with an inverse dynamics controller, enabling precise control over task-space commands to joint-level actions, thereby improving velocity tracking, foot touchdown location, and height control. Further advancements include a hybrid framework that merges HZD-based residual deep planning with model-based regulators to correct errors in learned trajectories, showcasing robustness, training efficiency, and effective velocity tracking [25]. These frameworks have been successfully transferred from simulation to reality and validated on robots such as Cassie.
However, the limitations imposed by residual learning constrained the agents’ capacity to explore a broader array of possibilities. Building on previous work [25], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [43]. In this approach, a purely learning-based HL planner interacts with an LL controller using an Inverse Dynamics with Quadratic Programming formulation (ID-QP). This policy adeptly captures dynamic walking gaits through the use of reduced-order states and simplifies the learning trajectory. Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalized across various bipedal platforms, including Digit, Cassie, and RABBIT.
In parallel, several research teams have focused on developing navigation planners specifically for toy-like humanoid robots, which provide greater physical stability compared to torque-driven or hydraulic bipedal robots as shown in Fig. 1. One notable study [46] implemented a visual navigation policy on the NAO robot, depicted in Fig. 1(a), utilizing RGB cameras as the primary sensory modality. This system has demonstrated successful zero-shot transfer to real-world scenarios, enabling the robot to adeptly navigate around obstacles. Further research [44] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a center of mass (CoM) trajectory. This configuration is coupled with a whole-body controller, facilitating dynamic activities like soccer shooting. Despite their platform’s stability, provided by large feet and a lightweight structure, these robots exhibit limited dynamic movement capabilities compared to full-sized humanoid robots. Consequently, this research primarily addresses navigation and task execution.
Regarding generalization, these frameworks have shown potential for adaptation across different types of bipedal and humanoid robots with minimal adjustments, demonstrating advanced user command tracking [43] and sophisticated navigation capabilities [44]. However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping. Furthermore, while these systems adeptly navigate complex terrains with obstacles, footstep planning alone is insufficient without concurrent enhancements to the robot’s overall locomotion capabilities. Moreover, the requisite communication between the two distinct layers of the hierarchical framework may introduce system complexities. Enhancing both navigation and dynamic locomotion capabilities within the HL planner remains a significant challenge.
Figure 3: Hierarchy Control Scheme Diagram: (a) A basic hierarchical scheme with two layers, where each module can be substituted with a learned policy. (b) A deep planning hybrid scheme, where the High-Level (HL) planner is learned. (c) A learning-based feedback control hybrid scheme, with a learned Low-Level (LL) controller. (d) A comprehensive DRL hierarchy control scheme, where both layers are learned.
III-B Feedback DRL control hybrid scheme
In contrast to the comprehensive approach of end-to-end policies discussed in Section II, which excels in handling versatile locomotion skills and complex terrains with minimal interface times, the Feedback DRL Control Hybrid Scheme integrates DRL policies as LL controllers. These LL controllers, replacing traditional model-based feedback mechanisms, work in conjunction with HL planners that process terrain information, plan future walking paths, and maintain robust locomotion stability.
For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [45]. Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness. A more dynamic approach involves online planning, which has shown greater adaptability and efficiency. One notable framework combines a conventional foot planner with an LL DRL policy [26], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands. Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies. However, this work has not yet been transferred from simulation to real-world applications. Later, a similar structure featuring an HL foot planner and an LL DRL policy was proposed [27]. This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
A recent development has shown that focusing solely on foot placement might restrict the stability and adaptability of locomotion, particularly in complex maneuvers. A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [47]. This system employs a residual learning architecture, where the DRL policy’s outputs are merged with the planner’s directives before being relayed to the PD controller. This integrated approach not only concerns itself with foot placement but also generates comprehensive trajectories for trunk position, orientation, and ankle yaw angle, enabling the robot to perform a wide array of locomotion skills including walking, squatting, turning, and stair climbing.
Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities. These policies have demonstrated high efficiency in locomotion and accurate reference tracking. Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited. Nonetheless, these learning algorithms often face challenges in environments characterized by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [48].
Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [27]. This capability effectively addresses the navigational limitations observed in end-to-end frameworks. Moreover, unlike the deep planning hybrid scheme where modifications post-policy establishment can be cumbersome, this hybrid scheme offers enhanced flexibility for on-the-fly adjustments.
Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models. Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimization and hinder rapid deployment and iterative enhancements [48].
III-C Learned hierarchy framework
The Learned Hierarchy Framework merges a learned HL planner with an LL controller, focusing initially on refining LL policies to ensure balance and basic locomotion capabilities. Subsequently, an HL policy is developed to direct the robot towards specific targets, encapsulating a structured approach to robotic autonomy.
The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [12]. In this setup, LL policies, informed by human motions or trajectories generated via Trajectory Optimization (TO), strive to track these trajectories as dictated by the HL planner while maintaining balance. An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths. This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal. The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [13].
However, despite its structured and layered approach, which allows for the reuse of learned behaviors to achieve long-term objectives, these frameworks have predominantly been validated only in simulations. The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviors, including stability issues like falling.
Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [29].
Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics. Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [12]. As discussed in the previous section, the benefits of integrating RL-based planners with RL-based controllers have been effectively demonstrated. This combination enables the framework to adeptly manage a diverse array of environments and tasks.
Within such a framework, the High-Level (HL) policy is optimized for strategic planning and achieving specific goals. This optimization allows for targeted enhancements depending on the tasks at hand. Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives.
Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored. The transition from simulation to real-world scenarios is fraught with challenges, not least because of the complexities involved in training and integrating two separate layers within the control hierarchy. Ensuring effective communication and cooperation between these layers is critical, requiring a meticulously defined communication interface to avoid operational discrepancies.
Additionally, the training process for each policy within the hierarchy demands considerable computational resources. The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalize to real-world conditions. This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
IV Challenges and future research directions
While learning-based frameworks for bipedal robots have demonstrated considerable potential, they have also clearly exposed the limitations inherent to each framework. Moreover, several critical areas remain largely unexplored, especially within the realm of legged robotics, where the pace of research on bipedal robots lags behind that of their quadruped counterparts. This discrepancy in research progress can be attributed to several factors, including the higher costs and less mature technology associated with bipedal robot hardware, as well as the inherent instability issues that bipedal designs face.
To gain a deeper understanding of these challenges and to outline potential future directions, it is instructive to first review existing research on quadruped robots. The insights gained from quadrupeds, which benefit from more robust research outputs and technological advancements, can provide valuable lessons for overcoming similar challenges in bipedal systems.
IV-A Recent progress with quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [49, 50], to systems that include the modeling of deformable terrain to enhance locomotion over compliant surfaces [51]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [52], and sophisticated acrobatic motions are achieved through imitation learning [53].
The domain of quadruped DRL has also seen significant advancements in complex hybrid frameworks that integrate vision-based systems. To date, two primary versions of such frameworks have been developed: one where a deep planning module is paired with model-based control [54], and another that combines model-based planning with low-level DRL control [48, 55]. The latter has shown substantial efficacy; it employs a model predictive control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [56] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway. These include mimicking real animals through motion capture data and imitation learning [57, 58], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [59, 60] or by using the robots’ legs [61]. Notably, the research presented in [60] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks. Establishing a unified framework could bridge this gap, an essential step given the integral role of bipedal robots in developing full humanoid systems. Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
IV-B Gaps and challenges
Despite numerous promising developments in the field of bipedal and humanoid robotics, significant gaps remain between current research outcomes and the ultimate goals. This discussion concentrates on the gaps in frameworks and algorithms rather than hardware, structured around two pivotal questions: 1) Is it possible to design a unified framework that achieves both generalization and precision? 2) Can we develop a straightforward end-to-end policy capable of managing all tasks efficiently?
IV-B1 Generalization versus precision
DRL has demonstrated potential in facilitating versatile locomotion skills [22]; however, challenges such as poor velocity tracking and issues with precise control often arise. While [43] shows that deep planning combined with model-based control can achieve precise velocity tracking, and [37] illustrates successful end-to-end control for precise jumping, the creation of a policy that effectively handles both diverse tasks and precise movements remains elusive. Furthermore, [41] introduces a foot constraint policy framework, enabling precise target tracking and accurate touchdown locations. Yet, there is still no framework that comprehensively addresses the dual demands of versatility and precision in locomotion.
The difficulty in simultaneously achieving precise control and a broad range of actions in bipedal locomotion using DRL stems from several factors:
  • Complex dynamics: Bipedal locomotion involves intricate dynamics, posing a significant challenge to maintaining both dynamic motion and precision.
  • Resource intensity: Executing diverse locomotion tasks requires considerable computational power and extensive data, necessitating high-quality hardware and efficient DRL algorithms.
  • Training conflicts: Training DRL systems to achieve both precision and versatility often leads to conflicts. Designing reward functions and training policies that satisfy both criteria is inherently complex.
These challenges underscore the need for innovative solutions that can bridge the gap between the capabilities of current frameworks and the ambitious goals of advanced bipedal and humanoid robotics.
IV-B2 Simplifying frameworks to overcome complex tasks
The envisioned ideal in robotic design is an end-to-end framework that enables robots to traverse various terrains using versatile locomotion skills. Although current research often focuses on enhancing frameworks by adding complex components to mitigate inherent limitations, such as the integration of a foot planner for omnidirectional locomotion and stair navigation, as demonstrated in [26, 27], simpler end-to-end frameworks have also proven effective. These frameworks adeptly navigate challenging terrains and perform a diverse range of locomotion tasks with fewer components [20, 22].
The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure. To achieve an optimal end-to-end framework, advancements in several key areas are essential:
  • Robust and efficient DRL algorithms: Development of algorithms that can manage high-dimensional and continuous control problems more effectively.
  • Specialized neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [42].
  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviors and strategic outcomes.
  • Advanced computational resources: Enhancement of computational capabilities to support more intensive training and faster inference, facilitating real-time decision-making in dynamic environments.
By focusing on these developmental areas, the potential to create a unified, efficient, and less complex framework for handling complex locomotion challenges in bipedal robots is significantly increased.
IV-C Future directions
The exploration of quadruped robotics has yielded substantial advancements, yet the full potential of bipedal robotics remains largely untapped. Building on the successes and innovative approaches observed in quadruped robots, several key future directions emerge that could significantly enhance bipedal and humanoid robotics.
IV-C1 Unified framework
Currently, no single framework exists that enables bipedal or humanoid robots to adeptly navigate all types of terrains, including stepping stones, stairs, deformable terrain, and slippery surfaces. A promising approach, as evidenced by recent work in quadruped robots [48], utilizes MPC to generate reference motions, which a low-level DRL policy then tracks. This method, coupled with the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module, simplifies the terrain representation into a height map, facilitating more effective navigation. This success encourages further exploration into hybrid frameworks that combine model-based methods with DRL, inheriting the strengths of both approaches, as discussed in Section III. However, hybrid frameworks present challenges such as training efficiency and system complexity, which demand considerable computational resources and extensive training periods.
Moreover, recent studies [20, 42, 22] have demonstrated the potential of end-to-end frameworks enhanced with vision-based information. These frameworks successfully navigate challenging terrains and execute dynamic motions, suggesting the feasibility of a unified framework capable of handling diverse environments and tasks. Training strategies such as curriculum learning and task randomization could be employed, utilizing visual height maps as inputs to the policy, enhancing the robot’s ability to adapt and perform in varied scenarios.
In addition, the introduction of a DRL end-to-end framework incorporating transformer models, as in [62], presents significant possibilities for integrating locomotion skills with language and vision capabilities. The use of large-scale models capable of processing and condensing extensive data sets into a coherent model could expand the robot’s range of capabilities, maintaining versatility across a broad spectrum of tasks.
The exploration of transformers and other large-scale models holds considerable promise for enhancing generalizability and adaptability in complex tasks, warranting further investigation into their potential applications in bipedal robotics.
IV-C2 Vision-based learning framework
Vision plays a critical role in enabling robots to navigate challenging terrains, such as blind drops, where tactile and other sensory inputs may not provide sufficient information. Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [38, 43]. Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal and humanoid robots [20, 42, 27].
Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
  • Height scanner mapping: This approach, evaluated in works like [42], involves using height maps generated by scanners to inform locomotion strategies. These maps provide detailed topographical data, allowing robots to plan steps on uneven or obstructed surfaces more effectively.
  • Direct vision inputs: Directly utilizing inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [46, 63]. Although previous studies like [46] have integrated visual navigation by feeding visual information to a High-Level (HL) planner, the potential of direct visual inputs to RL policies has not been fully explored.
Enhancing the capability of bipedal robots to directly interpret and utilize visual data without intermediary processing can revolutionize their adaptability and efficiency in real-world scenarios. The exploration of direct vision inputs to reinforcement learning policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
IV-C3 Bridge the gap from simulation to reality
While simulations offer a safe and cost-effective environment for developing robotics policies, the transition from simulation to real-world application often encounters significant challenges due to the approximations and simplifications made in simulations. Numerous sim-to-real frameworks [34, 64, 65] have shown high efficiency and performance, as detailed in Appendix B. Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments. Moreover, many studies [26, 66, 21] remain validated only in simulation settings.
IV-C4 Loco-manipulation tasks
Loco-manipulation, which combines locomotion and manipulation, presents opportunities for humanoid robots to excel beyond purely bipedal capabilities. Few studies have addressed this integrated task; one such study [67] demonstrated a ’box transportation’ framework. This framework decomposes the task into five distinct policies, each addressing different aspects of the transportation process. However, this approach lacks efficiency and does not incorporate vision-based information, suggesting substantial room for improvement. Moreover, the challenges of managing mobile tools like scooters [68] or dynamically interacting with objects such as balls [69] introduce further complexities.
Decomposing loco-manipulation tasks into multiple layers could simplify the challenges, allowing for more precise and flexible control by manually tuning individual components of the task [43]. This structured control approach provides a more coordinated response to complex interactions within the robot’s environment, facilitating the execution of task-specific commands.
Alternatively, an end-to-end framework may enable bipeds to perform a variety of tasks through task randomization and structured curriculum learning methods, progressively teaching the policy [35, 27, 22]. During training, such policies can also learn human-like movements from motion capture data [70, 18, 16], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
IV-C5 Designing reward functions
The development of effective reward functions is a critical challenge in the field of deep reinforcement learning (DRL) for bipedal robots. While periodic reward functions have been designed to facilitate cyclic movements like walking [19], there remains a significant gap in crafting reward functions for non-periodic actions such as jumping. These actions require distinct considerations for success and efficiency, yet current research lacks comprehensive methods for their reward structure. Furthermore, minimizing the need for extensive manual tuning while achieving high performance in DRL systems continues to be a substantial challenge, pointing to the need for more adaptive and automatically adjusting reward mechanisms.
IV-C6 Integrating large language models
The integration of Large Language Models (LLMs) into bipedal robotics opens new avenues for contextual understanding and task execution, significantly enhancing the robots’ interaction capabilities. LLMs, when implemented at the highest task level, offer substantial promise for improving human-robot interaction, making these systems more intuitive and responsive [71]. The potential applications of this technology are broad and impactful, spanning sectors such as industrial automation, where robots can perform complex assembly tasks; healthcare, offering assistance in patient care and rehabilitation; assistive devices, providing support for individuals with disabilities; search and rescue operations, where robust and adaptive decision-making is critical; and entertainment and education, where interactive and engaging experiences are key [72]. Each of these fields could benefit from the advanced capabilities of LLM-enhanced bipedal robots, particularly in environments requiring nuanced understanding and adaptability.
IV-D Applications in various fields
The advancements in bipedal locomotion technology hold significant promise for practical applications beyond the confines of laboratory environments. These robots, bolstered by AI, are poised to transform numerous sectors by enhancing operational capabilities and interaction with humans. The potential for humanoid robots in various fields is detailed in [72], emphasizing the integration of learning-based approaches for more effective implementation. Key areas include:
  1. 1.
    Industrial automation and manufacturing: The integration of humanoid robots in industrial settings can significantly enhance productivity and efficiency, freeing workers from repetitive and labor-intensive tasks. These robots, equipped with advanced loco-manipulation capabilities and the ability to cooperate with human teams, are particularly effective in assembly line operations, maintenance tasks, and the construction of complex machinery [73, 74]. Their articulated arms and floating bases provide unmatched flexibility, making them ideal for human-centric manufacturing environments. The humanoid robot Digit, for example, demonstrates remarkable stability and efficiency in industrial tasks over extended periods, as seen in video demonstrations [75]. Moreover, these robots are also suited for operation in high-risk environments such as underwater or areas with high radiation levels, significantly enhancing safety and operational capacity in these contexts.
  2. 2.
    Healthcare and assistive devices: In the healthcare sector, bipedal and humanoid robots contribute significantly to rehabilitation and assistive technologies. Exoskeletons enhanced with DRL methodologies are being used to train individuals to achieve more natural gait patterns, improving mobility and rehabilitation outcomes [76]. Beyond mere mobility aids, humanoid robots integrated with LLMs show promise in delivering medications, monitoring patient health, and assisting in surgeries [77]. The synergy between LLMs and loco-manipulation capabilities paves the way for more interactive, responsive support, aligning closely with the needs of personalized care. Additionally, the aging population can benefit from humanoid robots performing everyday tasks like house cleaning or delivery through simple voice commands, thereby enhancing the quality of life.
  3. 3.
    Search and rescue missions: Humanoid robots are exceptionally valuable in search and rescue operations, especially in disaster-stricken or hazardous environments where human presence is risky or impractical. Unlike traditional wheeled robots, humanoid robots can navigate complex terrains filled with debris, gaps, and elevated structures, making them indispensable in these scenarios. They also demonstrate potential for significant interaction and collaboration with human rescue teams. For instance, in environments with high nuclear radiation, humanoid robots can perform tasks that would be perilous for humans, handling delicate instruments and preventing human exposure to harmful conditions. This capability extends to other challenging environments such as underwater [78], outer space [79, 80], and other hazardous areas. However, the full realization of these applications remains constrained by the absence of a unified framework that can seamlessly navigate all terrains and fully integrate loco-manipulation and human interaction functionalities.
  4. 4.
    Entertainment and education: Humanoid robots have the potential to transform the realms of entertainment and education by providing highly interactive experiences. With their ability to integrate extensive knowledge bases, these robots can significantly enhance educational environments. They can assume the roles of butlers, teachers, or even babysitters, engaging with users in diverse activities. For example, robots can facilitate language learning [81], participate in storytelling, teach various academic subjects, or engage in the performing arts and games. In the sphere of entertainment, humanoid robots can act, dance, play ball games [69], and take part in interactive performances, captivating audiences of all ages with their versatility and dynamic capabilities.
    However, this in turn leads to a variety of ethical issues. First, interacting with humans involves collecting human daily behavior data and increases the risk of a data breach. Second, another concern is the increasing dependency of humans on robots, not just for assistance but also for emotional support. This will result in less human-to-human interaction and ultimately affect social constructs and emotional development. Third, advancements of humanoid robots will replace humans in various jobs, and eventually lead to unemployment issues.
    On the positive side, humanoid robots can provide invaluable assistance to people with disabilities or the elderly, offering companionship and reducing the care burden on families and healthcare systems. Furthermore, their application across diverse fields such as education, industry, and healthcare can bring about revolutionary changes, improving efficiency and safety while opening up new possibilities for technological integration. As we navigate these advancements, it is crucial to balance innovation with ethical considerations to ensure that the deployment of humanoid robots enhances societal well-being without compromising personal integrity or social dynamics.
V Conclusion
Despite significant advances in DRL for robotics, a considerable gap persists between current research achievements and the development of a unified framework capable of empowering robots to perform a broad spectrum of complex tasks efficiently. Presently, DRL research can be categorized into two primary control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have shown promising capabilities in executing diverse locomotion skills [22], climbing stairs [38], and navigating challenging terrains such as stepping stones [20]. Conversely, hybrid frameworks, which often integrate an HL planner or an LL model-based controller, offer enhanced capabilities, allowing for simultaneous management of locomotion and navigation tasks.
To bridge the existing gaps, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising. Such frameworks could simultaneously address issues of precision and generalization. Moreover, the advent of LLMs presents a transformative opportunity, potentially enabling the unification of language processing and visual functionalities within robotic systems. While numerous challenges remain—ranging from the technical intricacies of framework integration to real-world application—the steady progression in control framework refinement and DRL development provides a hopeful outlook. The vision of achieving an end-to-end unified framework, capable of mimicking human-like learning processes and enabling bipedal robots to handle a wide range of complex tasks, may soon move within reach.
Appendix A Deep reinforcement learning algorithms
The advancement and development of RL is crucial for bipedal locomotion. Specifically, advancements in deep learning provide deep neural networks serving as function approximators to empower RL with the capability to handle tasks characterized by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data. In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms. Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
Designing an effective neural network architecture is essential for tackling complex bipedal locomotion tasks. Multi-layer perceptrons (MLP), a fundamental neural network, excel in straightforward regression tasks with lower computational resource demands. A comprehensive comparison between MLP and the memory-based neural network, Long Short-Term Memory (LSTM) reveals that MLPs have an advantage in convergence speed for tasks [65]. However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modeling key physical properties vital for periodical gaits [19] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNN) specialize in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diverse range of neural network architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
Considering DRL alogirthms, recent bipedal locomotion studies focus on model-free reinforcement algorithms. Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model. Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimization) and value-based approaches. Value-based methods e.g. Q-learning, Deep Q-learning (DQN) [82] only excel in discrete action space and often struggle with high dimensional action space. In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
Figure 4: Diagram for RL algorithms catalogue
More advanced algorithms combine both policy-based methods and value-based methods. Actor-critic (AC) refers to a main idea simultaneously learning both a policy (actor) and a value function (critic), where it owns both advantages of both algorithms [83, 84]. Popular algorithms e.g. Trust region policy optimization (TRPO) [85] and PPO based on policy-based methods, borrow ideas from AC. Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [86], Twin Delayed Deep Deterministic Policy Gradients (TD3) [87], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [88], SAC (Soft Actor-Critic) [89]. Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario. There are several key factors to value these algorithms such as: sample efficiency, robustness and generalization, and implementation challenges. A comparative analysis work [90] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
In [91], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilizing fewer resources than TRPO. In terms of training time, PPO is much faster than SAC, and DDPG algorithms [90]. Besides, many works [19, 45, 36] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various neural network architectures have made PPO the most popular choice in this field. Various work has demonstrated that PPO can conduct the exploration of walking [19], jumping [37], stair climbing [38], and stepping stones [20], which demonstrates its efficiency, robustness and generalization.
Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimizing sampling efficiency. In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [21, 92]. TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [89]. Soft Actor-Critic (SAC) further the agent’s exploration capabilities and sample efficiency [89]. While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning. Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
Appendix B Bridging sim-to-real gap
Due to the large number of interactions needed for RL algorithms, training directly on robots can lead to costly damage to hardware and the environment. Consequently, training a policy in the simulation and then deploying it to the hardware illustrates significant potential and efficiency. However, the gap between simulation and the real world remains substantial, making sim-to-real challenging. To overcome the gap, several sim-to-real approaches are developed, including dynamics randomization [36], system identification [93, 94], periodic reward composition [66], learned actuators dynamics [93, 95], regulation feedback controller [96], adversarial motion prior [18, 17].
There are two primary approaches to training policies under domain randomization. One is end-to-end training with a history of robot measurements or I/O [36] and another is policy distillation, an expert policy with environmental insights guides a student policy that learns from internal sensory feedback, such as teacher-student policy [66], RMA [64].
Details of these sim-to-real transition approaches are shown below:
1) The dynamics randomization method involves systematically varying the physical parameters of the simulated environment-such as mass, inertia, or stiffness. 2) System Identification methods develop mathematical models of dynamics from observed data, enhancing the accuracy of robots’ properties within the models, such as mass and inertia, to ensure the model faithfully represents the system’s behavior. 3) The learned actuator dynamics method utilizes experimental data from actuators to develop a model of their dynamics, achieving sim-to-real by incorporating realistic actuator behavior within the training environment. It is noticeable that the higher-level planner can also learn from reference or non-reference. 4) periodic reward composition helps capture the essential locomotion information and the periodic gait pattern is more general to adapt to uncertainty and variation in the real world. 5) The regulation feedback controller manually tunes the setting of the controller to mitigate perturbations and gaps between the sim and the real, thereby enhancing the robustness and adaptation. Key aspects of sim-to-real, including system identification, state estimation with noise measurement, and the selection of state-and-action spaces are highlighted in [34].

### v3

@@
-Deep Reinforcement Learning for Bipedal Locomotion: A Brief Survey
+Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey
 Abstract
-Bipedal robots are garnering increasing global attention due to their potential applications and advancements in artificial intelligence, particularly in Deep Reinforcement Learning (DRL).
-While DRL has driven significant progress in bipedal locomotion, developing a comprehensive and unified framework capable of adeptly performing a wide range of tasks remains a challenge.
-This survey systematically categorizes, compares, and summarizes existing DRL frameworks for bipedal locomotion, organizing them into end-to-end and hierarchical control schemes.
-End-to-end frameworks are assessed based on their learning approaches, whereas hierarchical frameworks are dissected into layers that utilize either learning-based methods or traditional model-based approaches.
-This survey provides a detailed analysis of the composition, capabilities, strengths, and limitations of each framework type.
-Furthermore, we identify critical research gaps and propose future directions aimed at achieving a more integrated and efficient framework for bipedal locomotion, with potential broad applications in everyday life.
+Bipedal robots are gaining global recognition due to their potential applications and advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL).
+While DRL has significantly advanced bipedal locomotion, the development of a unified framework capable of handling a wide range of tasks remains an ongoing challenge.
+This survey systematically categorises, compares, and analyses existing DRL frameworks for bipedal locomotion, organising them into end-to-end and hierarchical control schemes.
+End-to-end frameworks are evaluated based on their learning approaches, while hierarchical frameworks are examined in terms of layered structures that integrate learning-based or traditional model-based methods.
+We provide a detailed evaluation of the composition, strengths, limitations, and capabilities of each framework.
+Additionally, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with wide-ranging applications in real-world environments.
 Index Terms:
@@
 I Introduction
-Humans navigate complex and varied environments, performing diverse locomotion tasks with only two legs.
-To facilitate dynamic bipedal locomotion, model-based methods were introduced in the 1980s and have since evolved significantly [1, 2, 3].
-These methods, characterized by rapid convergence, furnish a predictive framework for understanding environmental structures.
-However, they struggle to adapt in dynamically challenging environments that are difficult to model precisely.
-More recently, advancements in machine learning have provided novel approaches.
-Reinforcement learning (RL)-based methods, in particular, are adept at navigating the full dynamics of robot-environment interactions [4].
-Additionally, hybrid approaches that combine model-based and learning-based methods have been developed to leverage the advantages of both.
-Yet, the question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
+Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using just two legs.
+Bipedal robots, which closely mimic human form, have a distinct advantage over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains.
+Furthermore, bipedal humanoid robots are specifically designed to function in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use.
+This makes them highly adaptable to a wide range of tasks in such settings.
+As a result, bipedal robots hold significant potential for real-world applications [1].
+In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [2, 3, 4].
+Their agility is particularly advantageous in complex environments, such as multi-level workplaces.
+Bipedal robots are also well-suited for tasks that involve the use of human-designed tools, making them valuable for assisting in daily tasks, healthcare, and rehabilitation [5].
+Furthermore, they show considerable promise in search and rescue operations, where they can navigate hazardous and unpredictable terrains [6, 7, 8].
+Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [9, 10, 11].
+Early methods, such as the Linear Inverted Pendulum Model (LIPM) [12], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control.
+As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion.
+While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
+Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [13], which offers a distinct advantage.
+Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments.
+Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
 Figure 1: Common bipedal and humanoid robots used as platforms for testing DRL frameworks.
@@
 (e) Digit, a full human-sized 3D humanoid robot, an upgrade based on Cassie and actuated by torque control [18].
-To address this, we explore recent advancements in deep reinforcement learning (DRL)-based frameworks, categorizing control schemes into two primary types: (i) end-to-end and (ii) hierarchical.
-End-to-end frameworks map robot states directly to control outputs at the joint level, while hierarchical frameworks adopt a structured approach, decomposing decision-making into multiple layers.
-Here, a High-Level (HL) planner addresses navigation and path planning, while a Low-Level (LL) controller focuses on fundamental locomotion skills.
-The highest decision-making tier, the task level, receives direct input from task or user commands.
-Figure 2: Classification of DRL-based control schemes.
+Despite these advancements, a key question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
+If not, how should such a framework be designed, and what features or components should it include?
+To address these questions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical.
+These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
+End-to-end frameworks directly map robot states to joint-level control outputs, while hierarchical frameworks decompose decision-making into multiple layers.
+In hierarchical systems, a High-Level (HL) planner governs navigation and path planning, while a Low-Level (LL) controller handles fundamental locomotion tasks.
+The task-level decision-making tier interfaces directly with user commands or predefined tasks, forming a structured approach to robotic control.
 The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications.
-Although the application of RL to simple 2D bipedal robots began in 2004 [10, 11], it took several years before deep reinforcement learning (DRL) algorithms emerged.
+Although the application of RL to simple 2D bipedal robots began in 2004 [19, 20], it took several years before DRL algorithms emerged.
 These DRL-based methods have since shown promising results in physical simulators [21, 22, 23].
@@
 Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies.
-A notable framework employs a learned High-Level (HL) planner coupled with a Low-Level (LL) model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [23, 24, 25].
-Another innovative construction integrates a learned feedback controller with an HL planner, categorized under the DRL Feedback Control Hybrid Scheme [26, 27].
+A notable framework employs a learned HL planner coupled with a LL model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [32, 33, 34].
+Another innovative construction integrates a learned feedback controller with an HL planner, categorised under the DRL Feedback Control Hybrid Scheme [35, 36].
 Additionally, a learned hierarchical control scheme [37] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [21, 22, 38].
-While several review papers discuss RL for general robotics [4] and model-based methods for bipeds [1, 2, 3], none specifically focus on DRL-based frameworks for bipeds.
-This survey aims to address this gap by summarizing current research progress, highlighting the structure and capabilities of bipedal locomotion frameworks, and exploring future directions.
-We also catalogue DRL-based frameworks, as depicted in Fig.
+In order to provide a clearer understanding of the current landscape, we have categorised existing DRL frameworks, as depicted in Fig.
 2.
+Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure.
+Although bipedal locomotion tasks have evolved significantly in the past decade.
+Earlier frameworks primarily focused on basic walking stabilization and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation.
+Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [39], emphasizing real-world deployment and robustness.
+Although several reviews discuss RL for general robotics [13] and model-based methods for bipedal robots [9, 10, 11], none specifically focus on DRL-based frameworks for bipeds.
+To address this gap, this survey reviews relevant literature by applying the following selection criteria: (1) studies that examine DRL frameworks specifically for bipedal robots, (2) research involving both simulated and physical bipedal robots, (3) approaches that improve policy transfer from simulation to real-world environments, and (4) publications from the last five years (2018 - April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, ArXiv, and major robotics conferences such as CoRL, RSS, ICRA, IROS, and Humanoids.
+The search was conducted using the key terms “Deep Reinforcement Learning” or “Reinforcement Learning” in combination with “Bipedal Locomotion”, “Bipedal Walking”, “Biped Robot”, “Humanoid Robot”, or “Legged Robot”.
+The most relevant and impactful works were manually selected for further review.
+This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where necessary.
+For clarity, in this survey, the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
 The primary contributions of this survey are:
@@
   • The identification of current challenges and the provision of insightful future research directions.
-The paper is organized as follows: Section II focuses on end-to-end frameworks, categorized by learning approaches.
+The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches.
 Section III details hierarchical frameworks, classified into three main types.
-Section IV addresses existing gaps, ongoing challenges, and potential future research directions.
-Finally, Section V concludes the paper.
+Section IV presents the expected unified frameworks, which build upon proposed approaches while highlighting the challenges associated with their implementation.
+Section V addresses potential future research directions.
+Finally, Section VI concludes the survey.
+Figure 2: Classification of DRL-based control schemes.
 II End-to-end framework
-The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly translates sensory inputs—such as images, lidar data, or proprioceptive feedback [30]—along with user commands [19] or pre-defined references [31], into joint-level control actions.
-These actions encompass motor torques [32], positions, and velocities [15].
+The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, lidar data, or proprioceptive feedback [40], along with user commands [28] or pre-defined references [41], into joint-level control actions $\mathcal{U}$.
+Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function.
 This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
-End-to-end strategies primarily simplify the design of low-level tracking to basic elements, such as a Proportional-Derivative (PD) controller.
-These methods can be broadly categorized based on their reliance on prior knowledge into two types: reference-based and reference-free.
+End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a Proportional-Derivative (PD) controller.
+These methods can be broadly categorised based on their reliance on prior knowledge into two types: reference-based and reference-free.
 The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
@@
 Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
-Residual learning [33] [34] [35] Forward walk undirectional walk Omni-walk Add residential term to the known motor positions at the current time step.
+Residual learning [42] [43] Forward walk undirectional walk Add residual term to the known motor positions at the current time step.
 A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains.
@@
 Refer to caption
-Reference-free learning [19] [38] [20] Periodic motions Stepping stones Vision-based Learn locomotion skills from scratch without any prior knowledge.
+Reference-free learning [28] [46] [29] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge.
 A: High potential for gait exploration, high robust to the complicated terrain.
@@
 II-A Reference-based learning
-Reference-based learning utilizes prior knowledge, allowing the policy to develop locomotion skills by adhering to predefined references, which may be derived from trajectory optimization (TO) techniques or captured through motion capture systems.
-This method facilitates the acquisition of locomotion skills compared to alternative approaches, though it typically results in locomotion patterns that closely resemble the predefined references or motion clips, thus limiting the variety of gait patterns.
+Reference-based learning leverages prior knowledge generated offline through methods such as trajectory optimisation (TO) or motion capture systems.
+This predefined reference typically includes data related to the robot’s joint movements or pre-planned trajectories, serving as a foundation for the policy to develop locomotion skills by following these established motion patterns.
 Generally, this approach can be divided into two primary methods: (i) residual learning and (ii) guided learning.
 II-A1 Residual learning
-This method involves a framework that is aware of the current reference joint positions and applies offsets determined by the policy to modify motor commands at the current timestep.
-By utilizing predefined motion trajectories, the residual term acts as feedback control, compensating for errors and enabling the biped to achieve dynamic locomotion skills.
+The proposed framework utilises a policy that modifies motor commands by applying action offsets based on the current reference joint positions, allowing the biped robot to achieve dynamic locomotion through error compensation.
+The state space includes proprioceptive information such as trunk position, orientation, velocity, angular velocity, joint angles, and joint velocities, providing the necessary sensory data for real-time adjustments.
+Actions are defined by offsets, $\delta a$, which represent deviations from the predefined desired joint positions, $\hat{a}$, with the final motor commands represented as $a=\hat{a}+\delta a$.
+The reward function encourages the policy to optimise locomotion performance by considering (a) how closely the robot’s active joint angles match the reference angles, (b) how effectively the robot responds to user commands, and (c) additional terms that further enhance the stability of the robot’s movements.
+This holistic approach enables the biped robot to adapt to various dynamic conditions while maintaining balance and control.
 Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [42].
-This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimization (PPO) algorithms, as detailed in Appendix A.
+This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A.
 The policy receives the robot’s states and reference inputs, outputting a residual term that augments the reference at the current timestep.
-These modified references are then processed by a Proportional-Derivative (PD) controller to set the desired joint positions.
+These modified references are then processed by a PD controller to set the desired joint positions.
 While this framework enhanced the robot’s ability to perform tasks beyond standing [47], its physical deployment on a bipedal robot has not yet occurred, potentially rendering it impractical for managing walking at varying speeds and limiting movement to a single direction.
 To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied on a physical bipedal robot [43].
-This process and its key points are further explored in Appendix B.
 Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks.
 However, the robot’s movements remain constrained to merely walking forward or backward.
-A novel approach in residual learning was introduced to enable unidirectional walking, where the policy outputs a residual term added to the current positional states, facilitating gradual omnidirectional walking [35].
+A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [48].
+In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive.
+However, this also limits the policy’s ability to explore more diverse motions, restricting it to a single slow walking pattern.
+Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance.
+Compared to other learning approaches that directly output joint positions, it is highly sample-efficient [33].
+However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
 II-A2 Guided learning
-Guided learning trains policies to directly output the desired joint-level commands, eschewing the addition of a residual term.
-The reward structure in this approach is focused on closely imitating predefined references.
+Guided learning trains policies to directly output the desired joint-level commands as actions $a$, without relying on the addition of a residual term.
+The states space is as same as residual learning approach.
+In this approach, the reward structure is centred on accurately imitating predefined reference trajectories, ensuring precise alignment between the policy output and the reference motion.
 A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [44].
@@
 A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [41], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
-Despite these advancements, the parameterization of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances.
+Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances.
 To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [45].
 This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping maneuvers.
-An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviors [18].
+An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [27].
 This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
@@
 This shortfall stems from the framework’s dependency on adhering closely to the provided references, which narrows its learning capabilities.
-Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities.
+Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [41, 49].
 Additionally, this approach exhibits limited adaptability in responding effectively to unforeseen environmental changes or novel challenges.
@@
 This approach allows the policy to explore a wider range of gait patterns and adapt to unforeseen terrains, thereby enhancing innovation and flexibility within the learning process.
+The action space and observation space in this approach are similar to the guided-learning method; however, the reward structure differs significantly from the reference-based method.
+Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [49].
 The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models.
 A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [23].
-This framework introduced a novel term into the loss function and utilized a curriculum learning strategy to effectively shape gait patterns.
+This framework introduced a novel term into the loss function and utilised a curriculum learning strategy to effectively shape gait patterns.
 Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model, Cassie, though this has yet to be validated outside of simulation [50].
@@
 It has been further refined to adapt to diverse terrains and scenarios.
-For instance, robust blind walking on stairs was demonstrated through terrain randomization techniques in [38].
+For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [46].
 Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [51], thus enabling the robot to effectively navigate stepping stones [29].
-Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalizes terrain information [42].
+Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [52].
 This approach to learning enables the exploration of novel solutions and strategies that might not be achievable through mere imitation of existing behaviours.
@@
 Unlike end-to-end policies that directly map sensor inputs to motor outputs, hierarchical control schemes deconstruct locomotion challenges into discrete, manageable layers or stages of decision-making.
-Each layer within this structure is tasked with specific objectives, ranging from high-level navigation to fundamental locomotion skills.
+Each layer within this structure is tasked with specific objectives, ranging from navigation to fundamental locomotion skills.
 This division not only enhances the framework’s flexibility but also simplifies the problem-solving process for each policy.
-The architecture of a hierarchical framework typically comprises two principal modules: an HL planner and an LL controller.
+The architecture of a hierarchical framework typically comprises two principal modules: an High-Level (HL) planner and an Low-Level (LL) controller.
 This modular approach allows for the substitution of each component with either a model-based method or a learning-based policy, further enhancing adaptability and customisation to specific needs.
+Communication between the layers in a hierarchical framework is achieved through the transmission of commands.
+The HL planner sets abstract goals, which the LL controller translates into specific actions, such as calculating joint movements to follow a desired trajectory.
+In return, the robot sends sensor data back to the HL planner, enabling real-time adjustments.
+The tasks handled by different layers often operate on varying time scales, adding complexity to synchronising communication between the layers.
 Hierarchical frameworks can be classified into three distinct types based on the integration and function of their components:
   1. 1.
-    Deep planning hybrid scheme: This approach combines strategic, high-level planning with dynamic low-level execution, leveraging the strengths of both learning-based and traditional model-based methods.
+    Deep planning hybrid scheme: This approach combines strategic, HL planning with dynamic LL execution, leveraging the strengths of both learning-based and traditional model-based methods.
   2. 2.
-    Feedback DRL control hybrid scheme: It focuses on integrating direct feedback control mechanisms with deep reinforcement learning, allowing for real-time adjustments and enhanced responsiveness.
+    Feedback DRL control hybrid scheme: It focuses on integrating direct feedback control mechanisms with DRL, allowing for real-time adjustments and enhanced responsiveness.
   3. 3.
@@
 Deep Planning Hybrid Scheme [33] [53] [54] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
-A: Enhanced command tracking capabilities, generalized across different platforms, sampling efficiency, and robust.
-D: Complicate system and communication between layers, require precise model, lack generalization regarding different tasks.
+A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust.
+D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.
 Feedback DRL Control Hybrid Scheme [55] [35] [36] Gait library + feedback policy Footstep planner + feedback policy Model-based planner + feedback policy LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
@@
 This policy adeptly captures dynamic walking gaits through the use of reduced-order states and simplifies the learning trajectory.
-Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalized across various bipedal platforms, including Digit, Cassie, and RABBIT.
+Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalised across various bipedal platforms, including Digit, Cassie, and RABBIT.
 In parallel, several research teams have focused on developing navigation planners specifically for toy-like humanoid robots, which provide greater physical stability compared to torque-driven or hydraulic bipedal robots as shown in Fig.
@@
 One notable study [57] implemented a visual navigation policy on the NAO robot, depicted in Fig.
-1(a), utilizing RGB cameras as the primary sensory modality.
+1(a), utilising RGB cameras as the primary sensory modality.
 This system has demonstrated successful zero-shot transfer to real-world scenarios, enabling the robot to adeptly navigate around obstacles.
-Further research [44] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a center of mass (CoM) trajectory.
+Further research [54] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a centre of mass (CoM) trajectory.
 This configuration is coupled with a whole-body controller, facilitating dynamic activities like soccer shooting.
@@
 Consequently, this research primarily addresses navigation and task execution.
-Regarding generalization, these frameworks have shown potential for adaptation across different types of bipedal and humanoid robots with minimal adjustments, demonstrating advanced user command tracking [43] and sophisticated navigation capabilities [44].
+Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [53] and sophisticated navigation capabilities [54].
 However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping.
@@
 Enhancing both navigation and dynamic locomotion capabilities within the HL planner remains a significant challenge.
-Figure 3: Hierarchy Control Scheme Diagram: (a) A basic hierarchical scheme with two layers, where each module can be substituted with a learned policy.
-(b) A deep planning hybrid scheme, where the High-Level (HL) planner is learned.
-(c) A learning-based feedback control hybrid scheme, with a learned Low-Level (LL) controller.
-(d) A comprehensive DRL hierarchy control scheme, where both layers are learned.
+Figure 3: Hierarchy Control Scheme Diagram: presents a hierarchical control framework for a bipedal robot, consisting of a basic scheme and three variations.
+(1) Basic Scheme: The framework starts with a Task Command, followed by a HL Planner, and a LL Controller, which ultimately drives the Robot.
+Each module can be replaced by a learned policy, introducing adaptability in different control layers.
+(2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned.
+(b) A Feedback DRL control hybrid scheme, with a learned LL controller.
+(c) A Learned hierarchy control scheme, where both layers are learned.
 III-B Feedback DRL control hybrid scheme
@@
 However, this work has not yet been transferred from simulation to real-world applications.
-Later, a similar structure featuring an HL foot planner and an LL DRL policy was proposed [27].
+Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [36].
 This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
@@
 Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities.
-These policies have demonstrated high efficiency in locomotion and accurate reference tracking.
+These policies have demonstrated high efficiency in locomotion and accurate reference tracking [59].
 Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited.
-Nonetheless, these learning algorithms often face challenges in environments characterized by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [48].
+Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [59].
 Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [36].
@@
 Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models.
-Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimization and hinder rapid deployment and iterative enhancements [48].
+Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation and hinder rapid deployment and iterative enhancements [59].
 III-C Learned hierarchy framework
@@
 The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [21].
-In this setup, LL policies, informed by human motions or trajectories generated via Trajectory Optimization (TO), strive to track these trajectories as dictated by the HL planner while maintaining balance.
+In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance.
 An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths.
@@
 The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [22].
-However, despite its structured and layered approach, which allows for the reuse of learned behaviors to achieve long-term objectives, these frameworks have predominantly been validated only in simulations.
-The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviors, including stability issues like falling.
+However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks have predominantly been validated only in simulations.
+The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
 Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [38].
+To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [56].
+It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as a LL tracking controller.
+Additionally, binocular RGB vision input enhances perception, enabling precise loco-manipulation and dynamic locomotion tasks such as jumping, walking, folding clothes, and rearranging objects.
+This shadowing-based imitation learning approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
 Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics.
@@
 This combination enables the framework to adeptly manage a diverse array of environments and tasks.
-Within such a framework, the High-Level (HL) policy is optimized for strategic planning and achieving specific goals.
-This optimization allows for targeted enhancements depending on the tasks at hand.
+Within such a framework, HL policy is optimised for strategic planning and achieving specific goals.
+This optimisation allows for targeted enhancements depending on the tasks at hand.
 Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives.
 Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
-The transition from simulation to real-world scenarios is fraught with challenges, not least because of the complexities involved in training and integrating two separate layers within the control hierarchy.
-Ensuring effective communication and cooperation between these layers is critical, requiring a meticulously defined communication interface to avoid operational discrepancies.
-Additionally, the training process for each policy within the hierarchy demands considerable computational resources.
-The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalize to real-world conditions.
+Additionally, the training process for each policy within the hierarchy demands considerable computational resources [38].
+The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions.
 This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
-IV Challenges and future research directions
-While learning-based frameworks for bipedal robots have demonstrated considerable potential, they have also clearly exposed the limitations inherent to each framework.
-Moreover, several critical areas remain largely unexplored, especially within the realm of legged robotics, where the pace of research on bipedal robots lags behind that of their quadruped counterparts.
-This discrepancy in research progress can be attributed to several factors, including the higher costs and less mature technology associated with bipedal robot hardware, as well as the inherent instability issues that bipedal designs face.
-To gain a deeper understanding of these challenges and to outline potential future directions, it is instructive to first review existing research on quadruped robots.
-The insights gained from quadrupeds, which benefit from more robust research outputs and technological advancements, can provide valuable lessons for overcoming similar challenges in bipedal systems.
-IV-A Recent progress with quadruped robots
-While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems.
-The diversity of frameworks developed for quadrupeds ranges from model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [49, 50], to systems that include the modeling of deformable terrain to enhance locomotion over compliant surfaces [51].
-Furthermore, dynamic quadruped models facilitate highly adaptable policies [52], and sophisticated acrobatic motions are achieved through imitation learning [53].
-The domain of quadruped DRL has also seen significant advancements in complex hybrid frameworks that integrate vision-based systems.
-To date, two primary versions of such frameworks have been developed: one where a deep planning module is paired with model-based control [54], and another that combines model-based planning with low-level DRL control [48, 55].
-The latter has shown substantial efficacy; it employs a model predictive control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy.
-Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [56] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
-However, similar hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
-Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway.
-These include mimicking real animals through motion capture data and imitation learning [57, 58], as well as augmenting quadrupeds with manipulation abilities.
-This is achieved either by adding a manipulator [59, 60] or by using the robots’ legs [61].
-Notably, the research presented in [60] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
-Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks.
-Establishing a unified framework could bridge this gap, an essential step given the integral role of bipedal robots in developing full humanoid systems.
-Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
-IV-B Gaps and challenges
-Despite numerous promising developments in the field of bipedal and humanoid robotics, significant gaps remain between current research outcomes and the ultimate goals.
-This discussion concentrates on the gaps in frameworks and algorithms rather than hardware, structured around two pivotal questions: 1) Is it possible to design a unified framework that achieves both generalization and precision?
-2) Can we develop a straightforward end-to-end policy capable of managing all tasks efficiently?
-IV-B1 Generalization versus precision
+Besides, for the general hierarchical framework, the transition from simulation to real-world scenarios is challenging, particularly due to the complexities involved in coordinating two layers within the control hierarchy.
+Ensuring seamless communication and cooperation between the HL Planner and LL Controller is essential to avoid operational discrepancies.
+The primary challenges include: (1) Task division complexity—while the HL Planner handles strategy and provides abstract goals, the LL Controller manages precise execution, necessitating careful coordination to avoid functional overlap and conflicts.
+(2) Effective communication—the HL’s abstract goals must be accurately interpreted and converted by the LL into real-time actions, especially in dynamic environments.
+(3) Task allocation—clear division of responsibilities between layers is crucial to prevent redundancy and ensure smooth system performance.
+IV Towards a Unified Framework
+Previous Section II and Section III illustrate two distinct approaches for controlling bipedal robots.
+Whether it is end-to-end or hierarchical, their ultimate goal is to adeptly navigate all types of terrains Both these two finally towards to the unified frameworks, can be named as Humanoid Foundation Model (HFM) and Multi-layered Adaptive Models (MLAM), which are expanded based on end-to-end, and hierarchical approaches.
+In this section, we explore the conceptual design of HFM and MLAM, analyse the existing gaps and challenges in achieving these unified frameworks, and discuss recent advancements in quadruped robotics that offer valuable insights for the future development of bipedal locomotion.
+IV-A Humanoid Foundation Models
+Inspired by the Robot Foundation Models (RFMs) [60, 61], built on an end-to-end framework, enhance versatility by learning shared representations across different skills, HFM focus specifically on humanoid robots, which inherit advantages of RFMs but possess a higher number of DoFs and a human-like morphology, requiring more advanced motion control and adaptability.
+As foundation model, this structure bypasses the need for task decomposition, allowing the system to learn through large datasets and generalise across diverse locomotion tasks.
+HFM with large parameters are expected to be pre-trained on large-scale robotic datasets across various platforms and later fine-tuned using DRL to adapt to specific locomotive tasks.
+This enables humanoid robots to continuously refine their locomotion strategies and adapt to dynamic environments through interaction.
+For instance, AdA [62] demonstrates a framework where an agent is first pre-trained on extensive datasets and subsequently fine-tuned with RL, allowing it to optimise policy performance and achieve near-optimal control.
+Other foundation model work also provide thoughts about how to build up HFM.
+For example Current research on RFMs mainly about Vision-Language-Action (VLA) model on robotic arm platform, derived from VLM, VLA models directly map visual information to robot actions, enabling end-to-end decision-making without explicit intermediate representations.
+Recent work, such as RT-2 [63], expands upon the transformer model-based RT-1 [64], a vision-language model designed for robotic control.
+In RT-2, the entire model is trained using a combination of a VLM and robotic demonstrations.
+The actions need to be quantized as inputs and subsequently dequantized.
+Challenges: The challenges of RMFs can be derived from two aspects: (1) training datasets and (2) model inference time.
+  1. 1.
+    Training datasets: In terms of datasets, unlike Large Language Models (LLMs) or VLMs, HMF takes multimodal inputs with heterogeneous data formats that are not inherently aligned.
+While LLMs process textual prompts as both input and output, and VLMs typically map images and text bidirectionally, HFM for bipedal robot must handle diverse sensory data, including vision, proprioception, motion trajectories, and object interactions, making input representation and model alignment more complex.
+Which lead to input representation and feature alignment challenge, which means different data modalities have different feature distributions and dimensionalities, making direct fusion difficult.
+    Beside of directly making the training process difficult and high requirements of model construction, another challenge in datasets is datasets collection.
+Current datasets provide human motions for bipedal humanoid robot to imitate.
+For example, AMASS [65] consists solely of human motion capture data, while Motion-X [66] extends this by incorporating textual prompt labels corresponding to the motions.
+Although various references exist for robots to learn different skills, most datasets are not directly tailored for robots and require additional mapping to specific robotic platforms.
+  2. 2.
+    Model inference time: Bipedal robots require high-frequency action strategies to maintain stability, but real-time decision-making in HFM faces several limitations.
+First, RMFs using LLMs as backbones are computationally heavy, leading to slow inference due to expensive forward propagation.
+Second, autoregressive token generation introduces sequential dependencies, preventing parallel computation and further increasing latency.
+Third, onboard processors lack sufficient memory and computational power to run large models efficiently, making real-time execution challenging without model compression or hardware acceleration.
+IV-B Multi-layered Adaptive Model
+Unlike HFM, the MLAM follows a hierarchical framework, decomposing tasks into multiple sub-tasks.
+Each layer is responsible for a specific level of decision-making, with higher layers handling HL planning and lower layers executing fine-grained control.
+This structured approach enables the system to adapt dynamically and ultimately achieve the overall goal.
+Each module in the hierarchical model can be fine-tuned, replaced, or upgraded without requiring re-training of the entire model.
+The discrete model is structured into three hierarchical levels: strategy, planning, and control, each progressively refining task execution.
+Section III has demonstrated how DRL-based policies contribute to both the HL planner and LL controller.
+At the highest level, the strategy layer serves as the reasoning component, enabling the robot to comprehend tasks, infer objectives, and decompose long-term goal-based tasks into structured sub-tasks, which are then passed down for execution.
+For example, [67] employs an LLM as a reasoning layer to interpret user commands, break them into sequential task steps, and translate them into executable code for a quadruped robot.
+Challenge: In the previous Section LABEL:Sec:learned_heirarchical_framework, the challenges of the hierarchical framework were briefly discussed.
+However, when considering the unified framework—MLAM, additional challenges emerge that require further exploration.
+These challenges are:
+  1. 1.
+    Limited Visual Perception and Multi-Modal Integration: MLAM relies on real-time perception and decision-making, yet bipedal locomotion demands low-latency, high-precision sensory processing (e.g., vision, IMU, force sensors).
+Integrating these multi-modal inputs across different layers of the hierarchy remains a significant challenge.
+Each modality operates at different timescales and levels of abstraction, making seamless coordination and alignment difficult.
+Moreover, existing methods often fail to effectively capture cross-modal dependencies, limiting their applicability in dynamic scenarios.
+  2. 2.
+    Computational Efficiency and Real-Time Execution: MLAM must balance the quality of HL decision-making with the strict low-latency requirements of real-time control.
+This is especially problematic for HL reasoning models, such as LLMs, which are inherently designed for step-by-step reasoning and rely on autoregressive decoding.
+Generating decisions token-by-token significantly increases inference time, making these models unsuitable for tasks that demand rapid response times.
+These delays can lead to unstable behaviours, particularly in tasks requiring fast adaptation and precise coordination in dynamic environments.
+These challenges highlight the complexity of implementing MLAM and underscore the need for innovative solutions to address these limitations in real-world applications.
+IV-C Fundamental challenges
+The development of unified frameworks such as HFM and MLAM aims to address the challenges of bipedal locomotion in dynamic and unstructured environments.
+However, achieving such frameworks requires overcoming several fundamental obstacles.
+Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges.
+These challenges are explored in the following subsections.
+IV-C1 Generalisation versus precision
+The challenge of balancing generalisation and precision is a fundamental obstacle in achieving the unified frameworks of HFM and MLAM.
+For HFM, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments.
+For MLAM, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
 DRL has demonstrated potential in facilitating versatile locomotion skills [31]; however, challenges such as poor velocity tracking and issues with precise control often arise.
@@
 Yet, there is still no framework that comprehensively addresses the dual demands of versatility and precision in locomotion.
+For HFM, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks.
+In contrast, MLAM struggles to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
 The difficulty in simultaneously achieving precise control and a broad range of actions in bipedal locomotion using DRL stems from several factors:
@@
 Designing reward functions and training policies that satisfy both criteria is inherently complex.
-These challenges underscore the need for innovative solutions that can bridge the gap between the capabilities of current frameworks and the ambitious goals of advanced bipedal and humanoid robotics.
-IV-B2 Simplifying frameworks to overcome complex tasks
-The envisioned ideal in robotic design is an end-to-end framework that enables robots to traverse various terrains using versatile locomotion skills.
-Although current research often focuses on enhancing frameworks by adding complex components to mitigate inherent limitations, such as the integration of a foot planner for omnidirectional locomotion and stair navigation, as demonstrated in [26, 27], simpler end-to-end frameworks have also proven effective.
-These frameworks adeptly navigate challenging terrains and perform a diverse range of locomotion tasks with fewer components [20, 22].
+These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between HFM’s large-scale adaptability and MLAM’s hierarchical precision.
+IV-C2 Streamlined frameworks capable of handling complex tasks
+One of the critical challenges in unifying HFM and MLAM is designing frameworks that are both simple and capable of handling complex tasks.
+While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments.
+On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
+For HFM, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity.
+For MLAM, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency.
+Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
 The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure.
-To achieve an optimal end-to-end framework, advancements in several key areas are essential:
+To achieve an optimal framework, advancements in several key areas are essential:
   • Robust and efficient DRL algorithms: Development of algorithms that can manage high-dimensional and continuous control problems more effectively.
-  • Specialized neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [42].
-  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviors and strategic outcomes.
+  • Specialised neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [52].
+  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviours and strategic outcomes.
   • Advanced computational resources: Enhancement of computational capabilities to support more intensive training and faster inference, facilitating real-time decision-making in dynamic environments.
 By focusing on these developmental areas, the potential to create a unified, efficient, and less complex framework for handling complex locomotion challenges in bipedal robots is significantly increased.
-IV-C Future directions
-The exploration of quadruped robotics has yielded substantial advancements, yet the full potential of bipedal robotics remains largely untapped.
-Building on the successes and innovative approaches observed in quadruped robots, several key future directions emerge that could significantly enhance bipedal and humanoid robotics.
-IV-C1 Unified framework
-Currently, no single framework exists that enables bipedal or humanoid robots to adeptly navigate all types of terrains, including stepping stones, stairs, deformable terrain, and slippery surfaces.
-A promising approach, as evidenced by recent work in quadruped robots [48], utilizes MPC to generate reference motions, which a low-level DRL policy then tracks.
-This method, coupled with the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module, simplifies the terrain representation into a height map, facilitating more effective navigation.
-This success encourages further exploration into hybrid frameworks that combine model-based methods with DRL, inheriting the strengths of both approaches, as discussed in Section III.
-However, hybrid frameworks present challenges such as training efficiency and system complexity, which demand considerable computational resources and extensive training periods.
-Moreover, recent studies [20, 42, 22] have demonstrated the potential of end-to-end frameworks enhanced with vision-based information.
-These frameworks successfully navigate challenging terrains and execute dynamic motions, suggesting the feasibility of a unified framework capable of handling diverse environments and tasks.
-Training strategies such as curriculum learning and task randomization could be employed, utilizing visual height maps as inputs to the policy, enhancing the robot’s ability to adapt and perform in varied scenarios.
-In addition, the introduction of a DRL end-to-end framework incorporating transformer models, as in [62], presents significant possibilities for integrating locomotion skills with language and vision capabilities.
-The use of large-scale models capable of processing and condensing extensive data sets into a coherent model could expand the robot’s range of capabilities, maintaining versatility across a broad spectrum of tasks.
-The exploration of transformers and other large-scale models holds considerable promise for enhancing generalizability and adaptability in complex tasks, warranting further investigation into their potential applications in bipedal robotics.
-IV-C2 Vision-based learning framework
-Vision plays a critical role in enabling robots to navigate challenging terrains, such as blind drops, where tactile and other sensory inputs may not provide sufficient information.
-Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [38, 43].
-Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal and humanoid robots [20, 42, 27].
-Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
-  • Height scanner mapping: This approach, evaluated in works like [42], involves using height maps generated by scanners to inform locomotion strategies.
-These maps provide detailed topographical data, allowing robots to plan steps on uneven or obstructed surfaces more effectively.
-  • Direct vision inputs: Directly utilizing inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [46, 63].
-Although previous studies like [46] have integrated visual navigation by feeding visual information to a High-Level (HL) planner, the potential of direct visual inputs to RL policies has not been fully explored.
-Enhancing the capability of bipedal robots to directly interpret and utilize visual data without intermediary processing can revolutionize their adaptability and efficiency in real-world scenarios.
-The exploration of direct vision inputs to reinforcement learning policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
-IV-C3 Bridge the gap from simulation to reality
-While simulations offer a safe and cost-effective environment for developing robotics policies, the transition from simulation to real-world application often encounters significant challenges due to the approximations and simplifications made in simulations.
-Numerous sim-to-real frameworks [34, 64, 65] have shown high efficiency and performance, as detailed in Appendix B.
-Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments.
-Moreover, many studies [26, 66, 21] remain validated only in simulation settings.
-IV-C4 Loco-manipulation tasks
-Loco-manipulation, which combines locomotion and manipulation, presents opportunities for humanoid robots to excel beyond purely bipedal capabilities.
+Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks.
+The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion.
+To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities.
+These advancements will lead to the creation of two potential unified models: (1) a multi-layered adaptive model grounded in the hierarchical framework and (2) a large robot model built upon the end-to-end framework.
+The unified framework will ultimately support diverse locomotion tasks in complex environments.
+V Future Directions
+Section II and III introduced two mainstream learning-based frameworks, HFM and MLAM, and analysed their contributions toward achieving a unified framework.While Section IV provides a clearer understanding of the current state and the existing gap between these frameworks and the expected unified framework, further discussion is needed to outline specific directions for achieving this goal.
+In this section, we outline the essential components and future research directions necessary to bridge this gap.
+Specifically, we identify key elements that will drive the evolution from current framework limitations to a fully integrated unified framework.
+A summary of these fundamental elements is provided in Fig.
+4.
+V-A Future directions for bipedal robots
+V-A1 Loco-manipulation tasks
+With the rapid development of bipedal robotics, modern bipedal robots are now equipped with upper-body structures to enhance their capabilities beyond basic locomotion tasks.
+These advancements enable robots to perform upper-body-assisted locomotion and loco-manipulation tasks, expanding their adaptability in complex environments.
+Building on this evolution, upper-body-assisted locomotion enhances movement stability by integrating upper-lower body coordination, enabling robots to navigate challenging terrain more effectively.
+Beyond movement, manipulation tasks focus on precise upper-body control, requiring dexterous motion planning for object interaction and task execution.
+At the highest level of complexity, loco-manipulation integrates both locomotion and manipulation, demanding multi-contact interactions, dynamic balance maintenance, and real-time adaptation to unpredictable environments.
+These tasks necessitate full-body coordination and task-driven motion planning, making them essential benchmarks for advancing unified frameworks in humanoid robotics.
+Tasks include fundamental locomotion movements like flat walking and stepping over obstacles, progressing to upper-body-assisted locomotion challenges such as climbing ladders, wall-assisted walking, jumping over obstacles, and grasping for support.
+Beyond locomotion, precise upper-body control is essential for manipulation tasks like turning valves, driving vehicles, using tools, and connecting fire hoses.
+The most complex scenarios demand full-body coordination for loco-manipulation tasks such as carrying objects while walking, opening doors while navigating, and executing disaster-response operations.
+These benchmarks are critical for developing robust and adaptable humanoid systems, addressing key challenges posed in DARPA’s robotics initiatives [39].
+Therefore, loco-manipulation presents opportunities for humanoid robots to excel beyond purely bipedal capabilities.
 Few studies have addressed this integrated task; one such study [68] demonstrated a ’box transportation’ framework.
@@
 Moreover, the challenges of managing mobile tools like scooters [69] or dynamically interacting with objects such as balls [70] introduce further complexities.
-Decomposing loco-manipulation tasks into multiple layers could simplify the challenges, allowing for more precise and flexible control by manually tuning individual components of the task [43].
-This structured control approach provides a more coordinated response to complex interactions within the robot’s environment, facilitating the execution of task-specific commands.
-Alternatively, an end-to-end framework may enable bipeds to perform a variety of tasks through task randomization and structured curriculum learning methods, progressively teaching the policy [35, 27, 22].
+To address these challenges, a hierarchical control approach could provide a more structured and adaptable solution.
+Decomposing loco-manipulation into multiple layers would simplify task execution by allowing precise, modular control over individual components [53].
+This method enables manual fine-tuning of task-specific controllers, ensuring better stability and adaptability to environmental variations.
+Additionally, a structured hierarchical approach facilitates coordinated whole-body responses, improving the efficiency of loco-manipulation interactions.
+Alternatively, an end-to-end learning framework could offer a more generalised and scalable solution.
+By leveraging task randomisation and structured curriculum learning, bipedal robots could progressively learn loco-manipulation skills in a unified policy [48, 36, 31].
 During training, such policies can also learn human-like movements from motion capture data [71, 27, 25], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
-IV-C5 Designing reward functions
-The development of effective reward functions is a critical challenge in the field of deep reinforcement learning (DRL) for bipedal robots.
-While periodic reward functions have been designed to facilitate cyclic movements like walking [19], there remains a significant gap in crafting reward functions for non-periodic actions such as jumping.
-These actions require distinct considerations for success and efficiency, yet current research lacks comprehensive methods for their reward structure.
-Furthermore, minimizing the need for extensive manual tuning while achieving high performance in DRL systems continues to be a substantial challenge, pointing to the need for more adaptive and automatically adjusting reward mechanisms.
-IV-C6 Integrating large language models
-The integration of Large Language Models (LLMs) into bipedal robotics opens new avenues for contextual understanding and task execution, significantly enhancing the robots’ interaction capabilities.
-LLMs, when implemented at the highest task level, offer substantial promise for improving human-robot interaction, making these systems more intuitive and responsive [71].
-The potential applications of this technology are broad and impactful, spanning sectors such as industrial automation, where robots can perform complex assembly tasks; healthcare, offering assistance in patient care and rehabilitation; assistive devices, providing support for individuals with disabilities; search and rescue operations, where robust and adaptive decision-making is critical; and entertainment and education, where interactive and engaging experiences are key [72].
-Each of these fields could benefit from the advanced capabilities of LLM-enhanced bipedal robots, particularly in environments requiring nuanced understanding and adaptability.
-IV-D Applications in various fields
-The advancements in bipedal locomotion technology hold significant promise for practical applications beyond the confines of laboratory environments.
-These robots, bolstered by AI, are poised to transform numerous sectors by enhancing operational capabilities and interaction with humans.
-The potential for humanoid robots in various fields is detailed in [72], emphasizing the integration of learning-based approaches for more effective implementation.
-Key areas include:
-  1. 1.
-    Industrial automation and manufacturing: The integration of humanoid robots in industrial settings can significantly enhance productivity and efficiency, freeing workers from repetitive and labor-intensive tasks.
-These robots, equipped with advanced loco-manipulation capabilities and the ability to cooperate with human teams, are particularly effective in assembly line operations, maintenance tasks, and the construction of complex machinery [73, 74].
-Their articulated arms and floating bases provide unmatched flexibility, making them ideal for human-centric manufacturing environments.
-The humanoid robot Digit, for example, demonstrates remarkable stability and efficiency in industrial tasks over extended periods, as seen in video demonstrations [75].
-Moreover, these robots are also suited for operation in high-risk environments such as underwater or areas with high radiation levels, significantly enhancing safety and operational capacity in these contexts.
-  2. 2.
-    Healthcare and assistive devices: In the healthcare sector, bipedal and humanoid robots contribute significantly to rehabilitation and assistive technologies.
-Exoskeletons enhanced with DRL methodologies are being used to train individuals to achieve more natural gait patterns, improving mobility and rehabilitation outcomes [76].
-Beyond mere mobility aids, humanoid robots integrated with LLMs show promise in delivering medications, monitoring patient health, and assisting in surgeries [77].
-The synergy between LLMs and loco-manipulation capabilities paves the way for more interactive, responsive support, aligning closely with the needs of personalized care.
-Additionally, the aging population can benefit from humanoid robots performing everyday tasks like house cleaning or delivery through simple voice commands, thereby enhancing the quality of life.
-  3. 3.
-    Search and rescue missions: Humanoid robots are exceptionally valuable in search and rescue operations, especially in disaster-stricken or hazardous environments where human presence is risky or impractical.
-Unlike traditional wheeled robots, humanoid robots can navigate complex terrains filled with debris, gaps, and elevated structures, making them indispensable in these scenarios.
-They also demonstrate potential for significant interaction and collaboration with human rescue teams.
-For instance, in environments with high nuclear radiation, humanoid robots can perform tasks that would be perilous for humans, handling delicate instruments and preventing human exposure to harmful conditions.
-This capability extends to other challenging environments such as underwater [78], outer space [79, 80], and other hazardous areas.
-However, the full realization of these applications remains constrained by the absence of a unified framework that can seamlessly navigate all terrains and fully integrate loco-manipulation and human interaction functionalities.
-  4. 4.
-    Entertainment and education: Humanoid robots have the potential to transform the realms of entertainment and education by providing highly interactive experiences.
-With their ability to integrate extensive knowledge bases, these robots can significantly enhance educational environments.
-They can assume the roles of butlers, teachers, or even babysitters, engaging with users in diverse activities.
-For example, robots can facilitate language learning [81], participate in storytelling, teach various academic subjects, or engage in the performing arts and games.
-In the sphere of entertainment, humanoid robots can act, dance, play ball games [69], and take part in interactive performances, captivating audiences of all ages with their versatility and dynamic capabilities.
-    However, this in turn leads to a variety of ethical issues.
-First, interacting with humans involves collecting human daily behavior data and increases the risk of a data breach.
-Second, another concern is the increasing dependency of humans on robots, not just for assistance but also for emotional support.
-This will result in less human-to-human interaction and ultimately affect social constructs and emotional development.
-Third, advancements of humanoid robots will replace humans in various jobs, and eventually lead to unemployment issues.
-    On the positive side, humanoid robots can provide invaluable assistance to people with disabilities or the elderly, offering companionship and reducing the care burden on families and healthcare systems.
-Furthermore, their application across diverse fields such as education, industry, and healthcare can bring about revolutionary changes, improving efficiency and safety while opening up new possibilities for technological integration.
-As we navigate these advancements, it is crucial to balance innovation with ethical considerations to ensure that the deployment of humanoid robots enhances societal well-being without compromising personal integrity or social dynamics.
-V Conclusion
-Despite significant advances in DRL for robotics, a considerable gap persists between current research achievements and the development of a unified framework capable of empowering robots to perform a broad spectrum of complex tasks efficiently.
-Presently, DRL research can be categorized into two primary control schemes: end-to-end and hierarchical frameworks.
-End-to-end frameworks have shown promising capabilities in executing diverse locomotion skills [22], climbing stairs [38], and navigating challenging terrains such as stepping stones [20].
-Conversely, hybrid frameworks, which often integrate an HL planner or an LL model-based controller, offer enhanced capabilities, allowing for simultaneous management of locomotion and navigation tasks.
-To bridge the existing gaps, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising.
-Such frameworks could simultaneously address issues of precision and generalization.
+V-A2 Bridging the gap from simulation to reality
+Due to the large number of interactions needed for DRL algorithms, training directly on robots can lead to costly damage to hardware and the environment.
+Consequently, training a policy in the simulation which offer a safe and cost-effective environment for developing robotics policies, and then deploying it to the hardware illustrates significant potential and efficiency.
+Bridging this gap is a key step toward achieving a unified framework, as it allows for the implementation of more versatile and robust systems that can handle diverse tasks across different terrains.
+Furthermore, the ultimate goal of achieving zero-shot transfer, where a policy trained in simulation can be deployed without additional real-world fine-tuning, directly contributes to the unified framework’s adaptability and scalability.
+This capability would make the framework far more practical for handling the wide range of locomotion tasks expected from future bipedal robots.
+Creating a reliable simulation environment and conducting RL training is a complex task.
+Advanced simulators are essential for training policies using full-order dynamics, which better represent the complex interactions robots face in the real world.
+Simulators like Isaac Gym [72], RoboCup3D [73], OpenAI Gym [74], and MuJoCo [75] are widely used to train policies in environments that closely mimic physical conditions.
+These tools help minimise the sim-to-real gap, improving the success of policy transfer to real robots.
+Further details on these simulators can be found in Appendix B.
+However, the process of sim-to-real encounters significant challenges due to the approximations and simplifications made in simulations.
+The process of sim-to-real faces significant challenges due to the approximations and simplifications made in simulations.
+Two primary issues arise: (1) Inaccurate dynamic models, contribute to discrepancies between simulation environments and real-world dynamics.
+They often simplify or generalise complex physical interactions, leading to a gap in performance when transitioning to reality.
+(2) The real environment is more diverse and complex, with varying terrain, sensor noise, and unpredictable conditions not captured in simulations.
+Numerous sim-to-real frameworks [43, 76, 77] have shown high efficiency and performance.
+Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments.
+To bridge the sim-to-real gap, several approaches have been developed, including dynamics randomisation [44], system identification [78, 79], periodic reward composition [49], learned actuator dynamics [78, 80], regulation feedback controllers [81], and adversarial motion priors [27, 26].
+In domain randomisation, there are two main methods for training policies: (1) end-to-end training using robot measurement histories, and (2) policy distillation, where an expert policy (e.g., teacher-student [49]) guides a student policy that learns from internal sensory feedback, as seen in techniques like RMA [76].
+Future research in bridging the sim-to-real gap will focus on enhancing the transferability and robustness of policies across diverse environments, ultimately aiming for zero-shot transfer, where no real-world fine-tuning is required.
+This will not only involve improving hardware but also making simulation environments more reflective of real-world conditions.
+Additionally, advancing sophisticated vision-based perception and multimodal sensory integration will be crucial in further minimising the sim-to-real gap, making policies more adaptable to complex and unpredictable environments.
+V-A3 Integrating foundation models
+Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots as discussed in Section IV.
+Their integration into robotics extends beyond decision-making and offers new ways to enhance control, adaptability, and learning efficiency.
+HL task planning: Foundation models, particularly LLMs and VLMs, enable robots to process multimodal inputs, enhancing task planning and execution.
+By integrating visual and textual data, these models enhance context-awareness, allowing robots to make more informed, adaptive decisions.
+For instance, Vision-Language Model Predictive Control (VLM-PC) [82] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
+Similarly, in hierarchical frameworks, VLMs can act as an intermediary layer, processing sensory data before passing structured commands to lower-level controllers.
+By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
+Future research should focus on integrating LLMs with DRL, where LLMs guide HL strategy while DRL refines real-time execution, creating an adaptive, flexible control loop.
+This integration would enable robots to generalize across tasks, adapt dynamically, and make autonomous complex decisions.
+Language-Assisted DRL: The design of effective reward functions is a fundamental challenge in DRL for bipedal robots.
+While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem.
+These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming.
+Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
+LLMs can enhance RL by guiding policy optimisation and refining reward functions.
+A key challenge in DRL is reward function design, which typically requires manual tuning.
+LLMs offer a solution by dynamically adjusting reward signals based on linguistic descriptions of task success.
+For example, Kumar et al.
+[83] demonstrate how LLMs refine reward functions by incorporating linguistic feedback, improving hierarchical RL frameworks.
+The language model processes HL instructions, which are then translated into lower-level RL tasks, simplifying complex skill acquisition and human-robot interactions.
+Conversely, end-to-end approaches such as AnyBipe [84] use LLMs to autonomously adjust reward functions and control strategies, reducing manual intervention.
+This enables more autonomous learning, allowing the model to self-optimise in diverse environments and handle complex terrain navigation.
+Other implementations of foundation models in RL have been comprehensively reviewed in [60].
+Future research should further explore the synergy between LLMs and DRL, with a particular focus on adaptive reward shaping and self-optimising control policies.
+Advancing these methodologies could pave the way for more autonomous, self-learning humanoid robots, capable of dynamically refining their decision-making processes and adapting seamlessly to diverse environments with minimal human intervention.
+V-A4 Vision-based learning framework
+The vision-based learning framework plays a crucial role in advancing the unified framework for bipedal robots.
+Integrating vision into the unified system enhances the robot’s ability to perceive and interpret its environment, making it capable of handling more complex and dynamic tasks, such as navigating rough terrains or interacting with objects in real-time.
+Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [46, 53].
+This can be attributed to several factors: (1) Simplified Problem: Blind locomotion reduces task complexity, focusing directly on motor control strategies rather than processing visual inputs.
+(2) Computation and Training Complexity: Integrating vision requires significant computational resources and robust algorithms for real-time processing, terrain identification, and object detection.
+(3) Advanced Control Without Vision: Previous studies have already demonstrated robust blind locomotion using proprioceptive data, such as joint angles, IMU, with vision typically enhancing precision in locomotion [52] or path planning [57].
+However, recent advancements in computing power and computer vision technology—such as more efficient real-time visual processing algorithms and increasingly powerful hardware—have significantly accelerated the integration of vision-based learning into robotics.
+This shift has led to a surge in the popularity of vision-based approaches, allowing robots to more effectively interpret and adapt to complex environments, making them a central component in modern frameworks.
+Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal robots [29, 52, 36].
+Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
+  • Height scanner mapping: This approach, evaluated in works like [52], involves using height maps generated by scanners to inform locomotion strategies.
+These maps provide detailed topographical data, allowing robots to plan steps on uneven or obstructed surfaces more effectively.
+  • Direct vision inputs: Directly utilising inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [57, 85].
+Although previous studies like [57] have integrated visual navigation by feeding visual information to a HL planner, the potential of direct visual inputs to RL policies has not been fully explored.
+Enhancing the capability of bipedal robots to directly interpret and utilise visual data without intermediary processing can revolutionise their adaptability and efficiency in real-world scenarios.
+The exploration of direct vision inputs to RL-based policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
+V-B Insights from quadruped robots
+While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems.
+The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [86, 87], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [88].
+Furthermore, dynamic quadruped models facilitate highly adaptable policies [89], and sophisticated acrobatic motions are achieved through imitation learning [90].
+The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems.
+To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [91] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [59, 92] within a feedback DRL control hybrid scheme.
+The latter has shown substantial efficacy; it employs a Model Predictive Control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy.
+Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [93] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
+However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
+Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway.
+These include mimicking real animals through motion capture data and imitation learning [94, 95], as well as augmenting quadrupeds with manipulation abilities.
+This is achieved either by adding a manipulator [96, 97] or by using the robots’ legs [98].
+Notably, the research presented in [97] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
+Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers and high mechanical complexity can be attributed to the cause of this disparity between quadruped and bipedal robots.
+Establishing a unified framework could bridge this gap, an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems.
+Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
+VI Conclusion
+Despite significant progress in DRL for robotics, a significant gap remains between current achievements and the development of a unified framework that can efficiently handle a wide range of complex tasks.
+DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks.
+End-to-end frameworks have demonstrated success in handling diverse locomotion skills [31], climbing stairs [46], and navigating challenging terrains such as stepping stones [29].
+Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
+Each framework brings unique strengths to the unified framework.
+End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills.
+In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies.
+Together, these approaches form the basis for two potential unified frameworks: the Humanoid Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
+To bridge the existing gaps and achieve a unified framework, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising.
+Such frameworks could simultaneously address issues of precision and generalisation.
 Moreover, the advent of LLMs presents a transformative opportunity, potentially enabling the unification of language processing and visual functionalities within robotic systems.
 While numerous challenges remain—ranging from the technical intricacies of framework integration to real-world application—the steady progression in control framework refinement and DRL development provides a hopeful outlook.
-The vision of achieving an end-to-end unified framework, capable of mimicking human-like learning processes and enabling bipedal robots to handle a wide range of complex tasks, may soon move within reach.
+The vision of achieving such unified framework may soon move within reach.
 Appendix A Deep reinforcement learning algorithms
+Figure 5: Diagram for RL algorithms catalogue
 The advancement and development of RL is crucial for bipedal locomotion.
-Specifically, advancements in deep learning provide deep neural networks serving as function approximators to empower RL with the capability to handle tasks characterized by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data.
+Specifically, advancements in deep learning provide deep NNs serving as function approximators to empower RL with the capability to handle tasks characterised by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data.
 In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms.
 Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
-Designing an effective neural network architecture is essential for tackling complex bipedal locomotion tasks.
-Multi-layer perceptrons (MLP), a fundamental neural network, excel in straightforward regression tasks with lower computational resource demands.
-A comprehensive comparison between MLP and the memory-based neural network, Long Short-Term Memory (LSTM) reveals that MLPs have an advantage in convergence speed for tasks [65].
-However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modeling key physical properties vital for periodical gaits [19] and successful sim-to-real transfer in bipedal locomotion.
-Additionally, Convolutional Neural Networks (CNN) specialize in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial.
-This diverse range of neural network architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
-Considering DRL alogirthms, recent bipedal locomotion studies focus on model-free reinforcement algorithms.
+Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks.
+Multi-layer perceptrons (MLP), a fundamental NN, excel in straightforward regression tasks with lower computational resource demands.
+A comprehensive comparison between MLP and the memory-based NN, LSTM reveals that MLPs have an advantage in convergence speed for tasks [77].
+However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modelling key physical properties vital for periodical gaits [28] and successful sim-to-real transfer in bipedal locomotion.
+Additionally, Convolutional Neural Networks (CNN) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial.
+This diverse range of NN architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
+Considering DRL algorithms, recent bipedal locomotion studies focus on model-free RL algorithms.
 Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model.
 Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
-Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimization) and value-based approaches.
+Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches.
 Value-based methods e.g.
-Q-learning, Deep Q-learning (DQN) [82] only excel in discrete action space and often struggle with high dimensional action space.
+Q-learning, State-Action-Reward-State-Action (SARSA), Deep Q-learning (DQN) [99] only excel in discrete action space and often struggle with high dimensional action space.
+Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy.
+SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning.
+DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to the difficulty of accurate value estimation.
 In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
-Figure 4: Diagram for RL algorithms catalogue
 More advanced algorithms combine both policy-based methods and value-based methods.
@@
 Popular algorithms e.g.
-Trust region policy optimization (TRPO) [85] and PPO based on policy-based methods, borrow ideas from AC.
+Trust region policy optimisation (TRPO) [102] and PPO based on policy-based methods, borrow ideas from AC.
 Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [103], Twin Delayed Deep Deterministic Policy Gradients (TD3) [104], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [105], SAC (Soft Actor-Critic) [106].
 Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario.
-There are several key factors to value these algorithms such as: sample efficiency, robustness and generalization, and implementation challenges.
+There are several key factors to value these algorithms such as: sample efficiency, robustness and generalisation, and implementation challenges.
 A comparative analysis work [107] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
-In [91], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilizing fewer resources than TRPO.
+In [108], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilising fewer resources than TRPO.
 In terms of training time, PPO is much faster than SAC, and DDPG algorithms [107].
-Besides, many works [19, 45, 36] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various neural network architectures have made PPO the most popular choice in this field.
-Various work has demonstrated that PPO can conduct the exploration of walking [19], jumping [37], stair climbing [38], and stepping stones [20], which demonstrates its efficiency, robustness and generalization.
-Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimizing sampling efficiency.
+Besides, many works [28, 55, 44] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various NN architectures have made PPO the most popular choice in this field.
+Various work has demonstrated that PPO can conduct the exploration of walking [28], jumping [45], stair climbing [46], and stepping stones [29], which demonstrates its efficiency, robustness and generalisation.
+Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimising sampling efficiency.
 In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [30, 109].
 TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [106].
-Soft Actor-Critic (SAC) further the agent’s exploration capabilities and sample efficiency [89].
+SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3.
+Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences to train without requiring constant interactions with the environment.
+This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential.
+SAC is also known for its stability and strong performance across a wide range of tasks [106].
 While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning.
 Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
-Appendix B Bridging sim-to-real gap
-Due to the large number of interactions needed for RL algorithms, training directly on robots can lead to costly damage to hardware and the environment.
-Consequently, training a policy in the simulation and then deploying it to the hardware illustrates significant potential and efficiency.
-However, the gap between simulation and the real world remains substantial, making sim-to-real challenging.
-To overcome the gap, several sim-to-real approaches are developed, including dynamics randomization [36], system identification [93, 94], periodic reward composition [66], learned actuators dynamics [93, 95], regulation feedback controller [96], adversarial motion prior [18, 17].
-There are two primary approaches to training policies under domain randomization.
-One is end-to-end training with a history of robot measurements or I/O [36] and another is policy distillation, an expert policy with environmental insights guides a student policy that learns from internal sensory feedback, such as teacher-student policy [66], RMA [64].
-Details of these sim-to-real transition approaches are shown below:
-1) The dynamics randomization method involves systematically varying the physical parameters of the simulated environment-such as mass, inertia, or stiffness.
-2) System Identification methods develop mathematical models of dynamics from observed data, enhancing the accuracy of robots’ properties within the models, such as mass and inertia, to ensure the model faithfully represents the system’s behavior.
-3) The learned actuator dynamics method utilizes experimental data from actuators to develop a model of their dynamics, achieving sim-to-real by incorporating realistic actuator behavior within the training environment.
-It is noticeable that the higher-level planner can also learn from reference or non-reference.
-4) periodic reward composition helps capture the essential locomotion information and the periodic gait pattern is more general to adapt to uncertainty and variation in the real world.
-5) The regulation feedback controller manually tunes the setting of the controller to mitigate perturbations and gaps between the sim and the real, thereby enhancing the robustness and adaptation.
-Key aspects of sim-to-real, including system identification, state estimation with noise measurement, and the selection of state-and-action spaces are highlighted in [34].
+Appendix B Training Simulation Environment
+The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators.
+Creating a reliable simulation environment and conducting RL training is difficult.
+Through the literature, several simulators are available, including Isaac Gym [72], robocup3D [73], Openai Gym [74], MuJoCo [75].
+OpenAI developed Gym and Gymnasium providing a simplified environment for fast testing RL algorithms with the simplified bipedal model.
+Also, RoboCup serves as a platform for RL research and development.
+On the other hand, MuJoCo and Gazebo developed by DeepMind respectively are comprehensive physics engines that support robotics research and development, along with other applications needing physics simulation.
+Isaac Gym, a cutting-edge physics simulation and RL framework developed by NVIDIA, is a high-performance platform for training agents in intricate and lifelike environments.
+One of the most crucial things is the parallelisation strategy and GPU simulations.
+For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms with particular emphasis on simulations that require large numbers of environment instances executing in parallel.
+Running the physics simulation on GPU can result in significant speedups, especially for large scenes with thousands of individual actors.

### v4

@@
 IV Towards a Unified Framework
-Previous Section II and Section III illustrate two distinct approaches for controlling bipedal robots.
-Whether it is end-to-end or hierarchical, their ultimate goal is to adeptly navigate all types of terrains Both these two finally towards to the unified frameworks, can be named as Humanoid Foundation Model (HFM) and Multi-layered Adaptive Models (MLAM), which are expanded based on end-to-end, and hierarchical approaches.
-In this section, we explore the conceptual design of HFM and MLAM, analyse the existing gaps and challenges in achieving these unified frameworks, and discuss recent advancements in quadruped robotics that offer valuable insights for the future development of bipedal locomotion.
-IV-A Humanoid Foundation Models
-Inspired by the Robot Foundation Models (RFMs) [60, 61], built on an end-to-end framework, enhance versatility by learning shared representations across different skills, HFM focus specifically on humanoid robots, which inherit advantages of RFMs but possess a higher number of DoFs and a human-like morphology, requiring more advanced motion control and adaptability.
-As foundation model, this structure bypasses the need for task decomposition, allowing the system to learn through large datasets and generalise across diverse locomotion tasks.
-HFM with large parameters are expected to be pre-trained on large-scale robotic datasets across various platforms and later fine-tuned using DRL to adapt to specific locomotive tasks.
-This enables humanoid robots to continuously refine their locomotion strategies and adapt to dynamic environments through interaction.
-For instance, AdA [62] demonstrates a framework where an agent is first pre-trained on extensive datasets and subsequently fine-tuned with RL, allowing it to optimise policy performance and achieve near-optimal control.
-Other foundation model work also provide thoughts about how to build up HFM.
-For example Current research on RFMs mainly about Vision-Language-Action (VLA) model on robotic arm platform, derived from VLM, VLA models directly map visual information to robot actions, enabling end-to-end decision-making without explicit intermediate representations.
-Recent work, such as RT-2 [63], expands upon the transformer model-based RT-1 [64], a vision-language model designed for robotic control.
-In RT-2, the entire model is trained using a combination of a VLM and robotic demonstrations.
-The actions need to be quantized as inputs and subsequently dequantized.
-Challenges: The challenges of RMFs can be derived from two aspects: (1) training datasets and (2) model inference time.
-  1. 1.
-    Training datasets: In terms of datasets, unlike Large Language Models (LLMs) or VLMs, HMF takes multimodal inputs with heterogeneous data formats that are not inherently aligned.
-While LLMs process textual prompts as both input and output, and VLMs typically map images and text bidirectionally, HFM for bipedal robot must handle diverse sensory data, including vision, proprioception, motion trajectories, and object interactions, making input representation and model alignment more complex.
-Which lead to input representation and feature alignment challenge, which means different data modalities have different feature distributions and dimensionalities, making direct fusion difficult.
-    Beside of directly making the training process difficult and high requirements of model construction, another challenge in datasets is datasets collection.
-Current datasets provide human motions for bipedal humanoid robot to imitate.
-For example, AMASS [65] consists solely of human motion capture data, while Motion-X [66] extends this by incorporating textual prompt labels corresponding to the motions.
-Although various references exist for robots to learn different skills, most datasets are not directly tailored for robots and require additional mapping to specific robotic platforms.
-  2. 2.
-    Model inference time: Bipedal robots require high-frequency action strategies to maintain stability, but real-time decision-making in HFM faces several limitations.
-First, RMFs using LLMs as backbones are computationally heavy, leading to slow inference due to expensive forward propagation.
-Second, autoregressive token generation introduces sequential dependencies, preventing parallel computation and further increasing latency.
-Third, onboard processors lack sufficient memory and computational power to run large models efficiently, making real-time execution challenging without model compression or hardware acceleration.
-IV-B Multi-layered Adaptive Model
-Unlike HFM, the MLAM follows a hierarchical framework, decomposing tasks into multiple sub-tasks.
-Each layer is responsible for a specific level of decision-making, with higher layers handling HL planning and lower layers executing fine-grained control.
-This structured approach enables the system to adapt dynamically and ultimately achieve the overall goal.
-Each module in the hierarchical model can be fine-tuned, replaced, or upgraded without requiring re-training of the entire model.
-The discrete model is structured into three hierarchical levels: strategy, planning, and control, each progressively refining task execution.
-Section III has demonstrated how DRL-based policies contribute to both the HL planner and LL controller.
-At the highest level, the strategy layer serves as the reasoning component, enabling the robot to comprehend tasks, infer objectives, and decompose long-term goal-based tasks into structured sub-tasks, which are then passed down for execution.
-For example, [67] employs an LLM as a reasoning layer to interpret user commands, break them into sequential task steps, and translate them into executable code for a quadruped robot.
-Challenge: In the previous Section LABEL:Sec:learned_heirarchical_framework, the challenges of the hierarchical framework were briefly discussed.
-However, when considering the unified framework—MLAM, additional challenges emerge that require further exploration.
-These challenges are:
-  1. 1.
-    Limited Visual Perception and Multi-Modal Integration: MLAM relies on real-time perception and decision-making, yet bipedal locomotion demands low-latency, high-precision sensory processing (e.g., vision, IMU, force sensors).
-Integrating these multi-modal inputs across different layers of the hierarchy remains a significant challenge.
-Each modality operates at different timescales and levels of abstraction, making seamless coordination and alignment difficult.
-Moreover, existing methods often fail to effectively capture cross-modal dependencies, limiting their applicability in dynamic scenarios.
-  2. 2.
-    Computational Efficiency and Real-Time Execution: MLAM must balance the quality of HL decision-making with the strict low-latency requirements of real-time control.
-This is especially problematic for HL reasoning models, such as LLMs, which are inherently designed for step-by-step reasoning and rely on autoregressive decoding.
-Generating decisions token-by-token significantly increases inference time, making these models unsuitable for tasks that demand rapid response times.
-These delays can lead to unstable behaviours, particularly in tasks requiring fast adaptation and precise coordination in dynamic environments.
-These challenges highlight the complexity of implementing MLAM and underscore the need for innovative solutions to address these limitations in real-world applications.
+Previous Sections II and III outlined two distinct paradigms for controlling bipedal robot locomotion: end-to-end and hierarchical frameworks.
+While both paradigms have demonstrated significant success in addressing specific locomotion tasks, their ability to generalise across diverse robot morphologies, environmental conditions, and locomotion objectives remains limited.
+Consequently, this limitation motivates us to revisit the central question posed in Section I:
+Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
+If not, how should such a framework be designed, and what features or components should it include?
+To address this, we first explicitly define a unified framework for bipedal locomotion as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions, ultimately aiming to develop genuinely generalist locomotion policies that achieve robust zero-shot or few-shot generalisation to previously unseen scenarios, embodiments, and objectives.
+Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these unified criteria.
+Typically, these methods require significant redesign or retraining when introduced to new morphologies, tasks, or environmental conditions, highlighting the practical gap towards achieving genuine unification.
+To explore pathways towards unified frameworks, we identify two currently foreseeable and potentially viable directions that build upon the structural foundations of the previously discussed paradigms—end-to-end and hierarchical.
+These two directions offer distinct but complementary strategies for advancing towards scalable, generalisable locomotion architectures.
+  • Bipedal Foundation Models (BFMs) – large-scale, pre-trained models that map perception directly to action through representation learning.
+Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
+  • Multi-Layer Adaptive Frameworks (MLAFs) – modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs.
+MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
+In the following subsections, we present detailed conceptual analyses of BFMs and MLAFs, evaluate their potential to realise a unified framework, and subsequently discuss fundamental challenges inherent to existing frameworks.
+IV-A Bipedal Foundation Models
+Inspired by Robot Foundation Models [60, 61], BFMs are large-scale, general-purpose models tailored for bipedal locomotion.
+Trained on diverse datasets spanning various morphologies, terrains, and objectives, BFMs enable reasoning about novel scenarios and adaptation with minimal task-specific supervision.
+A second-stage DRL fine-tuning process further adapts these models to specific platforms, supporting versatile locomotion behaviours such as walking and turning.
+Building upon the end-to-end reinforcement learning methods introduced in Section II, which typically focus on task-specific policy training, BFMs retain RL as a core adaptation mechanism, but shift toward a unified and generalisable paradigm by introducing shared motion priors across tasks and morphologies.
+This reuse of RL in a fine-tuning role allows BFMs to ground HL priors into executable LL policies, while improving sample efficiency and transferability.
+The BFMs architecture includes a multi-modal embedding module, a latent backbone, and an action decoder.
+The embedding encodes morphology, task specs, and optional environment input into a shared space.
+The backbone converts these into a temporally structured latent representation conditioned on the task.
+The decoder outputs LL control commands adapted to the robot.
+Similar modular structures appear in RT-1 [62] and RT-2 [63], which map multimodal inputs to discrete actions using a shared transformer backbone.
+The model first leverages IL to pre-train generalisable motion priors from diverse demonstrations, resulting in latent representations that facilitate scenario reasoning and transfer across different robot morphologies.
+DRL then fine-tunes these priors by aligning them with the physical constraints, action spaces, and dynamics of specific platforms.
+Rather than learning from scratch, DRL enables efficient adaptation to new tasks.
+Lightweight strategies such as ControlNet-style modular layers such as AdaptNet [64]—allow task-specific adjustments with minimal parameter updates.
+Beyond modular adaptation, recent works have demonstrated that large-scale DRL fine-tuning significantly enhances generalisation and capability transfer.
+FLaRe [65] fine-tunes behaviour-cloned policies with sparse rewards to achieve high success rates on long-horizon mobile manipulation tasks, including generalisation to novel embodiments.
+Complementarily, MOTO [66] applies a model-based DRL framework that incorporates on-policy value expansion and epistemic uncertainty estimation, enabling effective offline-to-online adaptation from image observations.
+Other frameworks further highlight this paradigm.
+AdA [67] demonstrates that scaling DRL with memory-based policies allows for in-context adaptation to novel embodied environments.
+RoboFUME [68] combines VLM-based reward classification with DRL fine-tuning to autonomously adapt pre-trained manipulation policies without environment resets.
+Collectively, these approaches underscore DRL not only as a fine-tuning tool, but as a central mechanism for grounding foundation model priors into executable, platform-specific control policies.
+While DRL enables flexible and powerful adaptation of pre-trained motion priors, applying it in bipedal foundation models introduces several challenges.
+Fine-tuning often demands carefully shaped rewards, frequent environment resets, and large amounts of interaction data—factors that are especially costly and risky on physical bipedal robots [68].
+Additionally, DRL policies may overfit to narrow dynamics or forget generalisable priors acquired during pre-training [65, 67].
+These issues constrain the robustness, efficiency, and safety of DRL-based adaptation in real-world settings.
+Meanwhile, existing datasets such as AMASS [69] and Motion-X [70] remain human-centric and require substantial retargeting effort [25].
+The inference latency of large models also limits real-time deployment on hardware-constrained platforms [63].
+IV-B Multi-layered Adaptive Models
+Unlike BFMs, which rely on large-scale pre-training and shared latent representations, MLAMs adopt a modular hierarchical approach, extending conventional hierarchical frameworks in Section III with explicitly adaptive layers.
+MLAMs decompose the locomotion pipeline into three distinct adaptive tiers: (1) HL reasoning for task decomposition and goal interpretation, (2) mid-level planning for behaviour generation, and (3) LL control for precise command execution.
+This modularity allows each layer to be independently optimised, upgraded, or replaced without retraining the entire system.
+MLAMs distinguish themselves from conventional hierarchical approaches through their adaptive modularity, dynamic policy integration, and clear layer-wise semantic responsibilities.
+A key distinguishing feature of MLAMs lies in its capacity to dynamically compose adaptive modules for each control tier.
+Instead of generalising through unified embeddings as in BFMs, MLAMs achieve scalability by dynamically assembling specialised policies.
+Each layer processes context-specific inputs and outputs interpretable commands.
+The HL reasoning layer typically uses symbolic reasoning or Large Language Models (LLMs) [71, 72] to parse commands into sub-tasks.
+For instance, Vision-Language Model Predictive Control [73] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
+By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
+The mid-level planner selects or synthesises motions via learned motion libraries [55, 41] or DRL-based planners [74].
+The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands.
+These include locomotion primitives like walking and climbing [75], adaptive tracking controllers for whole-body tracking [56], and imitation-based skills such as kicking and dancing [22].
+Controllers at this level utilise RL, IL, or model-based methods [9, 76] and operate at high frequency to ensure robust real-time execution across varied conditions.
+An example of MLAMs architecture is demonstrated by recent hierarchical systems using LLMs for quadrupedal tasks [75].
+HL modules interpret instructions, mid-level planners select motion primitives, and LL controllers execute precise actions, enabling robust dynamic replanning and flexible real-world operation.
+Despite these strengths, MLAMs deployment faces challenges distinct from BFMs.
+MLAMs depend on real-time multi-modal perception such as vision, IMU, force sensors, complicating data alignment across layers due to differing timescales and abstraction levels [76].
+Additionally, HL reasoning modules like LLMs often incur computational latency from autoregressive inference, posing limitations for tasks needing rapid reactions [75].
+Thus, integrating multi-modal perception and ensuring real-time execution are primary challenges for effective MLAMs implementation.
+Taken together, MLAMs offer a complementary path to BFMs by emphasising modularity, interpretability, and adaptability through dynamic policy composition.
+While they avoid the scalability bottlenecks of monolithic models, their reliance on real-time multi-modal processing and multi-layer coordination introduces unique system-level challenges.
 IV-C Fundamental challenges
-The development of unified frameworks such as HFM and MLAM aims to address the challenges of bipedal locomotion in dynamic and unstructured environments.
+The development of unified frameworks such as BFMs and MLAMs aim to address the challenges of bipedal locomotion in dynamic and unstructured environments.
 However, achieving such frameworks requires overcoming several fundamental obstacles.
@@
 These challenges are explored in the following subsections.
-IV-C1 Generalisation versus precision
-The challenge of balancing generalisation and precision is a fundamental obstacle in achieving the unified frameworks of HFM and MLAM.
-For HFM, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments.
-For MLAM, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
+IV-C1 Generalisation and precision
+The challenge of simultaneously improving generalisation across diverse skills and precision in specific tasks, such as jumping precisely to a desired distance, remains a fundamental obstacle in realising unified frameworks like BFMs and MLAMs.
+For BFMs, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments.
+For MLAMs, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
 DRL has demonstrated potential in facilitating versatile locomotion skills [31]; however, challenges such as poor velocity tracking and issues with precise control often arise.
@@
 Yet, there is still no framework that comprehensively addresses the dual demands of versatility and precision in locomotion.
-For HFM, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks.
-In contrast, MLAM struggles to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
+For BFMs, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks.
+In contrast, MLAMs struggle to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
 The difficulty in simultaneously achieving precise control and a broad range of actions in bipedal locomotion using DRL stems from several factors:
@@
 Designing reward functions and training policies that satisfy both criteria is inherently complex.
-These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between HFM’s large-scale adaptability and MLAM’s hierarchical precision.
+These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between BFM’s large-scale adaptability and MLAM’s hierarchical precision.
 IV-C2 Streamlined frameworks capable of handling complex tasks
-One of the critical challenges in unifying HFM and MLAM is designing frameworks that are both simple and capable of handling complex tasks.
+One of the critical challenges in unifying BFMs and MLAMs is designing frameworks that are both simple and capable of handling complex tasks.
 While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments.
 On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
-For HFM, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity.
-For MLAM, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency.
+For BFMs, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity.
+For MLAMs, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency.
 Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
@@
 To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities.
-These advancements will lead to the creation of two potential unified models: (1) a multi-layered adaptive model grounded in the hierarchical framework and (2) a large robot model built upon the end-to-end framework.
+These advancements will lead to the creation of two potential unified models: (1) a Multi-Layered Adaptive Model grounded in the hierarchical framework and (2) a Bipedal Foundation Model built upon the end-to-end framework.
 The unified framework will ultimately support diverse locomotion tasks in complex environments.
@@
 V-A3 Integrating foundation models
-Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots as discussed in Section IV.
-Their integration into robotics extends beyond decision-making and offers new ways to enhance control, adaptability, and learning efficiency.
-HL task planning: Foundation models, particularly LLMs and VLMs, enable robots to process multimodal inputs, enhancing task planning and execution.
-By integrating visual and textual data, these models enhance context-awareness, allowing robots to make more informed, adaptive decisions.
-For instance, Vision-Language Model Predictive Control (VLM-PC) [82] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
-Similarly, in hierarchical frameworks, VLMs can act as an intermediary layer, processing sensory data before passing structured commands to lower-level controllers.
-By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
-Future research should focus on integrating LLMs with DRL, where LLMs guide HL strategy while DRL refines real-time execution, creating an adaptive, flexible control loop.
-This integration would enable robots to generalize across tasks, adapt dynamically, and make autonomous complex decisions.
-Language-Assisted DRL: The design of effective reward functions is a fundamental challenge in DRL for bipedal robots.
+Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots, as discussed in Section IV.
+Their integration into robotics extends beyond decision-making, offering new opportunities to enhance control, adaptability, and learning efficiency.
+One particularly promising direction is their use in RL, not merely for planning but for shaping the learning process itself—most notably in the design of reward functions, which remains a core challenge in bipedal locomotion tasks.
+The design of effective reward functions is a fundamental challenge in DRL for bipedal robots.
 While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem.
@@
 In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies.
-Together, these approaches form the basis for two potential unified frameworks: the Humanoid Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
+Together, these approaches form the basis for two potential unified frameworks: the Bipedal Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
 To bridge the existing gaps and achieve a unified framework, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising.
@@
 Creating a reliable simulation environment and conducting RL training is difficult.
-Through the literature, several simulators are available, including Isaac Gym [72], robocup3D [73], Openai Gym [74], MuJoCo [75].
-OpenAI developed Gym and Gymnasium providing a simplified environment for fast testing RL algorithms with the simplified bipedal model.
-Also, RoboCup serves as a platform for RL research and development.
-On the other hand, MuJoCo and Gazebo developed by DeepMind respectively are comprehensive physics engines that support robotics research and development, along with other applications needing physics simulation.
-Isaac Gym, a cutting-edge physics simulation and RL framework developed by NVIDIA, is a high-performance platform for training agents in intricate and lifelike environments.
+Through the literature, several simulators are available, including Isaac Gym [81], robocup3D [82], Openai Gym [83], MuJoCo [84], Orbit [117], Brax [118], Isaac Lab [119].
+OpenAI developed Gym and Gymnasium to provide lightweight environments for rapid testing of RL algorithms, including simplified bipedal locomotion models.
+RoboCup also serves as a benchmark platform for RL research and development in multi-agent settings.
+For physics-based simulation, MuJoCo developed by DeepMind, and Gazebo are widely used platforms that support a range of robotics research tasks.
+NVIDIA’s Isaac Gym, although now deprecated, played an important role as a high-performance GPU-based simulator for training agents in complex environments.
+Its successors, such as Isaac Lab and Orbit, continue to evolve as modern RL and robotics frameworks.
 One of the most crucial things is the parallelisation strategy and GPU simulations.

### v5

@@
 Furthermore, they show considerable promise in search and rescue operations, where they can navigate hazardous and unpredictable terrains [6, 7, 8].
-Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [9, 10, 11].
-Early methods, such as the Linear Inverted Pendulum Model (LIPM) [12], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control.
-As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion.
-While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
-Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [13], which offers a distinct advantage.
-Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments.
-Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
 Figure 1: Common bipedal and humanoid robots used as platforms for testing DRL frameworks.
@@
 (e) Digit, a full human-sized 3D humanoid robot, an upgrade based on Cassie and actuated by torque control [13].
+Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [14, 15, 16].
+Early methods, such as the Linear Inverted Pendulum Model (LIPM) [17], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control.
+As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion.
+Advanced methods such as Model Predictive Control (MPC) [18, 19] and Trajectory Optimisation (TO) [20, 21, 14] exploit a predefined dynamics model to solve constrained optimal-control problems planning footsteps, CoM motion, and contact forces.
+While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
+Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [22], which offers a distinct advantage.
+Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments.
+Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
 Despite these advancements, a key question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
 If not, how should such a framework be designed, and what features or components should it include?
-To address these questions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical.
+To provide a clear benchmark for answering this question, we first define a unified framework as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions.
+The ultimate aim is to develop genuinely generalist locomotion policies that achieve robust zero-shot generalisation to previously unseen scenarios, embodiments, and objectives.
+With this definition as our guide, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical.
 These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
@@
 2.
+Figure 2: Classification of DRL-based control schemes.
+The approaches are broadly categorised into two main paradigms: End-to-end frameworks, which learn a single policy from sensory inputs to motor commands, and Hierarchical frameworks, which decompose the control problem into multiple levels.
+Within the end-to-end paradigm, a key distinction is between Reference-free learning (learning from scratch) and Reference-based learning (tracking a predefined motion).
+Hierarchical structures include Hybrid control schemes, which synergistically combine learned components with traditional model-based controllers.
 Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure.
 Although bipedal locomotion tasks have evolved significantly in the past decade.
-Earlier frameworks primarily focused on basic walking stabilization and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation.
+Earlier frameworks primarily focused on basic walking stabilisation and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation.
 Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [42], emphasizing real-world deployment and robustness.
@@
 Section III details hierarchical frameworks, classified into three main types.
-Section IV presents the expected unified frameworks, which build upon proposed approaches while highlighting the challenges associated with their implementation.
-Section V addresses potential future research directions.
+Section IV presents specific limitations and challenges, highlighting their connection to the previously surveyed content.
+Section V addresses potential future pathways, opportunities, and two proposed conceptual models that extend the end-to-end and hierarchical paradigms.
 Finally, Section VI concludes the survey.
-Figure 2: Classification of DRL-based control schemes.
 II End-to-end framework
@@
 To facilitate an understanding of these distinctions, Table I provides a succinct overview of the frameworks discussed.
-TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework. The dashed line in the implementation flow chat refers to optional.
-Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
-Residual learning [42] [43] Forward walk undirectional walk Add residual term to the known motor positions at the current time step.
-A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains.
-Refer to caption
-Guided learning [44] [41] [45] [31] Forward walk Versatile walk Versatile jump Versatile motions Mimic the predefined reference and directly specifies joint-level command.
-A: Accelerate the learning process and robust to the terrains.
-D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment.
-Refer to caption
-Reference-free learning [28] [46] [29] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge.
-A: High potential for gait exploration, high robust to the complicated terrain.
-D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources.
-Refer to caption
-00footnotetext: Forward Walk involves the bipeds walking straight ahead.
+TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework.
+The dashed line in the implementation flow chat refers to optional.00footnotetext: Forward Walk involves the bipeds walking straight ahead.
 Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities.
@@
 Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
+Methods
+Works
+Capabilities
+Characteristic
+Advantages and Disadvantages
+Implementation Flow Chart
+Residual learning
+[45]
+[46]
+Forward walk
+undirectional walk
+Add residual term to the known motor positions at the current time step.
+A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains.
+Guided learning
+[47]
+[44]
+[48]
+[35]
+Forward walk
+Versatile walk
+Versatile jump
+Versatile motions
+Mimic the predefined reference and directly specifies joint-level command.
+A: Accelerate the learning process and robust to the terrains. D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment.
+Reference-free learning
+[32]
+[49]
+[33]
+Periodic motions
+Stepping stones walk
+Visual walk
+Learn locomotion skills from scratch without any prior knowledge.
+A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources.
 II-A Reference-based learning
@@
 TABLE II: Summary and comparison of Hierarchy framework
-Control Scheme Works Module Characteristic Advantages and Disadvantages
-Deep Planning Hybrid Scheme [33] [53] [54] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
+Control Scheme
+Works
+Module
+Characteristic
+Advantages and Disadvantages
+Deep Planning Hybrid Scheme
+[36]
+[56]
+[57]
+Deep planning + ID
+Deep planning + ID-QP
+Deep planning + WPG
+HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
 A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust.
 D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.
-Feedback DRL Control Hybrid Scheme [55] [35] [36] Gait library + feedback policy Footstep planner + feedback policy Model-based planner + feedback policy LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
-A: Short inference times, robust, navigation locomotion capabilities, interpretability.
-D: Complicated system and communication between layers, reducing sampling efficiency.
-Learned Hierarchy Framework [21] [38] [56] HL policy + LL policy HL policy + LL policy HL policy + LL policy Both HL planner and LL feedback controller are learned.
-LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills.
+Feedback DRL Control Hybrid Scheme
+[58]
+[38]
+[39]
+Gait library + feedback policy
+Footstep planner + feedback policy
+Model-based planner + feedback policy
+LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
+A: Short inference times, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, reducing sampling efficiency.
+Learned Hierarchy Framework
+[25]
+[41]
+[59]
+HL policy + LL policy
+HL policy + LL policy
+HL policy + LL policy
+Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills.
 A: provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy.
@@
 (2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned.
-(b) A Feedback DRL control hybrid scheme, with a learned LL controller.
+(b) A Feedback DRL control hybrid scheme, with a learned $L$ controller.
 (c) A Learned hierarchy control scheme, where both layers are learned.
@@
 (3) Task allocation—clear division of responsibilities between layers is crucial to prevent redundancy and ensure smooth system performance.
-IV Towards a Unified Framework
-Previous Sections II and III outlined two distinct paradigms for controlling bipedal robot locomotion: end-to-end and hierarchical frameworks.
-While both paradigms have demonstrated significant success in addressing specific locomotion tasks, their ability to generalise across diverse robot morphologies, environmental conditions, and locomotion objectives remains limited.
-Consequently, this limitation motivates us to revisit the central question posed in Section I:
-Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
-If not, how should such a framework be designed, and what features or components should it include?
-To address this, we first explicitly define a unified framework for bipedal locomotion as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions, ultimately aiming to develop genuinely generalist locomotion policies that achieve robust zero-shot or few-shot generalisation to previously unseen scenarios, embodiments, and objectives.
-Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these unified criteria.
-Typically, these methods require significant redesign or retraining when introduced to new morphologies, tasks, or environmental conditions, highlighting the practical gap towards achieving genuine unification.
-To explore pathways towards unified frameworks, we identify two currently foreseeable and potentially viable directions that build upon the structural foundations of the previously discussed paradigms—end-to-end and hierarchical.
-These two directions offer distinct but complementary strategies for advancing towards scalable, generalisable locomotion architectures.
-  • Bipedal Foundation Models (BFMs) – large-scale, pre-trained models that map perception directly to action through representation learning.
-Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
-  • Multi-Layer Adaptive Frameworks (MLAFs) – modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs.
-MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
-In the following subsections, we present detailed conceptual analyses of BFMs and MLAFs, evaluate their potential to realise a unified framework, and subsequently discuss fundamental challenges inherent to existing frameworks.
-IV-A Bipedal Foundation Models
-Inspired by Robot Foundation Models [60, 61], BFMs are large-scale, general-purpose models tailored for bipedal locomotion.
-Trained on diverse datasets spanning various morphologies, terrains, and objectives, BFMs enable reasoning about novel scenarios and adaptation with minimal task-specific supervision.
-A second-stage DRL fine-tuning process further adapts these models to specific platforms, supporting versatile locomotion behaviours such as walking and turning.
-Building upon the end-to-end reinforcement learning methods introduced in Section II, which typically focus on task-specific policy training, BFMs retain RL as a core adaptation mechanism, but shift toward a unified and generalisable paradigm by introducing shared motion priors across tasks and morphologies.
-This reuse of RL in a fine-tuning role allows BFMs to ground HL priors into executable LL policies, while improving sample efficiency and transferability.
-The BFMs architecture includes a multi-modal embedding module, a latent backbone, and an action decoder.
-The embedding encodes morphology, task specs, and optional environment input into a shared space.
-The backbone converts these into a temporally structured latent representation conditioned on the task.
-The decoder outputs LL control commands adapted to the robot.
-Similar modular structures appear in RT-1 [62] and RT-2 [63], which map multimodal inputs to discrete actions using a shared transformer backbone.
-The model first leverages IL to pre-train generalisable motion priors from diverse demonstrations, resulting in latent representations that facilitate scenario reasoning and transfer across different robot morphologies.
-DRL then fine-tunes these priors by aligning them with the physical constraints, action spaces, and dynamics of specific platforms.
-Rather than learning from scratch, DRL enables efficient adaptation to new tasks.
-Lightweight strategies such as ControlNet-style modular layers such as AdaptNet [64]—allow task-specific adjustments with minimal parameter updates.
-Beyond modular adaptation, recent works have demonstrated that large-scale DRL fine-tuning significantly enhances generalisation and capability transfer.
-FLaRe [65] fine-tunes behaviour-cloned policies with sparse rewards to achieve high success rates on long-horizon mobile manipulation tasks, including generalisation to novel embodiments.
-Complementarily, MOTO [66] applies a model-based DRL framework that incorporates on-policy value expansion and epistemic uncertainty estimation, enabling effective offline-to-online adaptation from image observations.
-Other frameworks further highlight this paradigm.
-AdA [67] demonstrates that scaling DRL with memory-based policies allows for in-context adaptation to novel embodied environments.
-RoboFUME [68] combines VLM-based reward classification with DRL fine-tuning to autonomously adapt pre-trained manipulation policies without environment resets.
-Collectively, these approaches underscore DRL not only as a fine-tuning tool, but as a central mechanism for grounding foundation model priors into executable, platform-specific control policies.
-While DRL enables flexible and powerful adaptation of pre-trained motion priors, applying it in bipedal foundation models introduces several challenges.
-Fine-tuning often demands carefully shaped rewards, frequent environment resets, and large amounts of interaction data—factors that are especially costly and risky on physical bipedal robots [68].
-Additionally, DRL policies may overfit to narrow dynamics or forget generalisable priors acquired during pre-training [65, 67].
-These issues constrain the robustness, efficiency, and safety of DRL-based adaptation in real-world settings.
-Meanwhile, existing datasets such as AMASS [69] and Motion-X [70] remain human-centric and require substantial retargeting effort [25].
-The inference latency of large models also limits real-time deployment on hardware-constrained platforms [63].
-IV-B Multi-layered Adaptive Models
-Unlike BFMs, which rely on large-scale pre-training and shared latent representations, MLAMs adopt a modular hierarchical approach, extending conventional hierarchical frameworks in Section III with explicitly adaptive layers.
-MLAMs decompose the locomotion pipeline into three distinct adaptive tiers: (1) HL reasoning for task decomposition and goal interpretation, (2) mid-level planning for behaviour generation, and (3) LL control for precise command execution.
-This modularity allows each layer to be independently optimised, upgraded, or replaced without retraining the entire system.
-MLAMs distinguish themselves from conventional hierarchical approaches through their adaptive modularity, dynamic policy integration, and clear layer-wise semantic responsibilities.
-A key distinguishing feature of MLAMs lies in its capacity to dynamically compose adaptive modules for each control tier.
-Instead of generalising through unified embeddings as in BFMs, MLAMs achieve scalability by dynamically assembling specialised policies.
-Each layer processes context-specific inputs and outputs interpretable commands.
-The HL reasoning layer typically uses symbolic reasoning or Large Language Models (LLMs) [71, 72] to parse commands into sub-tasks.
-For instance, Vision-Language Model Predictive Control [73] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
-By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
-The mid-level planner selects or synthesises motions via learned motion libraries [55, 41] or DRL-based planners [74].
-The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands.
-These include locomotion primitives like walking and climbing [75], adaptive tracking controllers for whole-body tracking [56], and imitation-based skills such as kicking and dancing [22].
-Controllers at this level utilise RL, IL, or model-based methods [9, 76] and operate at high frequency to ensure robust real-time execution across varied conditions.
-An example of MLAMs architecture is demonstrated by recent hierarchical systems using LLMs for quadrupedal tasks [75].
-HL modules interpret instructions, mid-level planners select motion primitives, and LL controllers execute precise actions, enabling robust dynamic replanning and flexible real-world operation.
-Despite these strengths, MLAMs deployment faces challenges distinct from BFMs.
-MLAMs depend on real-time multi-modal perception such as vision, IMU, force sensors, complicating data alignment across layers due to differing timescales and abstraction levels [76].
-Additionally, HL reasoning modules like LLMs often incur computational latency from autoregressive inference, posing limitations for tasks needing rapid reactions [75].
-Thus, integrating multi-modal perception and ensuring real-time execution are primary challenges for effective MLAMs implementation.
-Taken together, MLAMs offer a complementary path to BFMs by emphasising modularity, interpretability, and adaptability through dynamic policy composition.
-While they avoid the scalability bottlenecks of monolithic models, their reliance on real-time multi-modal processing and multi-layer coordination introduces unique system-level challenges.
-IV-C Fundamental challenges
-The development of unified frameworks such as BFMs and MLAMs aim to address the challenges of bipedal locomotion in dynamic and unstructured environments.
-However, achieving such frameworks requires overcoming several fundamental obstacles.
-Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges.
-These challenges are explored in the following subsections.
-IV-C1 Generalisation and precision
-The challenge of simultaneously improving generalisation across diverse skills and precision in specific tasks, such as jumping precisely to a desired distance, remains a fundamental obstacle in realising unified frameworks like BFMs and MLAMs.
-For BFMs, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments.
-For MLAMs, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
-DRL has demonstrated potential in facilitating versatile locomotion skills [31]; however, challenges such as poor velocity tracking and issues with precise control often arise.
-While [53] shows that deep planning combined with model-based control can achieve precise velocity tracking, and [45] illustrates successful end-to-end control for precise jumping, the creation of a policy that effectively handles both diverse tasks and precise movements remains elusive.
-Furthermore, [51] introduces a foot constraint policy framework, enabling precise target tracking and accurate touchdown locations.
-Yet, there is still no framework that comprehensively addresses the dual demands of versatility and precision in locomotion.
-For BFMs, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks.
-In contrast, MLAMs struggle to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
-The difficulty in simultaneously achieving precise control and a broad range of actions in bipedal locomotion using DRL stems from several factors:
-  • Complex dynamics: Bipedal locomotion involves intricate dynamics, posing a significant challenge to maintaining both dynamic motion and precision.
-  • Resource intensity: Executing diverse locomotion tasks requires considerable computational power and extensive data, necessitating high-quality hardware and efficient DRL algorithms.
-  • Training conflicts: Training DRL systems to achieve both precision and versatility often leads to conflicts.
-Designing reward functions and training policies that satisfy both criteria is inherently complex.
-These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between BFM’s large-scale adaptability and MLAM’s hierarchical precision.
-IV-C2 Streamlined frameworks capable of handling complex tasks
-One of the critical challenges in unifying BFMs and MLAMs is designing frameworks that are both simple and capable of handling complex tasks.
-While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments.
-On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
-For BFMs, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity.
-For MLAMs, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency.
-Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
-The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure.
-To achieve an optimal framework, advancements in several key areas are essential:
-  • Robust and efficient DRL algorithms: Development of algorithms that can manage high-dimensional and continuous control problems more effectively.
-  • Specialised neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [52].
-  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviours and strategic outcomes.
-  • Advanced computational resources: Enhancement of computational capabilities to support more intensive training and faster inference, facilitating real-time decision-making in dynamic environments.
-By focusing on these developmental areas, the potential to create a unified, efficient, and less complex framework for handling complex locomotion challenges in bipedal robots is significantly increased.
-Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks.
-The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion.
-To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities.
-These advancements will lead to the creation of two potential unified models: (1) a Multi-Layered Adaptive Model grounded in the hierarchical framework and (2) a Bipedal Foundation Model built upon the end-to-end framework.
-The unified framework will ultimately support diverse locomotion tasks in complex environments.
-V Future Directions
-Section II and III introduced two mainstream learning-based frameworks, HFM and MLAM, and analysed their contributions toward achieving a unified framework.While Section IV provides a clearer understanding of the current state and the existing gap between these frameworks and the expected unified framework, further discussion is needed to outline specific directions for achieving this goal.
-In this section, we outline the essential components and future research directions necessary to bridge this gap.
-Specifically, we identify key elements that will drive the evolution from current framework limitations to a fully integrated unified framework.
-A summary of these fundamental elements is provided in Fig.
-4.
-V-A Future directions for bipedal robots
-V-A1 Loco-manipulation tasks
-With the rapid development of bipedal robotics, modern bipedal robots are now equipped with upper-body structures to enhance their capabilities beyond basic locomotion tasks.
-These advancements enable robots to perform upper-body-assisted locomotion and loco-manipulation tasks, expanding their adaptability in complex environments.
-Building on this evolution, upper-body-assisted locomotion enhances movement stability by integrating upper-lower body coordination, enabling robots to navigate challenging terrain more effectively.
-Beyond movement, manipulation tasks focus on precise upper-body control, requiring dexterous motion planning for object interaction and task execution.
-At the highest level of complexity, loco-manipulation integrates both locomotion and manipulation, demanding multi-contact interactions, dynamic balance maintenance, and real-time adaptation to unpredictable environments.
-These tasks necessitate full-body coordination and task-driven motion planning, making them essential benchmarks for advancing unified frameworks in humanoid robotics.
-Tasks include fundamental locomotion movements like flat walking and stepping over obstacles, progressing to upper-body-assisted locomotion challenges such as climbing ladders, wall-assisted walking, jumping over obstacles, and grasping for support.
-Beyond locomotion, precise upper-body control is essential for manipulation tasks like turning valves, driving vehicles, using tools, and connecting fire hoses.
-The most complex scenarios demand full-body coordination for loco-manipulation tasks such as carrying objects while walking, opening doors while navigating, and executing disaster-response operations.
-These benchmarks are critical for developing robust and adaptable humanoid systems, addressing key challenges posed in DARPA’s robotics initiatives [39].
-Therefore, loco-manipulation presents opportunities for humanoid robots to excel beyond purely bipedal capabilities.
-Few studies have addressed this integrated task; one such study [77] demonstrated a ’box transportation’ framework.
-This framework decomposes the task into five distinct policies, each addressing different aspects of the transportation process.
-However, this approach lacks efficiency and does not incorporate vision-based information, suggesting substantial room for improvement.
-Moreover, the challenges of managing mobile tools like scooters [78] or dynamically interacting with objects such as balls [79] introduce further complexities.
-To address these challenges, a hierarchical control approach could provide a more structured and adaptable solution.
-Decomposing loco-manipulation into multiple layers would simplify task execution by allowing precise, modular control over individual components [53].
-This method enables manual fine-tuning of task-specific controllers, ensuring better stability and adaptability to environmental variations.
-Additionally, a structured hierarchical approach facilitates coordinated whole-body responses, improving the efficiency of loco-manipulation interactions.
-Alternatively, an end-to-end learning framework could offer a more generalised and scalable solution.
-By leveraging task randomisation and structured curriculum learning, bipedal robots could progressively learn loco-manipulation skills in a unified policy [48, 36, 31].
-During training, such policies can also learn human-like movements from motion capture data [80, 27, 25], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
-V-A2 Bridging the gap from simulation to reality
-Due to the large number of interactions needed for DRL algorithms, training directly on robots can lead to costly damage to hardware and the environment.
-Consequently, training a policy in the simulation which offer a safe and cost-effective environment for developing robotics policies, and then deploying it to the hardware illustrates significant potential and efficiency.
-Bridging this gap is a key step toward achieving a unified framework, as it allows for the implementation of more versatile and robust systems that can handle diverse tasks across different terrains.
-Furthermore, the ultimate goal of achieving zero-shot transfer, where a policy trained in simulation can be deployed without additional real-world fine-tuning, directly contributes to the unified framework’s adaptability and scalability.
-This capability would make the framework far more practical for handling the wide range of locomotion tasks expected from future bipedal robots.
-Creating a reliable simulation environment and conducting RL training is a complex task.
-Advanced simulators are essential for training policies using full-order dynamics, which better represent the complex interactions robots face in the real world.
-Simulators like Isaac Gym [81], RoboCup3D [82], OpenAI Gym [83], and MuJoCo [84] are widely used to train policies in environments that closely mimic physical conditions.
-These tools help minimise the sim-to-real gap, improving the success of policy transfer to real robots.
-Further details on these simulators can be found in Appendix B.
-However, the process of sim-to-real encounters significant challenges due to the approximations and simplifications made in simulations.
-The process of sim-to-real faces significant challenges due to the approximations and simplifications made in simulations.
-Two primary issues arise: (1) Inaccurate dynamic models, contribute to discrepancies between simulation environments and real-world dynamics.
-They often simplify or generalise complex physical interactions, leading to a gap in performance when transitioning to reality.
-(2) The real environment is more diverse and complex, with varying terrain, sensor noise, and unpredictable conditions not captured in simulations.
-Numerous sim-to-real frameworks [43, 85, 86] have shown high efficiency and performance.
-Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments.
-To bridge the sim-to-real gap, several approaches have been developed, including dynamics randomisation [44], system identification [87, 88], periodic reward composition [49], learned actuator dynamics [87, 89], regulation feedback controllers [90], and adversarial motion priors [27, 26].
-In domain randomisation, there are two main methods for training policies: (1) end-to-end training using robot measurement histories, and (2) policy distillation, where an expert policy (e.g., teacher-student [49]) guides a student policy that learns from internal sensory feedback, as seen in techniques like RMA [85].
-Future research in bridging the sim-to-real gap will focus on enhancing the transferability and robustness of policies across diverse environments, ultimately aiming for zero-shot transfer, where no real-world fine-tuning is required.
-This will not only involve improving hardware but also making simulation environments more reflective of real-world conditions.
-Additionally, advancing sophisticated vision-based perception and multimodal sensory integration will be crucial in further minimising the sim-to-real gap, making policies more adaptable to complex and unpredictable environments.
-V-A3 Integrating foundation models
-Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots, as discussed in Section IV.
-Their integration into robotics extends beyond decision-making, offering new opportunities to enhance control, adaptability, and learning efficiency.
-One particularly promising direction is their use in RL, not merely for planning but for shaping the learning process itself—most notably in the design of reward functions, which remains a core challenge in bipedal locomotion tasks.
-The design of effective reward functions is a fundamental challenge in DRL for bipedal robots.
-While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem.
-These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming.
-Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
-LLMs can enhance RL by guiding policy optimisation and refining reward functions.
-A key challenge in DRL is reward function design, which typically requires manual tuning.
-LLMs offer a solution by dynamically adjusting reward signals based on linguistic descriptions of task success.
-For example, Kumar et al.
-[91] demonstrate how LLMs refine reward functions by incorporating linguistic feedback, improving hierarchical RL frameworks.
-The language model processes HL instructions, which are then translated into lower-level RL tasks, simplifying complex skill acquisition and human-robot interactions.
-Conversely, end-to-end approaches such as AnyBipe [92] use LLMs to autonomously adjust reward functions and control strategies, reducing manual intervention.
-This enables more autonomous learning, allowing the model to self-optimise in diverse environments and handle complex terrain navigation.
-Other implementations of foundation models in RL have been comprehensively reviewed in [60].
-Future research should further explore the synergy between LLMs and DRL, with a particular focus on adaptive reward shaping and self-optimising control policies.
-Advancing these methodologies could pave the way for more autonomous, self-learning humanoid robots, capable of dynamically refining their decision-making processes and adapting seamlessly to diverse environments with minimal human intervention.
-V-A4 Vision-based learning framework
-The vision-based learning framework plays a crucial role in advancing the unified framework for bipedal robots.
-Integrating vision into the unified system enhances the robot’s ability to perceive and interpret its environment, making it capable of handling more complex and dynamic tasks, such as navigating rough terrains or interacting with objects in real-time.
-Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [46, 53].
-This can be attributed to several factors: (1) Simplified Problem: Blind locomotion reduces task complexity, focusing directly on motor control strategies rather than processing visual inputs.
-(2) Computation and Training Complexity: Integrating vision requires significant computational resources and robust algorithms for real-time processing, terrain identification, and object detection.
-(3) Advanced Control Without Vision: Previous studies have already demonstrated robust blind locomotion using proprioceptive data, such as joint angles, IMU, with vision typically enhancing precision in locomotion [52] or path planning [57].
-However, recent advancements in computing power and computer vision technology—such as more efficient real-time visual processing algorithms and increasingly powerful hardware—have significantly accelerated the integration of vision-based learning into robotics.
-This shift has led to a surge in the popularity of vision-based approaches, allowing robots to more effectively interpret and adapt to complex environments, making them a central component in modern frameworks.
-Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal robots [29, 52, 36].
-Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
-  • Height scanner mapping: This approach, evaluated in works like [52], involves using height maps generated by scanners to inform locomotion strategies.
-These maps provide detailed topographical data, allowing robots to plan steps on uneven or obstructed surfaces more effectively.
-  • Direct vision inputs: Directly utilising inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [57, 93].
-Although previous studies like [57] have integrated visual navigation by feeding visual information to a HL planner, the potential of direct visual inputs to RL policies has not been fully explored.
-Enhancing the capability of bipedal robots to directly interpret and utilise visual data without intermediary processing can revolutionise their adaptability and efficiency in real-world scenarios.
-The exploration of direct vision inputs to RL-based policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
-V-B Insights from quadruped robots
+IV Limitations and Challenges
+The end-to-end and hierarchical frameworks detailed in Sections II and  III represent the state-of-the-art in DRL-based bipedal locomotion, demonstrating significant capabilities on specific tasks.
+However, a significant gap persists between these specialised successes and the realisation of a truly unified framework as defined in Section  I.
+Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these criteria.
+The path to a unified framework is hindered by several interconnected issues.
+At a foundational level, a primary difficulty involves the limitations and challenges in achieving both generalisation and precision (Section IV-A).
+This is further complicated by the practical barrier of the ‘reality gap’ in transferring policies from simulation to physical robots (Section IV-B).
+Ultimately, these issues culminate in the critical challenges of ensuring safety and interpretability for robust deployment in real-world, safety-critical situations (Section IV-C).
+IV-A Generalisation and precision
+A central challenge in applying DRL to bipedal locomotion is the need to simultaneously achieve high generalisation across diverse skills and transverse to all kinds of terrains, and high precision in specific tasks.
+This remains a fundamental obstacle to realising truly unified and capable frameworks.
+This capability gap is evident in the current literature.
+Many approaches excel at generalisation, demonstrating policies that enable versatile skills like walking and jumping [44, 35] and can transfer to different terrains [49, 33, 63].
+However, these generalized policies often lack the fidelity required for high-precision tasks, such as exact foot placement [38, 33, 54] or maintaining a specific velocity with minimal error [52].
+Conversely, controllers specialized for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [48], yet they cannot generalize these capabilities to a broader range of tasks.
+Thus, the development of a single, unified framework that concurrently exhibits both broad competency and high fidelity remains a largely unresolved challenge.
+This difficulty in uniting generalisation and precision is not arbitrary but stems from several key limitations inherent in current DRL paradigms, whether related to framework design, task formulation, or the training process itself:
+  • Limited terrain and gait patterns: The failure to generalise is often a direct result of training on insufficiently diverse environments or with a restricted set of behaviours.
+Models trained on limited terrain are brittle when faced with novel surfaces, while a limited gait pattern library prevents them from adapting to tasks requiring new motor skills.
+  • Poor command tracking: The learning signals for generalisation and precision are often in direct conflict.
+generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations.
+In contrast, precision demands restrictive signals that minimize command tracking error.
+Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [34, 35].
+  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [40, 64, 65].
+This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only provides support for the success of the final task, thus the task, hence the community’s frequent resort to reward shaping and curriculum learning [53, 32, 58, 66].
+Consequently, the immense amount of data required for an agent to explore, stumble upon a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation just to make training tractable [67, 68, 25, 26].
+  • High-quality data scarcity: As highlighted in our discussion in Section LABEL:subsec:guided_learning, the scarcity of high-quality demonstrations is a key bottleneck.Such data provides essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while bypassing extensive, unsafe exploration [26, 25, 69].
+This scarcity stems from the fact that scalable human data is difficult to transfer due to embodiment mismatch [70, 71, 59], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [20, 21].
+These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly embodied in the designs of the field’s dominant control architectures.
+End-to-end frameworks attempt a holistic solution, learning a single, monolithic policy that must implicitly resolve all challenges simultaneously.
+While this approach can yield highly versatile and dynamic behaviours [63], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within the system’s unstable dynamics.
+This often results in a notable lack of the fidelity and precision that hierarchical systems can enforce [35].
+Conversely, hierarchical frameworks, for example, are a direct architectural response to the lack of skill compositionality.
+By employing a “divide and conquer” strategy, they use a HL policy to sequence a library of LL, often model-based, controllers.
+This structure is designed to enforce precision and manage the complex dynamics at a lower level [21, 54, 39].
+However, as noted, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [38].
+IV-B Reality gap: From simulation to the physical world
+Another challenge hindering the deployment of DRL policies on bipedal robots is the simulation-to-reality (sim-to-real) gap.
+This refers to the significant discrepancy between a policy’s performance in a physics simulator and its performance on actual hardware.
+This gap is a critical obstacle because training directly on physical robots is often impractical.
+The millions of environmental interactions required for DRL would lead to accelerated mechanical wear, risk of catastrophic failure, and require constant human supervision.
+While simulation offers a safe and efficient alternative, the ultimate goal of “zero-shot” transfer, where a policy works perfectly without any real-world fine-tuning, is rarely achieved.
+A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [25, 26, 72, 73, 34].
+Even when transfer is successful, it often comes with compromises.
+Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [74, 46].
+In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [49, 33, 75].
+This gap is caused by unavoidable differences between the virtual and physical worlds, which are especially problematic for dynamically unstable bipedal robots.
+  • Robot dynamics modelling and actuation: Simulators struggle to perfectly replicate the complex dynamics of a physical bipedal robot, whose inherent instability makes it particularly sensitive to modeling errors.
+Factors such as motor friction, gear backlash, and precise link inertia are often simplified.
+  • Contact and terrain modelling: Accurately simulating the intermittent foot-ground contact that bipeds rely on is extremely difficult.
+A mismatch between simulated and real-world friction or surface properties can cause unexpected slips or bounces, leading to a loss of balance.
+  • Sensing and state estimation: A simulated robot has access to perfect, noise-free state information.
+In the real world, these states must be estimated from noisy sensors like IMUs and joint encoders [46, 74].
+For a bipedal robot, precise state estimation is critical for maintaining balance.
+Simulators like Isaac Gym [67], RoboCup3D [76], OpenAI Gym [77], and MuJoCo [78] detailed in Appendix B are widely used to train policies that closely mimic real-world physical conditions.
+These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [46, 79, 80] have demonstrated efficient and high-performance results.
+Despite these advancements, a significant gap persists between simulation and reality, exacerbated by the approximations made in simulation and the unpredictability of physical environments.
+IV-C Safety-critical Locomotion
+Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [1, 15, 16].
+This includes ensuring the robot’s own integrity to prevent costly damage, as well as guaranteeing the safety of the surrounding environment and any humans within it.
+While many existing frameworks have demonstrated impressive locomotion skills, they often prioritize performance over these safety considerations.
+This creates a critical barrier that separates success in controlled laboratory settings from reliable operation in the unpredictable real world.
+  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) like joint angles and IMU data [47, 49, 32], creating a major safety risk.
+Lacking external perception, these robots cannot anticipate obstacles, slopes, or slippery surfaces, making them purely reactive and highly prone to failure.
+Despite these significant safety drawbacks, this approach is often adopted for several key reasons.
+Omitting vision simplifies the control problem to pure motor skills and avoids the computational complexity of real-time visual processing.
+Furthermore, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [55] or path planning [60], rather than a core requirement for basic stability.
+  • Lack of physical constraint satisfaction: Many DRL frameworks lack built-in mechanisms to guarantee physical constraint satisfaction.
+This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules, for example Safe-RL on humanoids [81], hybrid DRL with identified low-dimensional safety models [82], footstep-constrained DRL policies [54], and reactive DRL steppers operating under feasibility constraints on uneven terrain [72].
+This limitation makes it difficult to prevent the robot from exceeding joint limits, applying excessive torques, or causing self-collisions, particularly when reacting to unexpected events.
+This is a key area where constrained RL could be applied.
+In summary, the pursuit of performance in DRL has often sidelined critical safety issues.
+The prevalence of blind policies that cannot anticipate environmental hazards, combined with the lack of inherent mechanisms to enforce physical constraints, creates a significant risk that hinders real-world deployment.
+While these challenges are significant, they also define a clear path forward.
+The following section on Future Directions and Opportunities will explore specific research avenues, such as vision-based learning and safe reinforcement learning, aimed directly at overcoming these safety barriers and enabling the development of truly robust and reliable bipedal robots.
+Figure 4: Towards a Unified Framework: This figure illustrates the logical progression from current DRL frameworks to future unified systems.
+It identifies the current limitations of existing end-to-end and hierarchical approaches, which motivate the exploration of specific Future Pathways.
+These pathways inform the design of two proposed conceptual models (i) Multi-Layered Adaptive Model (MLAM) and (ii) Bipedal Foundation Model (BFM) which represent potential blueprints for achieving a generalist, unified framework.
+V Future Directions and Opportunities
+Following the analysis of the surveyed frameworks and their limitations, this section outlines a path forward for DRL-based bipedal locomotion by exploring both direct research avenues and emerging opportunities.
+We begin in Section V-A by detailing research directions that directly respond to the challenges identified in Section IV.
+Building on this foundation, Section V-B broadens the scope to explore synergistic opportunities from related fields, such as loco-manipulation and the application of foundation models.
+These discussions culminate in Section V-C, where we propose two conceptual models for a unified framework that represent the future evolution of the end-to-end and hierarchical paradigms.
+V-A Pathways for bipedal locomotion
+V-A1 Multi-skill learning
+A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [63, 83, 35].
+This research direction focuses on enabling robots to acquire, adapt, and deploy a broad and varied repertoire of motor skills, allowing them to handle unforeseen situations and operate effectively in unstructured environments.
+To achieve such versatility, researchers are pursuing several HL pathways, which can be broadly categorized into structured and holistic approaches.
+The structured approach focuses on explicit decomposition.
+A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, a HL planner generates reference trajectories, while a LL DRL controller executes them robustly [59, 56], has shown in Section III.
+Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [26].
+A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single compact generalist policy [84].
+V-A2 Traversing challenging environments
+The goal of versatile skill learning is to enable bipedal robots to traverse challenging, human-centric environments where their unique form has an advantage.
+Validating capabilities on such terrains serves a crucial dual purpose.
+It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [49, 52, 63].
+More critically, it benchmarks precision on treacherous paths such as stepping stones that demand exact foot placement [33, 19, 72].
+These environments are the ultimate test of both a robot’s skill repertoire and its control fidelity.
+V-A3 Efficient learning and reward shaping
+As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [40, 32].
+Addressing this bottleneck is a crucial research frontier that calls for both more efficient algorithms and more robust reward designs.
+To mitigate sample inefficiency for complex skills, several research pathways are being actively explored.
+A primary strategy is to leverage prior data rather than learning entirely from scratch.
+Leveraging prior knowledge provides strong guidance and reduce unsafe exploration by anchoring policies to feasible motion patterns [26, 45, 59].
+Curriculum learning further organizes training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [53, 51, 39].
+Complementing advances in algorithms is the design of effective and robust rewards.
+Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [68, 26].
+Phase aware objectives are well established for cyclic gaits such as walking [32], whereas reward design for nonperiodic skills such as jumping is less standardized and often task specific [48].
+Promising directions reduce manual effort by adding higher level guidance, including event based terms, goal conditioned objectives, and kinematic reference tracking [54, 28, 50].
+Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand crafted objectives with implicit ones inferred from demonstrations [85].
+Together, these directions seek to minimize skill specific tuning and improve the transferability and reliability of learned locomotion policies.
+V-A4 Motion retargeting
+As human-like agents, bipedal robots, especially humanoids, have the unique advantage of a morphology that is similar to our own.
+This presents a significant opportunity: the potential to learn from vast libraries of human motion data.
+While large-scale datasets like AMASS [86] and Motion-X [87] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [29].
+Therefore, Motion Retargeting emerges as a critical component to bridge this gap.
+The challenge of this pathway is not merely to transfer human movements to the robot, but to generate trajectories that are both high-quality in their stylistic fidelity and rigorously physically feasible, adhering to the robot’s unique dynamics and constraints.
+Successfully developing these retargeting methods provides a scalable solution for accessing the data needed to train the natural and versatile generalist policies of the future.
+V-A5 Bridging the gap from simulation to reality
+Strategies to bridge the sim-to-real gap generally follow two main philosophies.
+The first aims to train policies that are robust enough to tolerate the inevitable mismatch between simulation and reality.
+The second focuses on minimizing the gap itself by making the simulator a more faithful replica of the physical world.
+The first approach seeks to reduce the discrepancy by improving the simulation’s fidelity.
+This is often achieved through system identification (SI), where real-world robot data is used to fine-tune the simulator’s parameters to create a more accurate ”digital twin” [74, 88].
+This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [74, 89].
+Other methods, such as designing specialized feedback controllers [90], also contribute by making the system less sensitive to the remaining modeling errors.
+In contrast, the second philosophy accepts that simulations will always be imperfect and instead focuses on creating highly adaptive, robust policies.
+The primary method here is DR, which forces a policy to generalize by training it across a wide range of simulated physical variations.
+Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [79], or via policy distillation, where a privileged “teacher” guides a “student” policy [52] to have a knowledge of unknown information like friction .
+Additionally, techniques like adversarial motion priors [31, 30] are used to ensure the learned behaviours are not just robust but also physically plausible.
+Looking ahead, the ultimate goal remains achieving reliable zero-shot transfer, where no real-world fine-tuning is needed.
+Progress will depend on the co-development of higher-fidelity simulations, improved hardware, and more robust control policies inherently capable of handling real-world unpredictability.
+The synergy of these advancements will be crucial in finally closing the sim-to-real gap.
+V-A6 Perception-conditioned locomotion
+ntegrating exteroceptive sensors like cameras and LiDAR enables bipedal robots to proactively plan footsteps, avoid obstacles, and adapt to upcoming terrain.
+This shift from reactive to anticipatory control is essential for navigating unstructured real-world environments.
+The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [33, 55, 39].
+In contrast, LiDAR is an active sensing method that generates precise 3D point clouds of the terrain.
+While vision provides richer data but is sensitive to lighting, LiDAR offers robust geometric measurements without visual detail.
+Based on this sensory data, current research is exploring two primary pathways for processing perceptual information for control.
+The first involves creating an intermediate geometric representation, such as a height map from scanners [55].
+This provides the policy with structured topographical data for effective footstep planning.
+The second is a more end-to-end approach, which utilizes direct vision inputs like RGB or depth images as inputs to the RL policy for real-time decision-making [60, 91].
+The former offers interpretability, while the latter promises more nuanced, reactive behaviours learned directly from raw perception.
+Future progress requires advancing both pathways: building richer, semantic world representations and improving the efficiency of direct perception-to-action policies.
+Solving the underlying challenges of real-time processing and the perceptual sim-to-real gap will be crucial for enabling truly adaptive locomotion in complex, real-world scenarios.
+V-A7 Constrained learning
+While the previously discussed pathways focus on enhancing a robot’s capabilities, a parallel and equally critical frontier is ensuring that these capabilities are exercised safely and reliably.
+To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [81, 82].
+Soft constraints encourage desirable behaviour and penalize undesirable behaviour without forbidding it.
+They are well suited for preferences or efficiency goals, for example minimizing energy use, limiting peak torques, or promoting smooth motion [81].
+Hard constraints are inviolable rules that prevent catastrophic failures.
+They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [54, 90].
+A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [92].
+In practice, a robust and trustworthy bipedal robot will likely combine both ideas.
+Soft constraints help a policy learn efficient and natural gaits, while hard constraints guarantee that it will not take catastrophic actions.
+This combination supports the transition from systems that are merely capable in laboratory settings to agents that are reliable, predictable, and safe for real-world deployment.
+V-B Opportunities
+V-B1 Leveraging foundation models for locomotion learning
+The recent rise of Foundation Models (FMs), such as Large Language Models (LLMs) and Vision-Language Models (VLMs), presents a transformative opportunity for bipedal locomotion.
+Their powerful reasoning capabilities are unlocking new approaches that go beyond traditional control methods, primarily by enabling sophisticated HL task planning and by providing novel solutions to shape the learning process itself, particularly in automated reward design.
+As HL planners, FMs provide a reasoning engine that can bridge the gap between abstract human goals and LL motor execution.
+They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for a LL policy to follow.
+This has been demonstrated effectively in legged robotics where VLMs process raw sensory data to pass structured commands to motor controllers [93], creating a seamless link between strategic planning and physical action.
+Furthermore, FMs create a significant opportunity to overcome one of the most persistent bottlenecks in DRL: reward design.
+Instead of tedious manual tuning, LLMs can dynamically generate or refine reward functions based on linguistic descriptions of task success.
+Research has shown that LLMs can translate human feedback into reward adjustments [94] or even autonomously adjust rewards and control strategies to self-optimize for diverse terrains [95], drastically reducing human intervention.
+The foremost opportunity lies in the deeper synergy between these roles.
+The integration of the HL symbolic reasoning of FMs with the LL, real-time control of DRL could create a new class of highly adaptive and flexible robots.
+As this rapidly evolving field progresses, as reviewed in [96], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
+V-B2 Loco-manipulation tasks
+While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it can’t physically interact with its environment.
+The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction.
+Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [42].
+However, realizing this opportunity is a significant challenge.
+Early studies, such as a ‘box transportation’ framework [97], often rely on inefficient, multi-policy solutions that lack visual perception.
+Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [98, 99].
+These difficulties create significant research opportunities.
+One such opportunity lies in exploring hierarchical control approaches.
+By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [56].
+Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution.
+Using techniques like curriculum learning and imitation from human motion capture data [51, 39, 35, 100, 31, 29], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
+V-B3 Insights from quadruped robots
 While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems.
 The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [101, 102], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [103].
-Furthermore, dynamic quadruped models facilitate highly adaptable policies [97], and sophisticated acrobatic motions are achieved through imitation learning [98].
+Furthermore, dynamic quadruped models facilitate highly adaptable policies [104, 105], and sophisticated acrobatic motions are achieved through IL [106].
 The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems.
 To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [107] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [62, 108] within a feedback DRL control hybrid scheme.
-The latter has shown substantial efficacy; it employs a Model Predictive Control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy.
-Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [76] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
+The latter has shown substantial efficacy; it employs a MPC to generate reference motions, which are then followed by a LL feedback DRL policy.
+Additionally, the Terrain-aware Motion Generation for Legged Robots module [109] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
 However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
 Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway.
-These include mimicking real animals through motion capture data and imitation learning [101, 102], as well as augmenting quadrupeds with manipulation abilities.
+These include mimicking real animals through motion capture data and IL [110, 111], as well as augmenting quadrupeds with manipulation abilities.
 This is achieved either by adding a manipulator [112, 113] or by using the robots’ legs [114].
@@
 Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
+V-C Conceptual Models for Unified Frameworks
+Motivated by our survey and the current state of the art, we propose two conceptual models, intended as reference designs, toward a unified locomotion framework.
+They build on end-to-end and hierarchical paradigms and offer complementary routes to scalable, generalizable architectures, rather than fully realized systems.
+  • Bipedal Foundation Models (BFMs): large-scale, pre-trained models that map perception directly to action through representation learning.
+Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
+  • Multi-Layer Adaptive Frameworks (MLAMs): modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs.
+MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
+In the following sections, we will analyze each of these conceptual models in detail, evaluating their respective strengths and challenges in the pursuit of a unified framework.
+V-C1 Bipedal Foundation Models
+Inspired by Robot Foundation Models (RFMs) [96, 115], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion.
+A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments.
+Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots.
+Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [116].
+The proposed BFM paradigm would operate in two stages.
+First, IL on diverse datasets would establish the generalisable foundation.
+Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot.
+The potential of this approach is highlighted by recent works, with frameworks like FLaRe [117] enhancing generalisation for long-horizon tasks, MOTO [118] enabling effective offline-to-online adaptation from images, and AdA [119] demonstrating in-context adaptation to novel environments.
+Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
+However, realizing the BFM concept for bipeds presents significant challenges.
+The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [117, 119].
+Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
+V-C2 Multi-layered Adaptive Models
+As a complementary path to BFMs, we propose the concept of MLAMs.
+Rather than relying on large-scale pre-training, this conceptual framework would adopt a modular, hierarchical approach.
+The idea is to extend conventional hierarchical frameworks (discussed in Section III) with explicitly adaptive layers, allowing for the dynamic composition of specialized policies.
+The core principle of this concept would be modularity, enabling each layer to be independently optimized or replaced and providing interpretable outputs at each stage.
+A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier.
+Each layer processes context-specific inputs and outputs interpretable commands.
+The HL reasoning layer leverages large pre-trained model such as LLMs and VLMs [120, 121] to parse commands into sub-tasks.
+For instance, Vision-Language Model Predictive Control [93] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
+By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
+The mid-level planner selects or synthesises motions via learned motion libraries [58, 44] or DRL-based planners [122].
+The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands.
+These include locomotion primitives like walking and climbing [123], adaptive tracking controllers for whole-body tracking [59], and imitation-based skills such as kicking and dancing [26], by utilizing RL, IL, or model-based methods.
+This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [123].
+However, realizing the MLAM concept would introduce challenges distinct from BFMs.
+Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [109].
+Additionally, the computational latency incurred by HL reasoning modules like LLMs [123] would pose limitations for tasks needing rapid reactions.
 VI Conclusion
@@
 In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies.
-Together, these approaches form the basis for two potential unified frameworks: the Bipedal Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
-To bridge the existing gaps and achieve a unified framework, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising.
-Such frameworks could simultaneously address issues of precision and generalisation.
-Moreover, the advent of LLMs presents a transformative opportunity, potentially enabling the unification of language processing and visual functionalities within robotic systems.
-While numerous challenges remain—ranging from the technical intricacies of framework integration to real-world application—the steady progression in control framework refinement and DRL development provides a hopeful outlook.
-The vision of achieving such unified framework may soon move within reach.
+While DRL has enabled remarkable progress, our survey concludes that current frameworks confront key limitations, including the tension between multi-skill generalisation and task-specific precision, the persistent sim-to-real gap, and critical safety concerns.
+To address these challenges, this survey synthesizes specific pathways for future research and identifies key opportunities for cross-pollination from related fields, such as FMs, loco-manipulation, and quadrupedal robotics.
+These insights culminate in our proposal of two conceptual frameworks: the BFMs, extending the end-to-end paradigm, and the MLAMs, evolving from the hierarchical approach, which together offer distinct blueprints for the next generation of generalist bipedal controllers.
 Appendix A Deep reinforcement learning algorithms

### v6

@@
 Abstract
-Bipedal robots are gaining global recognition due to their potential applications and advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL).
+Bipedal robots are gaining global recognition due to their potential applications and the rapid advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL).
 While DRL has significantly advanced bipedal locomotion, the development of a unified framework capable of handling a wide range of tasks remains an ongoing challenge.
 This survey systematically categorises, compares, and analyses existing DRL frameworks for bipedal locomotion, organising them into end-to-end and hierarchical control schemes.
-End-to-end frameworks are evaluated based on their learning approaches, while hierarchical frameworks are examined in terms of layered structures that integrate learning-based or traditional model-based methods.
+End-to-end frameworks are evaluated based on their learning approaches, whereas hierarchical frameworks are examined in terms of their layered structures that integrate learning-based and traditional model-based methods.
 We provide a detailed evaluation of the composition, strengths, limitations, and capabilities of each framework.
-Additionally, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with wide-ranging applications in real-world environments.
+Furthermore, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with broad applicability in real-world environments.
 Index Terms:
@@
 I Introduction
-Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using just two legs.
-Bipedal robots, which closely mimic human form, have a distinct advantage over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains.
-Furthermore, bipedal humanoid robots are specifically designed to function in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use.
+Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using only two legs.
+Bipedal robots, which closely mimic the human form, possess distinct advantages over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains.
+Furthermore, bipedal humanoid robots are specifically designed to operate in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use.
 This makes them highly adaptable to a wide range of tasks in such settings.
-As a result, bipedal robots hold significant potential for real-world applications [1].
-In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [2, 3, 4].
-Their agility is particularly advantageous in complex environments, such as multi-level workplaces.
-Bipedal robots are also well-suited for tasks that involve the use of human-designed tools, making them valuable for assisting in daily tasks, healthcare, and rehabilitation [5].
-Furthermore, they show considerable promise in search and rescue operations, where they can navigate hazardous and unpredictable terrains [6, 7, 8].
-Figure 1: Common bipedal and humanoid robots used as platforms for testing DRL frameworks.
-(a) NAO, a toy-like 3D humanoid robot, actuated by servo motors [9].
-(b) Rabbit, a 2D bipedal robot, actuated by torque control [10].
-(c) Cassie, a 3D bipedal robot, also actuated by torque control [11].
-(d) ATLAS, a 3D humanoid robot, driven by hydraulics [12].
-(e) Digit, a full human-sized 3D humanoid robot, an upgrade based on Cassie and actuated by torque control [13].
-Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [14, 15, 16].
-Early methods, such as the Linear Inverted Pendulum Model (LIPM) [17], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control.
+Figure 1: Representative bipedal and humanoid robots illustrating the diversity of platforms for locomotion research and development.
+(a) Cassie: a torque-controlled bipedal robot designed for agile locomotion.
+(b) Digit: a full-sized humanoid robot evolved from Cassie and actuated by torque control.
+(c) H1: a full-size, electric, torque-controlled humanoid robot developed by Unitree Robotics.
+(d) G1: a compact humanoid robot from Unitree featuring lightweight design and high joint backdrivability.
+(e) Atlas: a fully electric humanoid robot developed by Boston Dynamics.
+As a result, bipedal robots hold significant potential for real-world applications [2024_tong_advancements_review].
+In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [industrial_robotics_applications, collaborative_industrial, 2023_vidoeo_demo_digit_locomaniputlation].
+Their agility is particularly advantageous in complex environments such as multi-level workplaces.
+Bipedal robots are also well suited to tasks that involve the use of human-designed tools, making them valuable for assisting in daily activities, healthcare, and rehabilitation [2019_bingjing_human_exoskeleton].
+Moreover, they show considerable promise in search-and-rescue operations, where they can navigate hazardous and unpredictable terrains [2015_underwaterrobot_review, quadSpace1, quadSpace2].
+Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot].
+Early methods, such as the Linear Inverted Pendulum Model (LIPM) [2013_3DSLIPmodel_running], provided simplified representations of the dynamics involved in bipedal motion, enabling easier analysis and control.
 As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion.
-Advanced methods such as Model Predictive Control (MPC) [18, 19] and Trajectory Optimisation (TO) [20, 21, 14] exploit a predefined dynamics model to solve constrained optimal-control problems planning footsteps, CoM motion, and contact forces.
-While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
-Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [22], which offers a distinct advantage.
-Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments.
-Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
-Despite these advancements, a key question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks?
-If not, how should such a framework be designed, and what features or components should it include?
-To provide a clear benchmark for answering this question, we first define a unified framework as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions.
-The ultimate aim is to develop genuinely generalist locomotion policies that achieve robust zero-shot generalisation to previously unseen scenarios, embodiments, and objectives.
-With this definition as our guide, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical.
-These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
+Advanced methods such as Model Predictive Control (MPC) [2022_Hou_drlasparametertoMPC, 2023_junheng_steppingstones_MPCWBC] and Trajectory Optimisation (TO) [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2017_biped_model-based_method_review] exploit predefined dynamic models to solve constrained optimal-control problems that plan footsteps, centre-of-mass (CoM) motion, and contact forces.
+While model-based approaches offer rapid convergence and predictive capabilities, they often struggle in dynamically complex and uncertain environments where adaptability is essential.
+Reinforcement learning (RL)-based methods, particularly deep reinforcement learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [2020_systematic_review], which provides a distinct advantage.
+Unlike model-based approaches, which rely on predefined dynamics and may fail under unforeseen conditions, DRL enables robots to autonomously discover control strategies through trial and error, achieving greater adaptability and robustness in diverse environments.
+In addition, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both paradigms.
+Despite these advancements, research in DRL-based locomotion remains highly fragmented, with inconsistencies in training pipelines, reward formulations, observation spaces, and evaluation setups that hinder systematic benchmarking and slow progress towards generalisable locomotion capabilities.
+Moreover, many methods are tailored to specific morphologies or tasks, offering limited transferability across embodiments and environments.
+This fragmentation motivates the following central research questions: To what extent has current research achieved generalisation and robustness across diverse morphologies, terrains, and locomotion tasks?
+If full generalisation has not yet been realised, how can existing DRL approaches be organised and extended towards a unified framework that enables such capability in bipedal robots?
+In this context, the present survey seeks to categorise emerging DRL frameworks for bipedal locomotion, identify their key limitations, and outline opportunities for integration and convergence towards unification.
+To address these aspects, we first clarify the ultimate goal of robot learning: to develop systems that exhibit generalisation, adaptability, and robustness across diverse morphologies, tasks, and environments.
+The unified framework is therefore not the final destination but a conceptual scaffold emerging from the consolidation of current DRL research efforts.
+Its role is to organise fragmented methodologies through shared interfaces, training conventions, and evaluation protocols, thereby promoting steady progress towards the broader goal of generalisable and adaptive robot-learning systems.
+Guided by these definitions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical.
 End-to-end frameworks directly map robot states to joint-level control outputs, while hierarchical frameworks decompose decision-making into multiple layers.
@@
 The task-level decision-making tier interfaces directly with user commands or predefined tasks, forming a structured approach to robotic control.
-The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications.
-Although the application of RL to simple 2D bipedal robots began in 2004 [23, 24], it took several years before DRL algorithms emerged.
-These DRL-based methods have since shown promising results in physical simulators [25, 26, 27].
-Agility Robotics introduced the first sim-to-real end-to-end learning framework in 2019, which was applied on the 3D torque-controlled bipedal robot Cassie, as shown in Fig.
-1(c) [11].
-Besides model-based reference learning, the policy can also incorporate motion capture data [28, 29, 30, 31], or start from scratch [32] to explore solutions freely.
-Recent studies demonstrate that end-to-end frameworks robustly handle complex and diverse tasks [33, 34, 35].
+The evolution of RL in bipedal robotics has largely advanced through the end-to-end learning paradigm.
+Early studies in 2004 applied simple policy-gradient methods to 2D bipeds [2004_Tedrake_RL_bipedal_policygradient, 2004_model-based_RL_learnpoincaremap_simple_bipedrobot], while later breakthroughs in DRL enabled policy training in high-fidelity physics simulators [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_wenhao_yu_DRL_withoutpredefine_symmetrygait_PPO_jointanglePD_].
+As robotic hardware matured, an increasing variety of bipedal and humanoid platforms emerged, supporting extensive evaluation of DRL-based locomotion across diverse morphologies, as illustrated in Fig.
+1.
+This evolution marked the transition from purely simulation-based training to simulation-to-real (sim-to-real) transfer, where policies trained in simulators are deployed on physical robots.
+In 2020, the first successful sim-to-real transfer of an end-to-end DRL locomotion policy was achieved on the 3D torque-controlled bipedal robot Cassie [2020_Xie_firstsim2real_].
+Subsequent work explored two principal learning paradigms: reference-based learning, which leverages TO-generated data or motion-capture data to guide policy training [2021_michael_velocity-basedcontrol_motioncapturedata, 2024_speratebody_expressive_allmotion, 2024_tang_humanmimic_adversarial, 2024_Zhang_wholebody_adversarial_motion_priors]; and reference-free learning, where policies are trained entirely from scratch to autonomously discover control strategies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie].
+These developments demonstrate that end-to-end frameworks can achieve robust and versatile locomotion skills across complex terrains and dynamic environments [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
 Similarly, hierarchical structures have garnered significant interest.
 Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies.
-A notable framework employs a learned HL planner coupled with a LL model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [21, 36, 37].
-Another innovative construction integrates a learned feedback controller with an HL planner, categorised under the DRL Feedback Control Hybrid Scheme [38, 39].
-Additionally, a learned hierarchical control scheme [40] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [25, 26, 41].
-In order to provide a clearer understanding of the current landscape, we have categorised existing DRL frameworks, as depicted in Fig.
+Hybrid architectures often integrate learning-based and model-based modules to combine adaptability with physical consistency.
+One representative design couples a learned HL planner with an LL model-based controller, forming a cascade-structure or deep-planning hybrid scheme [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller, 2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN].
+Alternatively, DRL feedback-control hybrids embed learned control policies within model-based feedback loops to enhance tracking precision and disturbance rejection [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
+Learned hierarchical control schemes [2017_DRL_survey] decompose locomotion into multiple layers, each focusing on specific functions such as navigation and fundamental locomotion skills [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
+To provide a clearer overview of the current landscape, we categorise existing DRL frameworks as shown in Fig.
 2.
 Figure 2: Classification of DRL-based control schemes.
-The approaches are broadly categorised into two main paradigms: End-to-end frameworks, which learn a single policy from sensory inputs to motor commands, and Hierarchical frameworks, which decompose the control problem into multiple levels.
-Within the end-to-end paradigm, a key distinction is between Reference-free learning (learning from scratch) and Reference-based learning (tracking a predefined motion).
+The approaches are broadly categorised into two main paradigms: end-to-end frameworks, which learn a single policy from sensory inputs to motor commands; and hierarchical frameworks, which decompose the control problem into multiple levels.
+Within the end-to-end paradigm, a key distinction is drawn between reference-free learning (learning from scratch) and reference-based learning (tracking a predefined motion).
 Hierarchical structures include hybrid control schemes, which synergistically combine learned components with traditional model-based controllers.
-Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure.
-Although bipedal locomotion tasks have evolved significantly in the past decade.
-Earlier frameworks primarily focused on basic walking stabilisation and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation.
-Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [42], emphasizing real-world deployment and robustness.
-Although several reviews discuss RL for general robotics [22] and model-based methods for bipedal robots [14, 15, 16], none specifically focus on DRL-based frameworks for bipeds.
-To address this gap, this survey reviews relevant literature by applying the following selection criteria: (1) studies that examine DRL frameworks specifically for bipedal robots, (2) research involving both simulated and physical bipedal robots, (3) approaches that improve policy transfer from simulation to real-world environments, and (4) publications from the last five years (2018 - April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, ArXiv, and major robotics conferences such as CoRL, RSS, ICRA, IROS, and Humanoids.
+Current progress across both end-to-end and hierarchical paradigms indicates that a unified framework for DRL-based bipedal locomotion is still far from being realised.
+Establishing such a framework is essential for consolidating diverse learning pipelines, standardising evaluation metrics, and enabling transferable locomotion capabilities across different robot morphologies.
+As locomotion tasks become increasingly complex, ranging from basic stabilisation to dynamic parkour and loco-manipulation, the need for consistent benchmarking has intensified.The DARPA Robotics Challenge exemplified this trend by introducing one of the first large-scale evaluation platforms for bipedal humanoids performing real-world locomotion and manipulation tasks [2015_DAPRA_Challenge_1], highlighting the importance of robustness and practical deployment.
+Although several reviews discuss RL for general robotics [2020_systematic_review] and model-based methods for bipedal robots [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot], none specifically focus on DRL-based frameworks for bipeds.
+To address this gap, this survey reviews relevant literature according to the following selection criteria: (1) studies that investigate DRL frameworks specifically for bipedal robots; (2) research involving both simulated and physical bipedal robots; (3) approaches that improve policy transfer from simulation to real-world environments; and (4) publications from the last five years (2018–April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, arXiv, and major robotics venues such as CoRL, RSS, ICRA, IROS, and Humanoids.
 The search was conducted using the key terms “deep reinforcement learning” or “reinforcement learning” in combination with “bipedal locomotion”, “bipedal walking”, “biped robot”, “humanoid robot”, or “legged robot”.
 The most relevant and impactful works were manually selected for further review.
-This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where necessary.
-For clarity, in this survey, the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
+This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where appropriate.
+For clarity, throughout this paper the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
 The primary contributions of this survey are:
   • A comprehensive summary and cataloguing of DRL-based frameworks for bipedal locomotion.
-  • A detailed comparison of each control scheme, highlighting their strengths, limitations, and characteristics.
+  • A detailed comparison of each control scheme, highlighting their strengths, limitations, and distinctive characteristics.
   • The identification of current challenges and the provision of insightful future research directions.
-The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches.
-Section III details hierarchical frameworks, classified into three main types.
-Section IV presents specific limitations and challenges, highlighting their connection to the previously surveyed content.
-Section V addresses potential future pathways, opportunities, and two proposed conceptual models that extend the end-to-end and hierarchical paradigms.
-Finally, Section VI concludes the survey.
+The survey is organised as follows: Section II discusses end-to-end frameworks, categorised by learning approaches; Section III presents hierarchical frameworks, classified into three main types; Section IV outlines key limitations and challenges, linking them to the preceding discussions; Section V explores potential pathways, opportunities, and two proposed conceptual models that extend the end-to-end and hierarchical paradigms; finally, Section VI concludes the survey.
 II End-to-end framework
-The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, lidar data, or proprioceptive feedback [43], along with user commands [32] or pre-defined references [44], into joint-level control actions $\mathcal{U}$.
+The end-to-end DRL framework represents a holistic approach in which a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, LiDAR data, or proprioceptive feedback [2017_compare_actionspace_RL], together with user commands [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] or pre-defined references [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], into joint-level control actions $\mathcal{U}$.
 Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function.
 This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
-End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a Proportional-Derivative (PD) controller.
-These methods can be broadly categorised based on their reliance on prior knowledge into two types: reference-based and reference-free.
+End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a proportional–derivative (PD) controller.
+These methods can be broadly categorised according to their reliance on prior knowledge into two types: reference-based and reference-free.
 The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
-In the following sections, we will delve into various representation frameworks, exploring their characteristics, limitations, and strengths in comprehensive detail.
+The following sections delve into various representation frameworks, exploring their characteristics, limitations, and strengths in comprehensive detail.
 To facilitate an understanding of these distinctions, Table I provides a succinct overview of the frameworks discussed.
-TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework.
-The dashed line in the implementation flow chat refers to optional.00footnotetext: Forward Walk involves the bipeds walking straight ahead.
-Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities.
-Omni-Walk grants the bipeds the ability to walk in any direction.
+TABLE I: Summary and comparison of reference-based and reference-free learning approaches for the end-to-end framework.
+The dashed line in the implementation flow chart indicates optional steps.00footnotetext: Forward Walk involves bipeds walking straight ahead.
+Unidirectional Walk enables bipeds to move forward and backward within a range of desired velocities.
+Omnidirectional Walk grants bipeds the ability to walk in any direction.
 Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities.
 Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping.
-Versatile Jump refers to jump towards different desired targets.
+Versatile Jump refers to jumping towards different desired targets.
 Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
@@
 Residual learning
-[45]
-[46]
+[2018_xie_OSUDRL_feedback_DRL_reference-based]
+[2020_Xie_firstsim2real_]
 Forward walk
-undirectional walk
-Add residual term to the known motor positions at the current time step.
-A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains.
+unidirectional walk
+Adding a residual term to the known motor positions at the current time step.
+A: Fast convergence speed D: Requires high-quality predefined reference, limits to specific motions, and lacks robustness to complex terrains.
 Guided learning
-[47]
-[44]
-[48]
-[35]
+[2020_Siekmann_drl_e2e_s2r_cassie]
+[2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter]
+[2023_UCB_Cassie_RL_Jumping]
+[2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining]
 Forward walk
@@
 Versatile motions
-Mimic the predefined reference and directly specifies joint-level command.
-A: Accelerate the learning process and robust to the terrains. D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment.
+Mimic the predefined reference and directly specifies joint-level commands.
+A: Accelerates the learning process and is robust across terrains. D: Limits to the predefined motions and lacks adaptability to unforeseen changes in environment.
 Reference-free learning
-[32]
-[49]
-[33]
+[2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie]
+[2021_siekmann_blind_DRL_stair]
+[2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark]
 Periodic motions
@@
 Learn locomotion skills from scratch without any prior knowledge.
-A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources.
+A: High potential for gait exploration, highly robust to complex terrain D: Requires intensive reward shaping for gait patterns and is relatively expensive in computational resources
 II-A Reference-based learning
-Reference-based learning leverages prior knowledge generated offline through methods such as trajectory optimisation (TO) or motion capture systems.
+Reference-based learning leverages prior knowledge generated offline through methods such as TO or motion capture systems.
 This predefined reference typically includes data related to the robot’s joint movements or pre-planned trajectories, serving as a foundation for the policy to develop locomotion skills by following these established motion patterns.
@@
 This holistic approach enables the biped robot to adapt to various dynamic conditions while maintaining balance and control.
-Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [45].
+Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [2018_xie_OSUDRL_feedback_DRL_reference-based].
 This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A.
@@
 These modified references are then processed by a PD controller to set the desired joint positions.
-While this framework enhanced the robot’s ability to perform tasks beyond standing [50], its physical deployment on a bipedal robot has not yet occurred, potentially rendering it impractical for managing walking at varying speeds and limiting movement to a single direction.
-To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied on a physical bipedal robot [46].
+Although this framework has improved the robot’s ability to perform tasks beyond standing [2018_yang_push_recover_wholebody], it has yet to be physically deployed on a bipedal robot.
+As a result, it remains impractical for managing walking at varying speeds and is constrained to movement in a single direction.
+To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied to a physical bipedal robot [2020_Xie_firstsim2real_].
 Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks.
 However, the robot’s movements remain constrained to merely walking forward or backward.
-A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [51].
+A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [2021_Diego_deepwalk_curriculum_learning_sim2real].
 In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive.
@@
 Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance.
-Compared to other learning approaches that directly output joint positions, it is highly sample-efficient [36].
+Compared to other learning approaches that directly output joint positions, it is highly sample efficient [2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller].
 However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
@@
 Guided learning trains policies to directly output the desired joint-level commands as actions $a$, without relying on the addition of a residual term.
-The states space is as same as residual learning approach.
+The state space is the same as the residual-learning approach.
 In this approach, the reward structure is centred on accurately imitating predefined reference trajectories, ensuring precise alignment between the policy output and the reference motion.
-A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [47].
+A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [2020_Siekmann_drl_e2e_s2r_cassie].
 In this framework, the action space directly maps to the joint angles, and desired joint positions are managed by joint PD controllers.
-The framework also incorporates a Long Short-Term Memory (LSTM) network, as detailed in Appendix A, which is synchronised with periodic time inputs.
+The framework also incorporates a Long Short-Term Memory (LSTM) network, as detailed in the Appendix A, which is synchronised with periodic time inputs.
 However, this model is limited to a single locomotion goal: forward walking.
-A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [44], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
+A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
 Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances.
-To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [48].
-This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping maneuvers.
-An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [31].
+To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [2023_UCB_Cassie_RL_Jumping].
+This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping manoeuvrers.
+An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [2024_Zhang_wholebody_adversarial_motion_priors].
 This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
-While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [35] based on the foundational work in [48].
+While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining] based on the foundational work in [2023_UCB_Cassie_RL_Jumping].
 This framework enhances the learning process by incorporating a wide range of locomotion skills and introducing a dual I/O history approach, marking a significant breakthrough in creating a robust, versatile, and dynamic end-to-end framework.
@@
 Guided learning methods expedite the learning process by leveraging expert knowledge and demonstrating the capacity to achieve versatile and robust locomotion skills.
-Through the comprehensive evaluation [35], it is demonstrated that guided learning employs references without complete dependence on them.
+Through the comprehensive evaluation [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], it is demonstrated that guided learning employs references without complete dependence on them.
 Conversely, residual learning exhibits failures or severe deviations when predicated on references of inferior quality.
 This shortfall stems from the framework’s dependency on adhering closely to the provided references, which narrows its learning capabilities.
-Nonetheless, reference-based learning reliance on predefined trajectories confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [44, 52].
-Additionally, this approach exhibits limited adaptability in responding effectively to unforeseen environmental changes or novel challenges.
+However, the benefits of reference-based learning come with inherent limitations.
+Reliance on predefined trajectories often confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction].
+Moreover, such methods exhibit reduced adaptability when confronted with novel environments or unforeseen perturbations.
+These limitations are further compounded by the difficulty of acquiring high-quality and task-relevant demonstrations.
+Common sources of prior knowledge include TO [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], human motion capture [2024_speratebody_expressive_allmotion], teleoperation [2023_seo_deep_loco_manipulation, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and scripted controllers [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+While informative, these demonstrations often require adaptation due to embodiment mismatch or limited generalisability.
+Motion retargeting [2019_iit_wholebody_retargeting, 2024_tang_humanmimic_adversarial, 2017_Ayusawa_motionretarget_optimization], as one of promising direction, addresses this by converting human-centric motions into robot-feasible trajectories, yet it still struggles with preserving fidelity and adapting across morphologies.
+Ultimately, the success of guided learning relies not only on using references but on accessing high-quality, adaptable demonstrations that generalise across tasks and platforms—highlighting a key challenge in advancing robust policy learning.
 II-B Reference-free learning
@@
 The action space and observation space in this approach are similar to the guided-learning method; however, the reward structure differs significantly from the reference-based method.
-Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [52].
+Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction].
 The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models.
-A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [27].
+A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [2018_wenhao_yu_DRL_withoutpredefine_symmetrygait_PPO_jointanglePD_].
 This framework introduced a novel term into the loss function and utilised a curriculum learning strategy to effectively shape gait patterns.
-Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model, Cassie, though this has yet to be validated outside of simulation [53].
-Considering the practical implementation of theoretical models, significant efforts have been directed towards developing sim-to-real frameworks in robotics studies.
-A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [32].
+Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model (Cassie), though this has yet to be validated outside of simulation [2020_zhaoming_drl_steppingstones_PPOwithactorcritic_referencefree_simulation].
+Considering the practical implementation of this approach, significant efforts have been made to develop sim-to-real reference-free frameworks, and their potential has been further explored on physical robots.
+A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie].
 This framework employs periodic rewards to facilitate initial training within simulations before successfully transitioning to a physical robot.
 It has been further refined to adapt to diverse terrains and scenarios.
-For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [49].
-Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [54], thus enabling the robot to effectively navigate stepping stones [33].
-Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [55].
+For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [2021_siekmann_blind_DRL_stair].
+Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation], thus enabling the robot to effectively navigate stepping stones [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark].
+Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [2023_duan_OSUDRL_heightmap_visionbased_hybrid].
 This approach to learning enables the exploration of novel solutions and strategies that might not be achievable through mere imitation of existing behaviours.
@@
 III Hierarchy framework
+Figure 3: Hierarchical control scheme diagram.
+This figure illustrates a hierarchical control framework for a bipedal robot, comprising a basic scheme and three variations.
+(1) Basic scheme: The framework begins with a task command, followed by an HL planner and a LL controller, which ultimately drives the robot.
+Each module can be replaced with a learned policy, introducing adaptability across different control layers.
+(2) Variations (from left to right): (a) a deep planning hybrid scheme, in which the HL planner is learned; (b) a feedback DRL control hybrid scheme, with a learned $L$ controller; and (c) a learned hierarchical control scheme, where both layers are learned.
 Unlike end-to-end policies that directly map sensor inputs to motor outputs, hierarchical control schemes deconstruct locomotion challenges into discrete, manageable layers or stages of decision-making.
@@
 This division not only enhances the framework’s flexibility but also simplifies the problem-solving process for each policy.
-The architecture of a hierarchical framework typically comprises two principal modules: an High-Level (HL) planner and an Low-Level (LL) controller.
+The architecture of a hierarchical framework typically comprises two principal modules: an HL planner and an LL controller.
 This modular approach allows for the substitution of each component with either a model-based method or a learning-based policy, further enhancing adaptability and customisation to specific needs.
@@
 Each type offers unique capabilities and exhibits distinct characteristics, albeit with limitations primarily due to the complexities involved in integrating diverse modules and their interactions.
-For a concise overview, Table 3 summarises the various frameworks, detailing their respective strengths, limitations, and primary characteristics.
+For a concise overview, Table II summarises the various frameworks, detailing their respective strengths, limitations, and primary characteristics.
 The subsequent sections will delve deeper into each of these frameworks, providing a thorough analysis of their operational mechanics and their application in real-world scenarios.
-TABLE II: Summary and comparison of Hierarchy framework
+TABLE II: Summary and comparison of hierarchical frameworks.
 Control Scheme
@@
 Module
-Characteristic
+Characteristics
 Advantages and Disadvantages
 Deep Planning Hybrid Scheme
-[36]
-[56]
-[57]
+[2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller]
+[2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller]
+[2024_gaspard_footstepnet_soccer]
 Deep planning + ID
@@
 HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
-A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust.
-D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.
+A: Enhanced command tracking capabilities, generalised across different platforms, sampling-efficient, and robust.
+D: Complicated system and communication between layers, requires a precise model, and lacks generalisation across different tasks.
 Feedback DRL Control Hybrid Scheme
-[58]
-[38]
-[39]
+[2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real]
+[2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P]
+[2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]
 Gait library + feedback policy
@@
 LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
-A: Short inference times, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, reducing sampling efficiency.
+A: Short inference time, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, which reduces sampling efficiency.
 Learned Hierarchy Framework
-[25]
-[41]
-[59]
+[2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.]
+[2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot]
+[2024_stanford_humanplus_hierarchical_humanoidshadowingimitation]
 HL policy + LL policy
@@
 HL policy + LL policy
-Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills.
+Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learns navigation skills.
 A: Provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy.
-D: inefficienty sim-to-real, complicated interface between layers, training expensively.
+D: Inefficient sim-to-real transfer, complicated interface between layers, and an expensive training process.
 III-A Deep planning hybrid scheme
@@
 Several studies have demonstrated the integration of an HL planner policy with a model-based controller to achieve tasks in world space.
-A notable framework optimises task space level performance, eschewing direct joint level and balancing considerations [36].
+A notable framework optimises task space level performance, eschewing direct joint-level and balancing considerations [2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller].
 This system combines a residual learning planner with an inverse dynamics controller, enabling precise control over task-space commands to joint-level actions, thereby improving velocity tracking, foot touchdown location, and height control.
-Further advancements include a hybrid framework that merges HZD-based residual deep planning with model-based regulators to correct errors in learned trajectories, showcasing robustness, training efficiency, and effective velocity tracking [37].
+Further advancements include a hybrid framework that merges HZD-based residual deep planning with model-based regulators to correct errors in learned trajectories, showcasing robustness, training efficiency, and effective velocity tracking [2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN].
 These frameworks have been successfully transferred from simulation to reality and validated on robots such as Cassie.
 However, the limitations imposed by residual learning constrained the agents’ capacity to explore a broader array of possibilities.
-Building on previous work [37], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [56].
+Building on previous work [2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller].
 In this approach, a purely learning-based HL planner interacts with an LL controller using an Inverse Dynamics with Quadratic Programming formulation (ID-QP).
@@
 Demonstrating robustness and training efficiency, this framework has outperformed other models and was successfully generalised across various bipedal platforms, including Digit, Cassie, and RABBIT.
-In parallel, several research teams have focused on developing navigation planners specifically for toy-like humanoid robots, which provide greater physical stability compared to torque-driven or hydraulic bipedal robots as shown in Fig.
-1.
-One notable study [60] implemented a visual navigation policy on the NAO robot, depicted in Fig.
-1(a), utilising RGB cameras as the primary sensory modality.
-This system has demonstrated successful zero-shot transfer to real-world scenarios, enabling the robot to adeptly navigate around obstacles.
-Further research [57] has explored complex dynamic motion tasks, such as playing soccer, by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a centre of mass (CoM) trajectory.
-This configuration is coupled with a whole-body controller, facilitating dynamic activities like soccer shooting.
-Despite their platform’s stability, provided by large feet and a lightweight structure, these robots exhibit limited dynamic movement capabilities compared to full-sized humanoid robots.
-Consequently, this research primarily addresses navigation and task execution.
-Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [56] and sophisticated navigation capabilities [57].
+In parallel, several research teams have focused on developing navigation and locomotion planners for humanoid robots, leveraging onboard visual perception and learned control strategies.
+Recent work [2024_gaspard_footstepnet_soccer] explored complex dynamic motion tasks such as playing soccer by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a CoM trajectory.
+This configuration, coupled with a whole-body controller, enables dynamic activities like soccer shooting.
+Although these systems demonstrate promising coordination between perception, planning, and control, they remain limited in dynamic movement capability compared to full-sized humanoid robots, and thus primarily address navigation and task-level execution.
+Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller] and sophisticated navigation capabilities [2024_gaspard_footstepnet_soccer].
 However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping.
@@
 Enhancing both navigation and dynamic locomotion capabilities within the HL planner remains a significant challenge.
-Figure 3: Hierarchy Control Scheme Diagram: presents a hierarchical control framework for a bipedal robot, consisting of a basic scheme and three variations.
-(1) Basic Scheme: The framework starts with a Task Command, followed by a HL Planner, and a LL Controller, which ultimately drives the Robot.
-Each module can be replaced by a learned policy, introducing adaptability in different control layers.
-(2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned.
-(b) A Feedback DRL control hybrid scheme, with a learned $L$ controller.
-(c) A Learned hierarchy control scheme, where both layers are learned.
 III-B Feedback DRL control hybrid scheme
-In contrast to the comprehensive approach of end-to-end policies discussed in Section II, which excels in handling versatile locomotion skills and complex terrains with minimal interface times, the Feedback DRL Control Hybrid Scheme integrates DRL policies as LL controllers.
+In contrast to the comprehensive approach of end-to-end policies discussed in Section II, which excels in handling versatile locomotion skills and complex terrains with minimal inference time, the Feedback DRL Control Hybrid Scheme integrates DRL policies as LL controllers.
 These LL controllers, replacing traditional model-based feedback mechanisms, work in conjunction with HL planners that process terrain information, plan future walking paths, and maintain robust locomotion stability.
-For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [58].
+For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real].
 Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness.
 A more dynamic approach involves online planning, which has shown greater adaptability and efficiency.
-One notable framework combines a conventional foot planner with an LL DRL policy [38], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands.
+One notable framework combines a conventional foot planner with an LL DRL policy [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands.
 Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies.
 However, this work has not yet been transferred from simulation to real-world applications.
-Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [39].
+Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
 This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
 A recent development has shown that focusing solely on foot placement might restrict the stability and adaptability of locomotion, particularly in complex maneuvers.
-A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [61].
+A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [2024_Li_model_based_LLDRL_hybrid_control_scheme].
 This system employs a residual learning architecture, where the DRL policy’s outputs are merged with the planner’s directives before being relayed to the PD controller.
@@
 Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities.
-These policies have demonstrated high efficiency in locomotion and accurate reference tracking [62].
-Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited.
-Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [62].
-Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [39].
+These policies have demonstrated high efficiency in locomotion and accurate reference tracking [quadRLLL].
+Despite their extensive capabilities, such policies generally require short inference time, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited.
+Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [quadRLLL].
+Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
 This capability effectively addresses the navigational limitations observed in end-to-end frameworks.
@@
 Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models.
-Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation and hinder rapid deployment and iterative enhancements [62].
+Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation, and hinder rapid deployment and iterative enhancements [quadRLLL].
 III-C Learned hierarchy framework
@@
 Subsequently, an HL policy is developed to direct the robot towards specific targets, encapsulating a structured approach to robotic autonomy.
-The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [25].
+The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.].
 In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance.
@@
 This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal.
-The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [26].
-However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks have predominantly been validated only in simulations.
+The framework was later enhanced to include imitation learning (IL), facilitating the replication of dynamic human-like movements within the simulation environment [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks has been validated only in simulations.
 The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
-Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [41].
-To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [59].
-It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as a LL tracking controller.
+Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
+To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation].
+It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as an LL tracking controller.
 Additionally, binocular RGB vision input enhances perception, enabling precise loco-manipulation and dynamic locomotion tasks such as jumping, walking, folding clothes, and rearranging objects.
-This shadowing-based imitation learning approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
+This shadowing-based IL approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
 Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics.
-Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [25].
+Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.].
 As discussed in the previous section, the benefits of integrating RL-based planners with RL-based controllers have been effectively demonstrated.
 This combination enables the framework to adeptly manage a diverse array of environments and tasks.
-Within such a framework, HL policy is optimised for strategic planning and achieving specific goals.
+Within such a framework, the HL policy is optimised for strategic planning and achieving specific goals.
 This optimisation allows for targeted enhancements depending on the tasks at hand.
@@
 Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
-Additionally, the training process for each policy within the hierarchy demands considerable computational resources [41].
+Additionally, the training process for each policy within the hierarchy demands considerable computational resources [2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
 The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions.
@@
 IV Limitations and Challenges
-The end-to-end and hierarchical frameworks detailed in Sections II and  III represent the state-of-the-art in DRL-based bipedal locomotion, demonstrating significant capabilities on specific tasks.
-However, a significant gap persists between these specialised successes and the realisation of a truly unified framework as defined in Section  I.
-Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these criteria.
-The path to a unified framework is hindered by several interconnected issues.
+The end-to-end and hierarchical frameworks detailed in Sections II and III represent the state of the art in DRL-based bipedal locomotion, demonstrating remarkable capabilities on specific tasks.
+However, a substantial gap remains between these task-oriented successes and the broader goal of achieving generalisation and adaptability across diverse morphologies, tasks, and environments.
+Bridging this gap requires more than incremental improvements—it demands the establishment of a unified framework that consolidates interfaces, training conventions, and evaluation protocols to systematically address the underlying limitations of current DRL pipelines.
+As outlined in the following sections, the core challenges underlying this gap can be grouped into three interrelated aspects.
 At a foundational level, a primary difficulty involves the limitations and challenges in achieving both generalisation and precision (Section IV-A).
-This is further complicated by the practical barrier of the ‘reality gap’ in transferring policies from simulation to physical robots (Section IV-B).
+This is further complicated by the practical barrier of the sim-to-real gap in transferring policies from simulation to physical robots (Section IV-B).
 Ultimately, these issues culminate in the critical challenges of ensuring safety and interpretability for robust deployment in real-world, safety-critical situations (Section IV-C).
 IV-A Generalisation and precision
-A central challenge in applying DRL to bipedal locomotion is the need to simultaneously achieve high generalisation across diverse skills and transverse to all kinds of terrains, and high precision in specific tasks.
+A central challenge in applying DRL to bipedal locomotion is the need to simultaneously achieve high generalisation across diverse skills and traverse to all kinds of terrains, and high precision in specific tasks.
 This remains a fundamental obstacle to realising truly unified and capable frameworks.
 This capability gap is evident in the current literature.
-Many approaches excel at generalisation, demonstrating policies that enable versatile skills like walking and jumping [44, 35] and can transfer to different terrains [49, 33, 63].
-However, these generalized policies often lack the fidelity required for high-precision tasks, such as exact foot placement [38, 33, 54] or maintaining a specific velocity with minimal error [52].
-Conversely, controllers specialized for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [48], yet they cannot generalize these capabilities to a broader range of tasks.
-Thus, the development of a single, unified framework that concurrently exhibits both broad competency and high fidelity remains a largely unresolved challenge.
+Many approaches excel at generalisation, demonstrating policies that enable versatile skills such as walking and jumping [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining] and can transfer to different terrains [2021_siekmann_blind_DRL_stair, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_radosavovic_realworld_transformer_NNmodel].
+However, these generalised policies often lack the fidelity required for high-precision tasks such as exact foot placement [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation] or maintaining a specific velocity with minimal error [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction].
+Conversely, controllers specialised for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [2023_UCB_Cassie_RL_Jumping], yet they cannot generalise these capabilities to a broader range of tasks.
+Thus, the development of a single unified framework that concurrently exhibits both broad competency and high fidelity remains largely unresolved.
 This difficulty in uniting generalisation and precision is not arbitrary but stems from several key limitations inherent in current DRL paradigms, whether related to framework design, task formulation, or the training process itself:
   • Limited terrain and gait patterns: The failure to generalise is often a direct result of training on insufficiently diverse environments or with a restricted set of behaviours.
-Models trained on limited terrain are brittle when faced with novel surfaces, while a limited gait pattern library prevents them from adapting to tasks requiring new motor skills.
+Models trained on limited terrain are brittle when faced with novel surfaces, while a limited gait pattern library prevents adaptation to tasks requiring new motor skills.
   • Poor command tracking: The learning signals for generalisation and precision are often in direct conflict.
-generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations.
-In contrast, precision demands restrictive signals that minimize command tracking error.
-Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [34, 35].
-  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [40, 64, 65].
-This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only provides support for the success of the final task, thus the task, hence the community’s frequent resort to reward shaping and curriculum learning [53, 32, 58, 66].
-Consequently, the immense amount of data required for an agent to explore, stumble upon a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation just to make training tractable [67, 68, 25, 26].
-  • High-quality data scarcity: As highlighted in our discussion in Section LABEL:subsec:guided_learning, the scarcity of high-quality demonstrations is a key bottleneck.Such data provides essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while bypassing extensive, unsafe exploration [26, 25, 69].
-This scarcity stems from the fact that scalable human data is difficult to transfer due to embodiment mismatch [70, 71, 59], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [20, 21].
-These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly embodied in the designs of the field’s dominant control architectures.
-End-to-end frameworks attempt a holistic solution, learning a single, monolithic policy that must implicitly resolve all challenges simultaneously.
-While this approach can yield highly versatile and dynamic behaviours [63], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within the system’s unstable dynamics.
-This often results in a notable lack of the fidelity and precision that hierarchical systems can enforce [35].
-Conversely, hierarchical frameworks, for example, are a direct architectural response to the lack of skill compositionality.
-By employing a “divide and conquer” strategy, they use a HL policy to sequence a library of LL, often model-based, controllers.
-This structure is designed to enforce precision and manage the complex dynamics at a lower level [21, 54, 39].
-However, as noted, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [38].
-IV-B Reality gap: From simulation to the physical world
-Another challenge hindering the deployment of DRL policies on bipedal robots is the simulation-to-reality (sim-to-real) gap.
+Generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations, whereas precision demands restrictive signals that minimise command-tracking error.
+Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
+  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [2017_DRL_survey, 2017_Schulman_PPO_instructure, 2023_Omur_comparison_RLalgorithm_bipedallocomotion].
+This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only supports the success of the final task.
+Consequently, the immense amount of data required for an agent to explore, discover a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation merely to make training tractable [2021_Sim_Isaac, 2017_heess_emergence_animation_PPO_comprarisonrlalgorithm_differenttarrains, 2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+  • High-quality data scarcity: As highlighted in Section II-A2, the scarcity of high-quality demonstrations is a key bottleneck.
+Such data provide essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while avoiding unsafe exploration [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2020_motioncapture_referencebased_e2e_simulation_pybullet_augmentedrewarddesign].
+This scarcity stems from the difficulty of transferring scalable human data due to embodiment mismatch [2017_Ayusawa_motionretarget_optimization, 2019_iit_wholebody_retargeting, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy].
+These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly reflected in the design of the field’s dominant control architectures.
+End-to-end frameworks attempt a holistic solution, learning a single monolithic policy that must implicitly resolve all challenges simultaneously.
+While this approach can yield highly versatile and dynamic behaviours [2023_radosavovic_realworld_transformer_NNmodel], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within unstable system dynamics.
+This often results in a lack of the fidelity and precision that hierarchical systems can enforce [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
+Conversely, hierarchical frameworks are a direct architectural response to the lack of skill compositionality.
+By employing a “divide and conquer” strategy, they use an HL policy to sequence a library of LL, often model-based, controllers.
+This structure enforces precision and manages complex dynamics at a lower level [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
+However, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P].
+IV-B Challenges in transferring from simulation to reality
+Another challenge hindering the deployment of DRL policies on bipedal robots is the sim-to-real gap.
 This refers to the significant discrepancy between a policy’s performance in a physics simulator and its performance on actual hardware.
 This gap is a critical obstacle because training directly on physical robots is often impractical.
-The millions of environmental interactions required for DRL would lead to accelerated mechanical wear, risk of catastrophic failure, and require constant human supervision.
+The millions of environmental interactions required for DRL would lead to accelerated mechanical wear, a risk of catastrophic failure, and require constant human supervision.
 While simulation offers a safe and efficient alternative, the ultimate goal of “zero-shot” transfer, where a policy works perfectly without any real-world fine-tuning, is rarely achieved.
-A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [25, 26, 72, 73, 34].
+A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2021_DeepQ_DNQ_stepping_animationlonly_steppingstones, 2017_Josh_drl_animation_imitationlearning_GRAIL_TRPO_MUJOCO, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly].
 Even when transfer is successful, it often comes with compromises.
-Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [74, 46].
-In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [49, 33, 75].
+Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [2019_Yu_Phy_system_identification, 2020_Xie_firstsim2real_].
+In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [2021_siekmann_blind_DRL_stair, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2020_park_recoverstability_drl_PPO_referencebased_simulation].
 This gap is caused by unavoidable differences between the virtual and physical worlds, which are especially problematic for dynamically unstable bipedal robots.
-  • Robot dynamics modelling and actuation: Simulators struggle to perfectly replicate the complex dynamics of a physical bipedal robot, whose inherent instability makes it particularly sensitive to modeling errors.
+  • Robot dynamics modelling and actuation: Simulators struggle to replicate the complex dynamics of a physical bipedal robot, whose inherent instability makes it particularly sensitive to modelling errors.
 Factors such as motor friction, gear backlash, and precise link inertia are often simplified.
-  • Contact and terrain modelling: Accurately simulating the intermittent foot-ground contact that bipeds rely on is extremely difficult.
-A mismatch between simulated and real-world friction or surface properties can cause unexpected slips or bounces, leading to a loss of balance.
+  • Contact and terrain modelling: Accurately simulating intermittent foot–ground contact is extremely difficult.
+A mismatch between simulated and real-world friction or surface properties can cause unexpected slips or bounces, leading to loss of balance.
   • Sensing and state estimation: A simulated robot has access to perfect, noise-free state information.
-In the real world, these states must be estimated from noisy sensors like IMUs and joint encoders [46, 74].
+In the real world, these states must be estimated from noisy sensors such as IMUs and joint encoders [2020_Xie_firstsim2real_, 2019_Yu_Phy_system_identification].
 For a bipedal robot, precise state estimation is critical for maintaining balance.
-Simulators like Isaac Gym [67], RoboCup3D [76], OpenAI Gym [77], and MuJoCo [78] detailed in Appendix B are widely used to train policies that closely mimic real-world physical conditions.
-These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [46, 79, 80] have demonstrated efficient and high-performance results.
+Simulators such as Isaac Gym [2021_Sim_Isaac], RoboCup3D [2003_Sim_Robocup], OpenAI Gym [2016_Sim_OpenAI], and MuJoCo [2012_Sim_Mujoco], detailed in Appendix B, are widely used to train policies that closely mimic real-world physical conditions.
+These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [2020_Xie_firstsim2real_, 2022_Ashish_RMA_sim2real_UCB, 2023_Rohan_sim2real_currentdecompose] have demonstrated efficient and high-performance results.
 Despite these advancements, a significant gap persists between simulation and reality, exacerbated by the approximations made in simulation and the unpredictability of physical environments.
 IV-C Safety-critical locomotion
-Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [1, 15, 16].
+Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [2024_tong_advancements_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot].
 This includes ensuring the robot’s own integrity to prevent costly damage, as well as guaranteeing the safety of the surrounding environment and any humans within it.
-While many existing frameworks have demonstrated impressive locomotion skills, they often prioritize performance over these safety considerations.
+While many existing frameworks have demonstrated impressive locomotion skills, they often prioritise performance over these safety considerations.
 This creates a critical barrier that separates success in controlled laboratory settings from reliable operation in the unpredictable real world.
-  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) like joint angles and IMU data [47, 49, 32], creating a major safety risk.
+  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) such as joint angles and IMU data [2020_Siekmann_drl_e2e_s2r_cassie, 2021_siekmann_blind_DRL_stair, 2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], creating a major safety risk.
 Lacking external perception, these robots cannot anticipate obstacles, slopes, or slippery surfaces, making them purely reactive and highly prone to failure.
-Despite these significant safety drawbacks, this approach is often adopted for several key reasons.
-Omitting vision simplifies the control problem to pure motor skills and avoids the computational complexity of real-time visual processing.
-Furthermore, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [55] or path planning [60], rather than a core requirement for basic stability.
+Despite these significant safety drawbacks, this approach is often adopted for several reasons: omitting vision simplifies the control problem to pure motor skills and avoids the computational cost of real-time visual processing.
+Moreover, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [2023_duan_OSUDRL_heightmap_visionbased_hybrid] or path planning [2018_Tsunekawa_visualnavigation_hierarchy_DDPG-sim2real_NAO], rather than a core requirement for basic stability.
   • Lack of physical constraint satisfaction: Many DRL frameworks lack built-in mechanisms to guarantee physical constraint satisfaction.
-This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules, for example Safe-RL on humanoids [81], hybrid DRL with identified low-dimensional safety models [82], footstep-constrained DRL policies [54], and reactive DRL steppers operating under feasibility constraints on uneven terrain [72].
+This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules—for example, Safe-RL on humanoids [2020_Javier_safereinforcementlearning], hybrid DRL with identified low-dimensional safety models [2022_Li_UCB_safteystability_model-based_DRL_PPO], footstep-constrained DRL policies [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation], and reactive DRL steppers operating under feasibility constraints on uneven terrain [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones].
 This limitation makes it difficult to prevent the robot from exceeding joint limits, applying excessive torques, or causing self-collisions, particularly when reacting to unexpected events.
@@
 In summary, the pursuit of performance in DRL has often sidelined critical safety issues.
-The prevalence of blind policies that cannot anticipate environmental hazards, combined with the lack of inherent mechanisms to enforce physical constraints, creates a significant risk that hinders real-world deployment.
-While these challenges are significant, they also define a clear path forward.
-The following section on Future Directions and Opportunities will explore specific research avenues, such as vision-based learning and safe reinforcement learning, aimed directly at overcoming these safety barriers and enabling the development of truly robust and reliable bipedal robots.
+The prevalence of blind policies that cannot anticipate environmental hazards, combined with the lack of inherent mechanisms to enforce physical constraints, creates significant risk and hinders real-world deployment.
+While these challenges are considerable, they also define a clear path forward.
+The following section on Future Directions and Opportunities explores specific research avenues, such as vision-based learning and safe reinforcement learning, aimed at overcoming these safety barriers and enabling the development of truly robust and reliable bipedal robots.
 Figure 4: Towards a Unified Framework: This figure illustrates the logical progression from current DRL frameworks to future unified systems.
@@
 V-A Pathways for bipedal locomotion
+In relation to the research question introduced in Section I, progress in DRL-based bipedal locomotion should be assessed not only through conventional metrics such as reward and success rate but also by broader system-level measures.
+These include generalisation breadth (across skills, terrains, and morphologies), precision in fidelity-critical tasks (e.g., command-tracking error and foot-placement accuracy), safety and constraint compliance (joint, torque, and contact feasibility), and efficiency or deployability (sample efficiency and on-robot inference latency).
+These dimensions build directly upon the challenges outlined in Section IV and together define the key pathways for advancing bipedal locomotion.
+The following subsections elaborate on these pathways, each addressing one or more of the above aspects to guide progress towards more generalisable and robust control systems.
 V-A1 Multi-skill learning
-A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [63, 83, 35].
+A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [2023_radosavovic_realworld_transformer_NNmodel, 2024_radosavovic_humanoid_locomotiontokenprediction_transformer, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
 This research direction focuses on enabling robots to acquire, adapt, and deploy a broad and varied repertoire of motor skills, allowing them to handle unforeseen situations and operate effectively in unstructured environments.
-To achieve such versatility, researchers are pursuing several HL pathways, which can be broadly categorized into structured and holistic approaches.
+To achieve such versatility, researchers are pursuing several HL pathways, which can be broadly categorised into structured and holistic approaches.
 The structured approach focuses on explicit decomposition.
-A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, a HL planner generates reference trajectories, while a LL DRL controller executes them robustly [59, 56], has shown in Section III.
-Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [26].
-A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single compact generalist policy [84].
+A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, an HL planner generates reference trajectories, while an LL DRL controller executes them robustly [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation, 2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller], as shown in Section III.
+Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single, compact generalist policy [huang2024diffuseloco].
 V-A2 Traversing challenging environments
-The goal of versatile skill learning is to enable bipedal robots to traverse challenging, human-centric environments where their unique form has an advantage.
+The goal of versatile skill learning is to enable bipedal robots to traverse challenging, human-centric environments where their unique form offers an advantage.
 Validating capabilities on such terrains serves a crucial dual purpose.
-It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [49, 52, 63].
-More critically, it benchmarks precision on treacherous paths such as stepping stones that demand exact foot placement [33, 19, 72].
+It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [2021_siekmann_blind_DRL_stair, 2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction, 2023_radosavovic_realworld_transformer_NNmodel].
+More critically, it benchmarks precision on treacherous paths such as stepping stones, which demand exact foot placement [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_junheng_steppingstones_MPCWBC, 2021_DeepQ_DNQ_stepping_animationlonly_steppingstones].
 These environments are the ultimate test of both a robot’s skill repertoire and its control fidelity.
 V-A3 Efficient learning and reward shaping
-As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [40, 32].
+As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [2017_DRL_survey, 2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie].
 Addressing this bottleneck is a crucial research frontier that calls for both more efficient algorithms and more robust reward designs.
@@
 A primary strategy is to leverage prior data rather than learning entirely from scratch.
-Leveraging prior knowledge provides strong guidance and reduce unsafe exploration by anchoring policies to feasible motion patterns [26, 45, 59].
-Curriculum learning further organizes training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [53, 51, 39].
+Leveraging prior knowledge provides strong guidance and reduces unsafe exploration by anchoring policies to feasible motion patterns [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_xie_OSUDRL_feedback_DRL_reference-based, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation].
+Curriculum learning further organises training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [2020_zhaoming_drl_steppingstones_PPOwithactorcritic_referencefree_simulation, 2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
 Complementing advances in algorithms is the design of effective and robust rewards.
-Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [68, 26].
-Phase aware objectives are well established for cyclic gaits such as walking [32], whereas reward design for nonperiodic skills such as jumping is less standardized and often task specific [48].
-Promising directions reduce manual effort by adding higher level guidance, including event based terms, goal conditioned objectives, and kinematic reference tracking [54, 28, 50].
-Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand crafted objectives with implicit ones inferred from demonstrations [85].
-Together, these directions seek to minimize skill specific tuning and improve the transferability and reliability of learned locomotion policies.
+Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [2017_heess_emergence_animation_PPO_comprarisonrlalgorithm_differenttarrains, 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+Phase-aware objectives are well established for cyclic gaits such as walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], whereas reward design for non-periodic skills such as jumping is less standardised and often task specific [2023_UCB_Cassie_RL_Jumping].
+Promising directions reduce manual effort by adding higher-level guidance, including event-based terms, goal-conditioned objectives, and kinematic reference tracking [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2021_michael_velocity-basedcontrol_motioncapturedata, 2018_yang_push_recover_wholebody].
+Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand-crafted objectives with implicit ones inferred from demonstrations [2016_GAIL_imitation_learning].
+Together, these directions seek to minimise skill-specific tuning and improve the transferability and reliability of learned locomotion policies.
 V-A4 Motion retargeting
-As human-like agents, bipedal robots, especially humanoids, have the unique advantage of a morphology that is similar to our own.
+As human-like agents, bipedal robots—especially humanoids—have the unique advantage of a morphology that is similar to our own.
 This presents a significant opportunity: the potential to learn from vast libraries of human motion data.
-While large-scale datasets like AMASS [86] and Motion-X [87] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [29].
+While large-scale datasets such as AMASS [2019_Naureen_AMASS] and Motion-X [2024_motion_X_datasets] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [2024_speratebody_expressive_allmotion].
 Therefore, motion retargeting emerges as a critical component to bridge this gap.
-The challenge of this pathway is not merely to transfer human movements to the robot, but to generate trajectories that are both high-quality in their stylistic fidelity and rigorously physically feasible, adhering to the robot’s unique dynamics and constraints.
+The challenge of this pathway is not merely to transfer human movements to the robot, but to generate trajectories that are both high in stylistic fidelity and physically feasible, adhering to the robot’s unique dynamics and constraints.
 Successfully developing these retargeting methods provides a scalable solution for accessing the data needed to train the natural and versatile generalist policies of the future.
@@
 Strategies to bridge the sim-to-real gap generally follow two main philosophies.
-The first aims to train policies that are robust enough to tolerate the inevitable mismatch between simulation and reality.
-The second focuses on minimizing the gap itself by making the simulator a more faithful replica of the physical world.
+The first aims to train policies robust enough to tolerate the inevitable mismatch between simulation and reality, while the second focuses on minimising the gap itself by making the simulator a more faithful replica of the physical world.
 The first approach seeks to reduce the discrepancy by improving the simulation’s fidelity.
-This is often achieved through system identification (SI), where real-world robot data is used to fine-tune the simulator’s parameters to create a more accurate ”digital twin” [74, 88].
-This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [74, 89].
-Other methods, such as designing specialized feedback controllers [90], also contribute by making the system less sensitive to the remaining modeling errors.
+This is often achieved through system identification (SI), where real-world robot data are used to fine-tune the simulator’s parameters to create a more accurate “digital twin” [2019_Yu_Phy_system_identification, 2023_masuda_simtoreal_identification].
+This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [2019_Yu_Phy_system_identification, 2019_simtoreal_actuatornet_quadrupedrobot].
+Other methods, such as designing specialised feedback controllers [2021_Casillo_cyberbotics_hierarchyscheme_bezierpolynominals], also contribute by making the system less sensitive to residual modelling errors.
 In contrast, the second philosophy accepts that simulations will always be imperfect and instead focuses on creating highly adaptive, robust policies.
 The primary method here is DR, which forces a policy to generalize by training it across a wide range of simulated physical variations.
-Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [79], or via policy distillation, where a privileged “teacher” guides a “student” policy [52] to have a knowledge of unknown information like friction .
-Additionally, techniques like adversarial motion priors [31, 30] are used to ensure the learned behaviours are not just robust but also physically plausible.
+Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [2022_Ashish_RMA_sim2real_UCB], or via policy distillation, where a privileged “teacher” guides a “student” policy [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction] to have a knowledge of unknown information like friction .
+Additionally, techniques like adversarial motion priors [2024_Zhang_wholebody_adversarial_motion_priors, 2024_tang_humanmimic_adversarial] are used to ensure the learned behaviours are not just robust but also physically plausible.
 Looking ahead, the ultimate goal remains achieving reliable zero-shot transfer, where no real-world fine-tuning is needed.
@@
 V-A6 Perception-conditioned locomotion
-ntegrating exteroceptive sensors like cameras and LiDAR enables bipedal robots to proactively plan footsteps, avoid obstacles, and adapt to upcoming terrain.
+Integrating exteroceptive sensors such as cameras and LiDAR enables bipedal robots to proactively plan footsteps, avoid obstacles, and adapt to upcoming terrain.
 This shift from reactive to anticipatory control is essential for navigating unstructured real-world environments.
-The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [33, 55, 39].
+The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_duan_OSUDRL_heightmap_visionbased_hybrid, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
 In contrast, LiDAR is an active sensing method that generates precise 3D point clouds of the terrain.
@@
 Based on this sensory data, current research is exploring two primary pathways for processing perceptual information for control.
-The first involves creating an intermediate geometric representation, such as a height map from scanners [55].
+The first involves creating an intermediate geometric representation, such as a height map from scanners [2023_duan_OSUDRL_heightmap_visionbased_hybrid].
 This provides the policy with structured topographical data for effective footstep planning.
-The second is a more end-to-end approach, which utilizes direct vision inputs like RGB or depth images as inputs to the RL policy for real-time decision-making [60, 91].
+The second is a more end-to-end approach, which utilises direct vision inputs such as RGB or depth images as inputs to the RL policy for real-time decision-making [2018_Tsunekawa_visualnavigation_hierarchy_DDPG-sim2real_NAO, 2022_byravan_nerf2real_basketball].
 The former offers interpretability, while the latter promises more nuanced, reactive behaviours learned directly from raw perception.
@@
 While the previously discussed pathways focus on enhancing a robot’s capabilities, a parallel and equally critical frontier is ensuring that these capabilities are exercised safely and reliably.
-To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [81, 82].
-Soft constraints encourage desirable behaviour and penalize undesirable behaviour without forbidding it.
-They are well suited for preferences or efficiency goals, for example minimizing energy use, limiting peak torques, or promoting smooth motion [81].
+To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [2020_Javier_safereinforcementlearning, 2022_Li_UCB_safteystability_model-based_DRL_PPO].
+Soft constraints encourage desirable behaviour and penalise undesirable behaviour without forbidding it.
+They are well suited to preferences or efficiency goals, for example minimising energy use, limiting peak torques, or promoting smooth motion [2020_Javier_safereinforcementlearning].
 Hard constraints are inviolable rules that prevent catastrophic failures.
-They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [54, 90].
-A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [92].
+They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2021_Casillo_cyberbotics_hierarchyscheme_bezierpolynominals].
+A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [2016_Quan_gaitlibrary].
 In practice, a robust and trustworthy bipedal robot will likely combine both ideas.
@@
 V-B1 Leveraging foundation models for locomotion learning
-The recent rise of Foundation Models (FMs), such as Large Language Models (LLMs) and Vision-Language Models (VLMs), presents a transformative opportunity for bipedal locomotion.
+The recent rise of Foundation Models (FMs), such as Large Language Models (LLMs) and Vision–Language Models (VLMs), presents a transformative opportunity for bipedal locomotion.
 Their powerful reasoning capabilities are unlocking new approaches that go beyond traditional control methods, primarily by enabling sophisticated HL task planning and by providing novel solutions to shape the learning process itself, particularly in automated reward design.
 As HL planners, FMs provide a reasoning engine that can bridge the gap between abstract human goals and LL motor execution.
-They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for a LL policy to follow.
-This has been demonstrated effectively in legged robotics where VLMs process raw sensory data to pass structured commands to motor controllers [93], creating a seamless link between strategic planning and physical action.
+They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for an LL policy to follow.
+This has been demonstrated effectively in legged robotics, where VLMs process raw sensory data to pass structured commands to motor controllers [2024_VLM_PC], creating a seamless link between strategic planning and physical action.
 Furthermore, FMs create a significant opportunity to overcome one of the most persistent bottlenecks in DRL: reward design.
 Instead of tedious manual tuning, LLMs can dynamically generate or refine reward functions based on linguistic descriptions of task success.
-Research has shown that LLMs can translate human feedback into reward adjustments [94] or even autonomously adjust rewards and control strategies to self-optimize for diverse terrains [95], drastically reducing human intervention.
+Research has shown that LLMs can translate human feedback into reward adjustments [2023_kumar_wordscontrol_LLM] or even autonomously adjust rewards and control strategies to self-optimise for diverse terrains [2024_LLM_AnyBipe], drastically reducing human intervention.
 The foremost opportunity lies in the deeper synergy between these roles.
 The integration of the HL symbolic reasoning of FMs with the LL, real-time control of DRL could create a new class of highly adaptive and flexible robots.
-As this rapidly evolving field progresses, as reviewed in [96], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
+As this rapidly evolving field progresses, as reviewed in [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
 V-B2 Loco-manipulation tasks
-While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it can’t physically interact with its environment.
+While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it cannot physically interact with its environment.
 The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction.
-Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [42].
-However, realizing this opportunity is a significant challenge.
-Early studies, such as a ‘box transportation’ framework [97], often rely on inefficient, multi-policy solutions that lack visual perception.
-Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [98, 99].
+Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [2015_DAPRA_Challenge_1].
+However, realising this opportunity is a significant challenge.
+Early studies, such as a ‘box transportation’ framework [2023_dao_OSUDRL_loco-manipulation_hybrid_simtoreal], often rely on inefficient, multi-policy solutions that lack visual perception.
+Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [2023_baltes_control_scooter, 2023_deepmind_haarnoja_playsoccer].
 These difficulties create significant research opportunities.
 One such opportunity lies in exploring hierarchical control approaches.
-By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [56].
+By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller].
 Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution.
-Using techniques like curriculum learning and imitation from human motion capture data [51, 39, 35, 100, 31, 29], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
+Using techniques like curriculum learning and imitation from human motion-capture data [2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining, 2023_seo_deep_loco_manipulation, 2024_Zhang_wholebody_adversarial_motion_priors, 2024_speratebody_expressive_allmotion], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
 V-B3 Insights from quadruped robots
 While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems.
-The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [101, 102], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [103].
-Furthermore, dynamic quadruped models facilitate highly adaptable policies [104, 105], and sophisticated acrobatic motions are achieved through IL [106].
-The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems.
-To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [107] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [62, 108] within a feedback DRL control hybrid scheme.
-The latter has shown substantial efficacy; it employs a MPC to generate reference motions, which are then followed by a LL feedback DRL policy.
-Additionally, the Terrain-aware Motion Generation for Legged Robots module [109] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
+The diversity of frameworks developed for quadrupeds ranges from end-to-end, model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [quadModelBased1, quadModelBased2], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [quadDeformable].
+Furthermore, dynamic quadruped models facilitate highly adaptable policies [quadModular, humphreys2024bio], and sophisticated acrobatic motions are achieved through IL [quadImitation2].
+The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks that integrate vision-based systems.
+To date, two primary versions of such hierarchical frameworks have been developed: one where a deep-planning module is paired with model-based control [quadRLPlanning] within a deep-planning hybrid scheme, and another that combines model-based planning with LL DRL control [quadRLLL, quadmpcref] within a feedback DRL control hybrid scheme.
+The latter has shown substantial efficacy; it employs an MPC to generate reference motions, which are then followed by an LL feedback DRL policy.
+Additionally, the Terrain-aware Motion Generation for Legged Robots module [TAMOLS] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
 However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
 Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are under way.
-These include mimicking real animals through motion capture data and IL [110, 111], as well as augmenting quadrupeds with manipulation abilities.
-This is achieved either by adding a manipulator [112, 113] or by using the robots’ legs [114].
-Notably, the research presented in [113] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
-Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers and high mechanical complexity can be attributed to the cause of this disparity between quadruped and bipedal robots.
-Establishing a unified framework could bridge this gap, an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems.
+These include mimicking real animals through motion-capture data and IL [quadAnimal1, quadAnimal2], as well as augmenting quadrupeds with manipulation abilities.
+This is achieved either by adding a manipulator [quadManip1, quadManip2] or by using the robots’ legs [quadManip3].
+Notably, the research presented in [quadManip2] demonstrates that loco-manipulation tasks can be effectively managed using a single, unified, end-to-end framework.
+Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers, and high mechanical complexity can be attributed to this disparity between quadruped and bipedal robots.
+Establishing a unified framework could bridge this gap—an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems.
 Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
 V-C Conceptual models for unified frameworks
-Motivated by our survey and the current state of the art, we propose two conceptual models, intended as reference designs, toward a unified locomotion framework.
-They build on end-to-end and hierarchical paradigms and offer complementary routes to scalable, generalizable architectures, rather than fully realized systems.
+Motivated by our survey and the current state of the art, we propose two conceptual models, intended as reference designs, towards a unified locomotion framework.
+They build on end-to-end and hierarchical paradigms and offer complementary routes to scalable, generalisable architectures, rather than fully realised systems.
   • Bipedal Foundation Models (BFMs): large-scale, pre-trained models that map perception directly to action through representation learning.
 Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
-  • Multi-Layer Adaptive Frameworks (MLAMs): modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs.
-MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
-In the following sections, we will analyze each of these conceptual models in detail, evaluating their respective strengths and challenges in the pursuit of a unified framework.
+  • Multi-Layer Adaptive Models (MLAMs): modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs.
+MLAMs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
+In the following sections, we will analyse each of these conceptual models in detail, evaluating their respective strengths and challenges in the pursuit of a unified framework.
 V-C1 Bipedal foundation models
-Inspired by Robot Foundation Models (RFMs) [96, 115], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion.
+Inspired by Robot Foundation Models (RFMs) [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture, 2023_generalpurposerobot_FMs], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion.
 A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments.
 Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots.
-Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [116].
+Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [2023_VLA_Rt-2].
 The proposed BFM paradigm would operate in two stages.
@@
 Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot.
-The potential of this approach is highlighted by recent works, with frameworks like FLaRe [117] enhancing generalisation for long-horizon tasks, MOTO [118] enabling effective offline-to-online adaptation from images, and AdA [119] demonstrating in-context adaptation to novel environments.
+The potential of this approach is highlighted by recent works, with frameworks like FLaRe [2024_FLaRe_RLfinetune_foundationmodel] enhancing generalisation for long-horizon tasks, MOTO [2024_MOTO_RLfinetune_foundationmodel] enabling effective offline-to-online adaptation from images, and AdA [2023_RL_finetune_RMFs] demonstrating in-context adaptation to novel environments.
 Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
-However, realizing the BFM concept for bipeds presents significant challenges.
-The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [117, 119].
+However, realising the BFM concept for bipeds presents significant challenges.
+The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [2024_FLaRe_RLfinetune_foundationmodel, 2023_RL_finetune_RMFs].
 Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
@@
 Rather than relying on large-scale pre-training, this conceptual framework would adopt a modular, hierarchical approach.
-The idea is to extend conventional hierarchical frameworks (discussed in Section III) with explicitly adaptive layers, allowing for the dynamic composition of specialized policies.
-The core principle of this concept would be modularity, enabling each layer to be independently optimized or replaced and providing interpretable outputs at each stage.
+The idea is to extend conventional hierarchical frameworks (discussed in Section III) with explicitly adaptive layers, allowing for the dynamic composition of specialised policies.
+The core principle of this concept would be modularity, enabling each layer to be independently optimised or replaced and providing interpretable outputs at each stage.
 A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier.
 Each layer processes context-specific inputs and outputs interpretable commands.
-The HL reasoning layer leverages large pre-trained model such as LLMs and VLMs [120, 121] to parse commands into sub-tasks.
-For instance, Vision-Language Model Predictive Control [93] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
+The HL reasoning layer leverages large pre-trained models such as LLMs and VLMs [2022_DoAsICan, 2022_codeaspolicy_google] to parse commands into sub-tasks.
+For instance, Vision–Language Model Predictive Control [2024_VLM_PC] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
 By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
-The mid-level planner selects or synthesises motions via learned motion libraries [58, 44] or DRL-based planners [122].
+The mid-level planner selects or synthesises motions via learned motion libraries [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter] or DRL-based planners [2021_Mohammadreza_DRL_improveparameters_forplanner].
 The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands.
-These include locomotion primitives like walking and climbing [123], adaptive tracking controllers for whole-body tracking [59], and imitation-based skills such as kicking and dancing [26], by utilizing RL, IL, or model-based methods.
-This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [123].
-However, realizing the MLAM concept would introduce challenges distinct from BFMs.
-Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [109].
-Additionally, the computational latency incurred by HL reasoning modules like LLMs [123] would pose limitations for tasks needing rapid reactions.
+These include locomotion primitives like walking and climbing [2024_yangOu_LLM_hierarchical_quadrupedrobot], adaptive tracking controllers for whole-body tracking [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and imitation-based skills such as kicking and dancing [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task], by utilising RL, IL, or model-based methods.
+This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [2024_yangOu_LLM_hierarchical_quadrupedrobot].
+However, realising the MLAM concept would introduce challenges distinct from BFMs.
+Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [TAMOLS].
+Additionally, the computational latency incurred by HL reasoning modules like LLMs [2024_yangOu_LLM_hierarchical_quadrupedrobot] would pose limitations for tasks needing rapid reactions.
 VI Conclusion
-Despite significant progress in DRL for robotics, a significant gap remains between current achievements and the development of a unified framework that can efficiently handle a wide range of complex tasks.
+Despite significant progress in DRL for robotics, a substantial gap remains between current achievements and the development of a unified framework capable of efficiently handling a wide range of complex tasks.
 DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks.
-End-to-end frameworks have demonstrated success in handling diverse locomotion skills [35], climbing stairs [49], and navigating challenging terrains such as stepping stones [33].
+End-to-end frameworks have demonstrated success in handling diverse locomotion skills [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], climbing stairs [2021_siekmann_blind_DRL_stair], and navigating challenging terrains such as stepping stones [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark].
 Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
-Each framework brings unique strengths to the unified framework.
-End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills.
+Each framework contributes unique strengths to the pursuit of a unified framework.
+End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility required for robots to acquire diverse locomotion skills.
 In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies.
-While DRL has enabled remarkable progress, our survey concludes that current frameworks confront key limitations, including the tension between multi-skill generalisation and task-specific precision, the persistent sim-to-real gap, and critical safety concerns.
-To address these challenges, this survey synthesizes specific pathways for future research and identifies key opportunities for cross-pollination from related fields, such as FMs, loco-manipulation, and quadrupedal robotics.
+While DRL has enabled remarkable progress, our survey concludes that current frameworks face key limitations, including the tension between multi-skill generalisation and task-specific precision, the persistent sim-to-real gap, and critical safety concerns.
+To address these challenges, this survey synthesises specific pathways for future research and identifies key opportunities for cross-pollination from related fields, such as FMs, loco-manipulation, and quadrupedal robotics.
 These insights culminate in our proposal of two conceptual frameworks: the BFMs, extending the end-to-end paradigm, and the MLAMs, evolving from the hierarchical approach, which together offer distinct blueprints for the next generation of generalist bipedal controllers.
@@
 Figure 5: Diagram for RL algorithms catalogue
-The advancement and development of RL is crucial for bipedal locomotion.
-Specifically, advancements in deep learning provide deep NNs serving as function approximators to empower RL with the capability to handle tasks characterised by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data.
-In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms.
-Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
+The advancement and development of RL are crucial for bipedal locomotion.
+Specifically, advances in deep learning provide deep NNs that serve as function approximators, enabling RL to handle tasks characterised by high-dimensional and continuous spaces by efficiently discovering condensed, low-dimensional representations of complex data.
+In comparison with other robots of different morphologies, such as wheeled robots, bipedal robots possess far higher DoFs and continuously interact with their environments, which results in greater demands on DRL algorithms.
+In particular, within legged locomotion, policy-gradient-based algorithms are prevalent in bipedal locomotion research.
 Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks.
-Multi-layer perceptrons (MLP), a fundamental NN, excel in straightforward regression tasks with lower computational resource demands.
-A comprehensive comparison between MLP and the memory-based NN, LSTM reveals that MLPs have an advantage in convergence speed for tasks [80].
-However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modelling key physical properties vital for periodical gaits [32] and successful sim-to-real transfer in bipedal locomotion.
-Additionally, Convolutional Neural Networks (CNN) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial.
-This diverse range of NN architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
-Considering DRL algorithms, recent bipedal locomotion studies focus on model-free RL algorithms.
+Multi-Layer Perceptrons (MLPs), a fundamental NN structure, excel in straightforward regression tasks with lower computational resource requirements.
+A comprehensive comparison between MLPs and the memory-based NN LSTM reveals that MLPs have an advantage in convergence speed for many tasks [2023_Rohan_sim2real_currentdecompose].
+However, LSTMs, as variants of Recurrent Neural Networks (RNNs), are adept at processing time-associated data, effectively relating different states across time and modelling key physical properties vital for periodic gaits [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] and successful sim-to-real transfer in bipedal locomotion.
+Additionally, Convolutional Neural Networks (CNNs) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial.
+This diversity of NN architectures highlights the importance of selecting an appropriate model based on the specific requirements of bipedal locomotion tasks.
+Considering DRL algorithms, recent bipedal locomotion studies have focused on model-free RL algorithms.
 Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model.
-Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
+Although model-free RL requires more computational samples and resources, it can train more robust policies that allow robots to traverse challenging environments.
 Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches.
-Value-based methods e.g.
-Q-learning, State-Action-Reward-State-Action (SARSA), Deep Q-learning (DQN) [72] only excel in discrete action space and often struggle with high dimensional action space.
+Value-based methods, e.g.
+Q-learning, State–Action–Reward–State–Action (SARSA), and Deep Q-learning (DQN) [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones], excel only in discrete action spaces and often struggle with high-dimensional action spaces.
 Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy.
 SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning.
-DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to the difficulty of accurate value estimation.
-In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
-More advanced algorithms combine both policy-based methods and value-based methods.
-Actor-critic (AC) refers to a main idea simultaneously learning both a policy (actor) and a value function (critic), where it owns both advantages of both algorithms [124, 125].
-Popular algorithms e.g.
-Trust region policy optimisation (TRPO) [126] and PPO based on policy-based methods, borrow ideas from AC.
-Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [66], Twin Delayed Deep Deterministic Policy Gradients (TD3) [127], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [128], SAC (Soft Actor-Critic) [129].
-Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario.
-There are several key factors to value these algorithms such as: sample efficiency, robustness and generalisation, and implementation challenges.
-A comparative analysis work [65] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
-In [64], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilising fewer resources than TRPO.
-In terms of training time, PPO is much faster than SAC, and DDPG algorithms [65].
-Besides, many works [32, 58, 47] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various NN architectures have made PPO the most popular choice in this field.
-Various work has demonstrated that PPO can conduct the exploration of walking [32], jumping [48], stair climbing [49], and stepping stones [33], which demonstrates its efficiency, robustness and generalisation.
-Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimising sampling efficiency.
-In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [34, 130].
-TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [129].
+DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to difficulties in accurate value estimation.
+In contrast, policy-based methods, such as policy-gradient techniques, can handle complex tasks but are generally less sample-efficient than value-based methods.
+More advanced algorithms combine both policy-based and value-based methods.
+The Actor–Critic (AC) framework simultaneously learns both a policy (actor) and a value function (critic), combining the advantages of both approaches [2015_continous_control_deep_reinforcement_learning, 2016_guided_learning_of_Control_Graphs_physics-based_characters].
+Popular algorithms such as Trust Region Policy Optimisation (TRPO) [2015_TRPO_foundamentalpaper] and PPO, based on policy-based methods, borrow ideas from AC.
+Moreover, other novel algorithms based on the AC framework include Deep Deterministic Policy Gradient (DDPG) [2023_DDPG_e2e_openaigym_mujoco], Twin Delayed Deep Deterministic Policy Gradients (TD3) [2019_stephen_TD3_instruction], A2C (Advantage Actor–Critic), A3C (Asynchronous Advantage Actor–Critic) [2022_Jie_A3C_sim2real], and SAC (Soft Actor–Critic) [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
+Each algorithm has its strengths for different tasks in bipedal locomotion scenarios.
+Several key factors determine their performance, such as sample efficiency, robustness and generalisation, and implementation complexity.
+A comparative analysis [2023_Omur_comparison_RLalgorithm_bipedallocomotion] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO, which attains relatively high scores.
+In [2017_Schulman_PPO_instructure], PPO demonstrates robustness and computational efficiency in complex scenarios such as bipedal locomotion, utilising fewer resources than TRPO.
+In terms of training time, PPO is much faster than SAC and DDPG [2023_Omur_comparison_RLalgorithm_bipedallocomotion].
+Moreover, many studies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2020_Siekmann_drl_e2e_s2r_cassie] have demonstrated its robustness and ease of implementation.
+Combined with its flexibility to integrate with various NN architectures, this has made PPO the most popular choice in the field.
+Numerous studies have shown that PPO can enable the exploration of walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], jumping [2023_UCB_Cassie_RL_Jumping], stair climbing [2021_siekmann_blind_DRL_stair], and stepping-stone traversal [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark], demonstrating its efficiency, robustness, and generalisation.
+Additionally, the DDPG algorithm integrates the Actor–Critic framework with DQN to facilitate off-policy training, further optimising sample efficiency.
+In certain scenarios, such as jumping, DDPG shows higher rewards and better learning performance than PPO [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2022_Chongben_pallel_DDPG_e23_silumation_only].
+TD3, developed from DDPG, improves upon the performance of both DDPG and SAC [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
 SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3.
-Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences to train without requiring constant interactions with the environment.
+Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences for training without requiring constant interaction with the environment.
 This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential.
-SAC is also known for its stability and strong performance across a wide range of tasks [129].
-While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning.
-Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
+SAC is also known for its stability and strong performance across a wide range of tasks [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
+While A2C offers improved efficiency and stability compared with A3C, the asynchronous update mechanism of A3C provides better exploration capability and accelerates learning.
+Although these algorithms demonstrate clear advantages, they are more challenging to apply owing to their complexity compared with PPO.
 Appendix B Training Simulation Environment
-The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators.
-Creating a reliable simulation environment and conducting RL training is difficult.
-Through the literature, several simulators are available, including Isaac Gym [67], robocup3D [76], Openai Gym [77], MuJoCo [78], Orbit [131], Brax [132], Isaac Lab [133].
+The development of DRL algorithms and sim-to-real techniques highlights the requirement for high-quality simulators.
+Creating a reliable simulation environment and conducting RL training is challenging.
+The literature shows that several simulators are available, including Isaac Gym [2021_Sim_Isaac], RoboCup3D [2003_Sim_Robocup], OpenAI Gym [2016_Sim_OpenAI], MuJoCo [2012_Sim_Mujoco], Orbit [2023_Orbit_simulator], Brax [2021_Brax], and Isaac Lab [isaaclab2023].
 OpenAI developed Gym and Gymnasium to provide lightweight environments for rapid testing of RL algorithms, including simplified bipedal locomotion models.
 RoboCup also serves as a benchmark platform for RL research and development in multi-agent settings.
-For physics-based simulation, MuJoCo developed by DeepMind, and Gazebo are widely used platforms that support a range of robotics research tasks.
+For physics-based simulation, MuJoCo, developed by DeepMind, and Gazebo are widely used platforms that support a range of robotics research tasks.
 NVIDIA’s Isaac Gym, although now deprecated, played an important role as a high-performance GPU-based simulator for training agents in complex environments.
 Its successors, such as Isaac Lab and Orbit, continue to evolve as modern RL and robotics frameworks.
-One of the most crucial things is the parallelisation strategy and GPU simulations.
-For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms with particular emphasis on simulations that require large numbers of environment instances executing in parallel.
-Running the physics simulation on GPU can result in significant speedups, especially for large scenes with thousands of individual actors.
+One of the most crucial aspects is the parallelisation strategy and GPU simulation.
+For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms, with particular emphasis on simulations requiring large numbers of environment instances executing in parallel.
+Running the physics simulation on a GPU can result in significant speed-ups, especially for large scenes with thousands of individual actors.

### v7

@@
 (e) Atlas: a fully electric humanoid robot developed by Boston Dynamics.
-As a result, bipedal robots hold significant potential for real-world applications [2024_tong_advancements_review].
-In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [industrial_robotics_applications, collaborative_industrial, 2023_vidoeo_demo_digit_locomaniputlation].
+As a result, bipedal robots hold significant potential for real-world applications [110].
+In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [26, 120, 1].
 Their agility is particularly advantageous in complex environments such as multi-level workplaces.
-Bipedal robots are also well suited to tasks that involve the use of human-designed tools, making them valuable for assisting in daily activities, healthcare, and rehabilitation [2019_bingjing_human_exoskeleton].
-Moreover, they show considerable promise in search-and-rescue operations, where they can navigate hazardous and unpredictable terrains [2015_underwaterrobot_review, quadSpace1, quadSpace2].
-Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot].
-Early methods, such as the Linear Inverted Pendulum Model (LIPM) [2013_3DSLIPmodel_running], provided simplified representations of the dynamics involved in bipedal motion, enabling easier analysis and control.
+Bipedal robots are also well suited to tasks that involve the use of human-designed tools, making them valuable for assisting in daily activities, healthcare, and rehabilitation [9].
+Moreover, they show considerable promise in search-and-rescue operations, where they can navigate hazardous and unpredictable terrains [11, 94, 87].
+Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [37, 91, 14].
+Early methods, such as the Linear Inverted Pendulum Model (LIPM) [113], provided simplified representations of the dynamics involved in bipedal motion, enabling easier analysis and control.
 As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion.
-Advanced methods such as Model Predictive Control (MPC) [2022_Hou_drlasparametertoMPC, 2023_junheng_steppingstones_MPCWBC] and Trajectory Optimisation (TO) [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2017_biped_model-based_method_review] exploit predefined dynamic models to solve constrained optimal-control problems that plan footsteps, centre-of-mass (CoM) motion, and contact forces.
+Advanced methods such as Model Predictive Control (MPC) [42, 59] and Trajectory Optimisation (TO) [40, 60, 37] exploit predefined dynamic models to solve constrained optimal-control problems that plan footsteps, centre-of-mass (CoM) motion, and contact forces.
 While model-based approaches offer rapid convergence and predictive capabilities, they often struggle in dynamically complex and uncertain environments where adaptability is essential.
-Reinforcement learning (RL)-based methods, particularly deep reinforcement learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [2020_systematic_review], which provides a distinct advantage.
+Reinforcement learning (RL)-based methods, particularly deep reinforcement learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [54], which provides a distinct advantage.
 Unlike model-based approaches, which rely on predefined dynamics and may fail under unforeseen conditions, DRL enables robots to autonomously discover control strategies through trial and error, achieving greater adaptability and robustness in diverse environments.
@@
 The evolution of RL in bipedal robotics has largely advanced through the end-to-end learning paradigm.
-Early studies in 2004 applied simple policy-gradient methods to 2D bipeds [2004_Tedrake_RL_bipedal_policygradient, 2004_model-based_RL_learnpoincaremap_simple_bipedrobot], while later breakthroughs in DRL enabled policy training in high-fidelity physics simulators [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_wenhao_yu_DRL_withoutpredefine_symmetrygait_PPO_jointanglePD_].
+Early studies in 2004 applied simple policy-gradient methods to 2D bipeds [108, 78], while later breakthroughs in DRL enabled policy training in high-fidelity physics simulators [84, 83, 125].
 As robotic hardware matured, an increasing variety of bipedal and humanoid platforms emerged, supporting extensive evaluation of DRL-based locomotion across diverse morphologies, as illustrated in Fig.
@@
 This evolution marked the transition from purely simulation-based training to simulation-to-real (sim-to-real) transfer, where policies trained in simulators are deployed on physical robots.
-In 2020, the first successful sim-to-real transfer of an end-to-end DRL locomotion policy was achieved on the 3D torque-controlled bipedal robot Cassie [2020_Xie_firstsim2real_].
-Subsequent work explored two principal learning paradigms: reference-based learning, which leverages TO-generated data or motion-capture data to guide policy training [2021_michael_velocity-basedcontrol_motioncapturedata, 2024_speratebody_expressive_allmotion, 2024_tang_humanmimic_adversarial, 2024_Zhang_wholebody_adversarial_motion_priors]; and reference-free learning, where policies are trained entirely from scratch to autonomously discover control strategies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie].
-These developments demonstrate that end-to-end frameworks can achieve robust and versatile locomotion skills across complex terrains and dynamic environments [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
+In 2020, the first successful sim-to-real transfer of an end-to-end DRL locomotion policy was achieved on the 3D torque-controlled bipedal robot Cassie [116].
+Subsequent work explored two principal learning paradigms: reference-based learning, which leverages TO-generated data or motion-capture data to guide policy training [107, 19, 104, 126]; and reference-free learning, where policies are trained entirely from scratch to autonomously discover control strategies [98].
+These developments demonstrate that end-to-end frameworks can achieve robust and versatile locomotion skills across complex terrains and dynamic environments [25, 105, 63].
 Similarly, hierarchical structures have garnered significant interest.
@@
 Hybrid architectures often integrate learning-based and model-based modules to combine adaptability with physical consistency.
-One representative design couples a learned HL planner with an LL model-based controller, forming a cascade-structure or deep-planning hybrid scheme [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller, 2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN].
-Alternatively, DRL feedback-control hybrids embed learned control policies within model-based feedback loops to enhance tracking precision and disturbance rejection [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
-Learned hierarchical control schemes [2017_DRL_survey] decompose locomotion into multiple layers, each focusing on specific functions such as navigation and fundamental locomotion skills [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
+One representative design couples a learned HL planner with an LL model-based controller, forming a cascade-structure or deep-planning hybrid scheme [60, 23, 17].
+Alternatively, DRL feedback-control hybrids embed learned control policies within model-based feedback loops to enhance tracking precision and disturbance rejection [101, 112].
+Learned hierarchical control schemes [3] decompose locomotion into multiple layers, each focusing on specific functions such as navigation and fundamental locomotion skills [84, 83, 127].
 To provide a clearer overview of the current landscape, we categorise existing DRL frameworks as shown in Fig.
@@
 Establishing such a framework is essential for consolidating diverse learning pipelines, standardising evaluation metrics, and enabling transferable locomotion capabilities across different robot morphologies.
-As locomotion tasks become increasingly complex, ranging from basic stabilisation to dynamic parkour and loco-manipulation, the need for consistent benchmarking has intensified.The DARPA Robotics Challenge exemplified this trend by introducing one of the first large-scale evaluation platforms for bipedal humanoids performing real-world locomotion and manipulation tasks [2015_DAPRA_Challenge_1], highlighting the importance of robustness and practical deployment.
-Although several reviews discuss RL for general robotics [2020_systematic_review] and model-based methods for bipedal robots [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot], none specifically focus on DRL-based frameworks for bipeds.
+As locomotion tasks become increasingly complex, ranging from basic stabilisation to dynamic parkour and loco-manipulation, the need for consistent benchmarking has intensified.The DARPA Robotics Challenge exemplified this trend by introducing one of the first large-scale evaluation platforms for bipedal humanoids performing real-world locomotion and manipulation tasks [4], highlighting the importance of robustness and practical deployment.
+Although several reviews discuss RL for general robotics [54] and model-based methods for bipedal robots [37, 91, 14], none specifically focus on DRL-based frameworks for bipeds.
 To address this gap, this survey reviews relevant literature according to the following selection criteria: (1) studies that investigate DRL frameworks specifically for bipedal robots; (2) research involving both simulated and physical bipedal robots; (3) approaches that improve policy transfer from simulation to real-world environments; and (4) publications from the last five years (2018–April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, arXiv, and major robotics venues such as CoRL, RSS, ICRA, IROS, and Humanoids.
@@
 II End-to-end framework
-The end-to-end DRL framework represents a holistic approach in which a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, LiDAR data, or proprioceptive feedback [2017_compare_actionspace_RL], together with user commands [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] or pre-defined references [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], into joint-level control actions $\mathcal{U}$.
+The end-to-end DRL framework represents a holistic approach in which a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, LiDAR data, or proprioceptive feedback [86], together with user commands [98] or pre-defined references [61], into joint-level control actions $\mathcal{U}$.
 Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function.
@@
 Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
-Methods
-Works
-Capabilities
-Characteristic
-Advantages and Disadvantages
-Implementation Flow Chart
-Residual learning
-[2018_xie_OSUDRL_feedback_DRL_reference-based]
-[2020_Xie_firstsim2real_]
-Forward walk
-unidirectional walk
-Adding a residual term to the known motor positions at the current time step.
+Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
+Residual learning [115] [116] Forward walk unidirectional walk Adding a residual term to the known motor positions at the current time step.
 A: Fast convergence speed D: Requires high-quality predefined reference, limits to specific motions, and lacks robustness to complex terrains.
-Guided learning
-[2020_Siekmann_drl_e2e_s2r_cassie]
-[2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter]
-[2023_UCB_Cassie_RL_Jumping]
-[2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining]
-Forward walk
-Versatile walk
-Versatile jump
-Versatile motions
-Mimic the predefined reference and directly specifies joint-level commands.
-A: Accelerates the learning process and is robust across terrains. D: Limits to the predefined motions and lacks adaptability to unforeseen changes in environment.
-Reference-free learning
-[2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie]
-[2021_siekmann_blind_DRL_stair]
-[2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark]
-Periodic motions
-Stepping stones walk
-Visual walk
-Learn locomotion skills from scratch without any prior knowledge.
-A: High potential for gait exploration, highly robust to complex terrain D: Requires intensive reward shaping for gait patterns and is relatively expensive in computational resources
+[Uncaptioned image]
+Guided learning [100] [61] [62] [63] Forward walk Versatile walk Versatile jump Versatile motions Mimic the predefined reference and directly specifies joint-level commands.
+A: Accelerates the learning process and is robust across terrains.
+D: Limits to the predefined motions and lacks adaptability to unforeseen changes in environment.
+[Uncaptioned image]
+Reference-free learning [98] [99] [25] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge.
+A: High potential for gait exploration, highly robust to complex terrain D: Requires intensive reward shaping for gait patterns and is relatively expensive in computational resources [Uncaptioned image]
 II-A Reference-based learning
@@
 This holistic approach enables the biped robot to adapt to various dynamic conditions while maintaining balance and control.
-Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [2018_xie_OSUDRL_feedback_DRL_reference-based].
+Introduced in 2018, a residual learning framework for the bipedal robot Cassie marked a significant advancement [115].
 This framework allowed the robot to walk forward by incorporating a policy trained via Proximal Policy Optimisation (PPO) algorithms, as detailed in Appendix A.
@@
 These modified references are then processed by a PD controller to set the desired joint positions.
-Although this framework has improved the robot’s ability to perform tasks beyond standing [2018_yang_push_recover_wholebody], it has yet to be physically deployed on a bipedal robot.
+Although this framework has improved the robot’s ability to perform tasks beyond standing [119], it has yet to be physically deployed on a bipedal robot.
 As a result, it remains impractical for managing walking at varying speeds and is constrained to movement in a single direction.
-To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied to a physical bipedal robot [2020_Xie_firstsim2real_].
+To transition this framework to a real robot, a sim-to-real strategy based on the previous model was demonstrated, where the policy, trained through a residual learning approach, was subsequently applied to a physical bipedal robot [116].
 Compared to model-based methods, this training policy achieves faster running speeds on the same platform, underlining the considerable potential of DRL-based frameworks.
 However, the robot’s movements remain constrained to merely walking forward or backward.
-A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [2021_Diego_deepwalk_curriculum_learning_sim2real].
+A unique residual learning approach was introduced to enable omnidirectional walking, where the policy adds a residual term to the current joint positions, allowing gradual omnidirectional walking [92].
 In this case, the desired reference is the robot’s current joint positions, which makes the approach distinctive.
@@
 Residual learning enhances an existing control policy by taking current joint positions or states and applying a residual action to adjust reference actions for better performance.
-Compared to other learning approaches that directly output joint positions, it is highly sample efficient [2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller].
+Compared to other learning approaches that directly output joint positions, it is highly sample efficient [23].
 However, when predefined references are unstable or of low quality, residual learning may struggle, especially on complex terrains, as the action space is bounded by the reference, limiting the ability to handle unpredictable or uneven terrains.
@@
 In this approach, the reward structure is centred on accurately imitating predefined reference trajectories, ensuring precise alignment between the policy output and the reference motion.
-A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [2020_Siekmann_drl_e2e_s2r_cassie].
+A sim-to-real framework that employs periodic references to initiate the training phase was proposed in [100].
 In this framework, the action space directly maps to the joint angles, and desired joint positions are managed by joint PD controllers.
@@
 However, this model is limited to a single locomotion goal: forward walking.
-A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
+A more diverse and robust walking DRL framework that includes a Hybrid Zero Dynamics (HZD) gait library was demonstrated [61], achieving a significant advancement by enabling a single end-to-end policy to facilitate walking, turning, and squatting.
 Despite these advancements, the parameterisation of reference motions introduces constraints that limit the flexibility of the learning process and the policy’s response to disturbances.
-To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [2023_UCB_Cassie_RL_Jumping].
+To broaden the capabilities of guided learning policies, a framework capable of handling multiple targets, including jumping, was developed [62].
 This approach introduced a novel policy structure that integrates long-term input/output (I/O) encoding, complemented by a multi-stage training methodology that enables the execution of complex jumping manoeuvrers.
-An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [2024_Zhang_wholebody_adversarial_motion_priors].
+An adversarial motion priors approach, employing a style reward mechanism, was also introduced to facilitate the acquisition of user-specified gait behaviours [126].
 This method improves the training of high-dimensional simulated agents by replacing complex hand-designed reward functions with more intuitive controls.
-While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining] based on the foundational work in [2023_UCB_Cassie_RL_Jumping].
+While previous works primarily focused on specific locomotion skills, a unified framework that accommodates both periodic and non-periodic motions was further developed [63] based on the foundational work in [62].
 This framework enhances the learning process by incorporating a wide range of locomotion skills and introducing a dual I/O history approach, marking a significant breakthrough in creating a robust, versatile, and dynamic end-to-end framework.
@@
 Guided learning methods expedite the learning process by leveraging expert knowledge and demonstrating the capacity to achieve versatile and robust locomotion skills.
-Through the comprehensive evaluation [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], it is demonstrated that guided learning employs references without complete dependence on them.
+Through the comprehensive evaluation [63], it is demonstrated that guided learning employs references without complete dependence on them.
 Conversely, residual learning exhibits failures or severe deviations when predicated on references of inferior quality.
@@
 However, the benefits of reference-based learning come with inherent limitations.
-Reliance on predefined trajectories often confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction].
+Reliance on predefined trajectories often confines the policy to specific gaits, restricting its capacity to explore a broader range of motion possibilities [61, 111].
 Moreover, such methods exhibit reduced adaptability when confronted with novel environments or unforeseen perturbations.
 These limitations are further compounded by the difficulty of acquiring high-quality and task-relevant demonstrations.
-Common sources of prior knowledge include TO [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], human motion capture [2024_speratebody_expressive_allmotion], teleoperation [2023_seo_deep_loco_manipulation, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and scripted controllers [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+Common sources of prior knowledge include TO [60, 36, 61, 63], human motion capture [19], teleoperation [97, 31], and scripted controllers [83].
 While informative, these demonstrations often require adaptation due to embodiment mismatch or limited generalisability.
-Motion retargeting [2019_iit_wholebody_retargeting, 2024_tang_humanmimic_adversarial, 2017_Ayusawa_motionretarget_optimization], as one of promising direction, addresses this by converting human-centric motions into robot-feasible trajectories, yet it still struggles with preserving fidelity and adapting across morphologies.
+Motion retargeting [82, 104, 6], as one of promising direction, addresses this by converting human-centric motions into robot-feasible trajectories, yet it still struggles with preserving fidelity and adapting across morphologies.
 Ultimately, the success of guided learning relies not only on using references but on accessing high-quality, adaptable demonstrations that generalise across tasks and platforms—highlighting a key challenge in advancing robust policy learning.
@@
 The action space and observation space in this approach are similar to the guided-learning method; however, the reward structure differs significantly from the reference-based method.
-Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction].
+Instead of focusing on imitating predefined motions, the reward emphasises learning efficient gait patterns by capturing the distinctive characteristics of bipedal locomotion [111].
 The concept of reference-free learning was initially explored using simulated physics engines with somewhat unrealistic bipedal models.
-A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [2018_wenhao_yu_DRL_withoutpredefine_symmetrygait_PPO_jointanglePD_].
+A pioneering framework, which focused on learning symmetric gaits from scratch without the use of motion capture data, was developed and validated within a simulated environment [125].
 This framework introduced a novel term into the loss function and utilised a curriculum learning strategy to effectively shape gait patterns.
-Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model (Cassie), though this has yet to be validated outside of simulation [2020_zhaoming_drl_steppingstones_PPOwithactorcritic_referencefree_simulation].
+Another significant advancement was made in developing a learning method that enabled a robot to navigate stepping stones using curriculum learning, focusing on a physical robot model (Cassie), though this has yet to be validated outside of simulation [117].
 Considering the practical implementation of this approach, significant efforts have been made to develop sim-to-real reference-free frameworks, and their potential has been further explored on physical robots.
-A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie].
+A notable example of such a framework accommodates various periodic motions, including walking, hopping, and galloping [98].
 This framework employs periodic rewards to facilitate initial training within simulations before successfully transitioning to a physical robot.
 It has been further refined to adapt to diverse terrains and scenarios.
-For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [2021_siekmann_blind_DRL_stair].
-Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation], thus enabling the robot to effectively navigate stepping stones [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark].
-Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [2023_duan_OSUDRL_heightmap_visionbased_hybrid].
+For instance, robust blind walking on stairs was demonstrated through terrain randomisation techniques in [99].
+Additionally, the integration of a vision system has enhanced the framework’s ability to precisely determine foot locations [24], thus enabling the robot to effectively navigate stepping stones [25].
+Subsequent developments include the incorporation of a vision system equipped with height maps, leading to an end-to-end framework that more effectively generalises terrain information [72].
 This approach to learning enables the exploration of novel solutions and strategies that might not be achievable through mere imitation of existing behaviours.
@@
 TABLE II: Summary and comparison of hierarchical frameworks.
-Control Scheme
-Works
-Module
-Characteristics
-Advantages and Disadvantages
-Deep Planning Hybrid Scheme
-[2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller]
-[2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller]
-[2024_gaspard_footstepnet_soccer]
-Deep planning + ID
-Deep planning + ID-QP
-Deep planning + WPG
-HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
+Control Scheme Works Module Characteristics Advantages and Disadvantages
+Deep Planning Hybrid Scheme [23] [15] [35] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
 A: Enhanced command tracking capabilities, generalised across different platforms, sampling-efficient, and robust.
 D: Complicated system and communication between layers, requires a precise model, and lacks generalisation across different tasks.
-Feedback DRL Control Hybrid Scheme
-[2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real]
-[2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P]
-[2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]
-Gait library + feedback policy
-Footstep planner + feedback policy
-Model-based planner + feedback policy
-LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
-A: Short inference time, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, which reduces sampling efficiency.
-Learned Hierarchy Framework
-[2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.]
-[2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot]
-[2024_stanford_humanplus_hierarchical_humanoidshadowingimitation]
-HL policy + LL policy
-HL policy + LL policy
-HL policy + LL policy
-Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learns navigation skills.
+Feedback DRL Control Hybrid Scheme [36] [101] [112] Gait library + feedback policy Footstep planner + feedback policy Model-based planner + feedback policy LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
+A: Short inference time, robust, navigation locomotion capabilities, interpretability.
+D: Complicated system and communication between layers, which reduces sampling efficiency.
+Learned Hierarchy Framework [84] [127] [31] HL policy + LL policy HL policy + LL policy HL policy + LL policy Both HL planner and LL feedback controller are learned.
+LL policy focuses on basic locomotion skills; on the other side, HL policy learns navigation skills.
 A: Provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy.
@@
 Several studies have demonstrated the integration of an HL planner policy with a model-based controller to achieve tasks in world space.
-A notable framework optimises task space level performance, eschewing direct joint-level and balancing considerations [2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller].
+A notable framework optimises task space level performance, eschewing direct joint-level and balancing considerations [23].
 This system combines a residual learning planner with an inverse dynamics controller, enabling precise control over task-space commands to joint-level actions, thereby improving velocity tracking, foot touchdown location, and height control.
-Further advancements include a hybrid framework that merges HZD-based residual deep planning with model-based regulators to correct errors in learned trajectories, showcasing robustness, training efficiency, and effective velocity tracking [2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN].
+Further advancements include a hybrid framework that merges HZD-based residual deep planning with model-based regulators to correct errors in learned trajectories, showcasing robustness, training efficiency, and effective velocity tracking [17].
 These frameworks have been successfully transferred from simulation to reality and validated on robots such as Cassie.
 However, the limitations imposed by residual learning constrained the agents’ capacity to explore a broader array of possibilities.
-Building on previous work [2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller].
+Building on previous work [17], a more efficient hybrid framework was developed, which learns from scratch without reliance on prior knowledge [15].
 In this approach, a purely learning-based HL planner interacts with an LL controller using an Inverse Dynamics with Quadratic Programming formulation (ID-QP).
@@
 In parallel, several research teams have focused on developing navigation and locomotion planners for humanoid robots, leveraging onboard visual perception and learned control strategies.
-Recent work [2024_gaspard_footstepnet_soccer] explored complex dynamic motion tasks such as playing soccer by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a CoM trajectory.
+Recent work [35] explored complex dynamic motion tasks such as playing soccer by integrating a learned policy with an online footstep planner that utilises weight positioning generation (WPG) to create a CoM trajectory.
 This configuration, coupled with a whole-body controller, enables dynamic activities like soccer shooting.
 Although these systems demonstrate promising coordination between perception, planning, and control, they remain limited in dynamic movement capability compared to full-sized humanoid robots, and thus primarily address navigation and task-level execution.
-Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller] and sophisticated navigation capabilities [2024_gaspard_footstepnet_soccer].
+Regarding generalisation, these frameworks have shown potential for adaptation across different types of bipedal robots with minimal adjustments, demonstrating advanced user command tracking [15] and sophisticated navigation capabilities [35].
 However, limitations are evident, notably the absence of capabilities for executing more complex and dynamic motions, such as jumping.
@@
 These LL controllers, replacing traditional model-based feedback mechanisms, work in conjunction with HL planners that process terrain information, plan future walking paths, and maintain robust locomotion stability.
-For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real].
+For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [36].
 Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness.
 A more dynamic approach involves online planning, which has shown greater adaptability and efficiency.
-One notable framework combines a conventional foot planner with an LL DRL policy [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands.
+One notable framework combines a conventional foot planner with an LL DRL policy [101], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands.
 Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies.
 However, this work has not yet been transferred from simulation to real-world applications.
-Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
+Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [112].
 This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
 A recent development has shown that focusing solely on foot placement might restrict the stability and adaptability of locomotion, particularly in complex maneuvers.
-A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [2024_Li_model_based_LLDRL_hybrid_control_scheme].
+A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [58].
 This system employs a residual learning architecture, where the DRL policy’s outputs are merged with the planner’s directives before being relayed to the PD controller.
@@
 Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities.
-These policies have demonstrated high efficiency in locomotion and accurate reference tracking [quadRLLL].
+These policies have demonstrated high efficiency in locomotion and accurate reference tracking [51].
 Despite their extensive capabilities, such policies generally require short inference time, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited.
-Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [quadRLLL].
-Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
+Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [51].
+Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [112].
 This capability effectively addresses the navigational limitations observed in end-to-end frameworks.
@@
 Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models.
-Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation, and hinder rapid deployment and iterative enhancements [quadRLLL].
+Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation, and hinder rapid deployment and iterative enhancements [51].
 III-C Learned hierarchy framework
@@
 Subsequently, an HL policy is developed to direct the robot towards specific targets, encapsulating a structured approach to robotic autonomy.
-The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.].
+The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [84].
 In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance.
@@
 This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal.
-The framework was later enhanced to include imitation learning (IL), facilitating the replication of dynamic human-like movements within the simulation environment [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+The framework was later enhanced to include imitation learning (IL), facilitating the replication of dynamic human-like movements within the simulation environment [83].
 However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks has been validated only in simulations.
 The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
-Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
-To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation].
+Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [127].
+To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [31].
 It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as an LL tracking controller.
@@
 Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics.
-Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.].
+Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [84].
 As discussed in the previous section, the benefits of integrating RL-based planners with RL-based controllers have been effectively demonstrated.
@@
 Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
-Additionally, the training process for each policy within the hierarchy demands considerable computational resources [2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
+Additionally, the training process for each policy within the hierarchy demands considerable computational resources [127].
 The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions.
@@
 This capability gap is evident in the current literature.
-Many approaches excel at generalisation, demonstrating policies that enable versatile skills such as walking and jumping [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining] and can transfer to different terrains [2021_siekmann_blind_DRL_stair, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_radosavovic_realworld_transformer_NNmodel].
-However, these generalised policies often lack the fidelity required for high-precision tasks such as exact foot placement [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation] or maintaining a specific velocity with minimal error [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction].
-Conversely, controllers specialised for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [2023_UCB_Cassie_RL_Jumping], yet they cannot generalise these capabilities to a broader range of tasks.
+Many approaches excel at generalisation, demonstrating policies that enable versatile skills such as walking and jumping [61, 63] and can transfer to different terrains [99, 25, 88].
+However, these generalised policies often lack the fidelity required for high-precision tasks such as exact foot placement [101, 25, 24] or maintaining a specific velocity with minimal error [111].
+Conversely, controllers specialised for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [62], yet they cannot generalise these capabilities to a broader range of tasks.
 Thus, the development of a single unified framework that concurrently exhibits both broad competency and high fidelity remains largely unresolved.
@@
 Generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations, whereas precision demands restrictive signals that minimise command-tracking error.
-Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
-  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [2017_DRL_survey, 2017_Schulman_PPO_instructure, 2023_Omur_comparison_RLalgorithm_bipedallocomotion].
+Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [105, 63].
+  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [3, 96, 5].
 This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only supports the success of the final task.
-Consequently, the immense amount of data required for an agent to explore, discover a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation merely to make training tractable [2021_Sim_Isaac, 2017_heess_emergence_animation_PPO_comprarisonrlalgorithm_differenttarrains, 2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
+Consequently, the immense amount of data required for an agent to explore, discover a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation merely to make training tractable [93, 39, 84, 83].
   • High-quality data scarcity: As highlighted in Section II-A2, the scarcity of high-quality demonstrations is a key bottleneck.
-Such data provide essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while avoiding unsafe exploration [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2020_motioncapture_referencebased_e2e_simulation_pybullet_augmentedrewarddesign].
-This scarcity stems from the difficulty of transferring scalable human data due to embodiment mismatch [2017_Ayusawa_motionretarget_optimization, 2019_iit_wholebody_retargeting, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy].
+Such data provide essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while avoiding unsafe exploration [83, 84, 118].
+This scarcity stems from the difficulty of transferring scalable human data due to embodiment mismatch [6, 82, 31], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [40, 60].
 These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly reflected in the design of the field’s dominant control architectures.
 End-to-end frameworks attempt a holistic solution, learning a single monolithic policy that must implicitly resolve all challenges simultaneously.
-While this approach can yield highly versatile and dynamic behaviours [2023_radosavovic_realworld_transformer_NNmodel], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within unstable system dynamics.
-This often results in a lack of the fidelity and precision that hierarchical systems can enforce [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
+While this approach can yield highly versatile and dynamic behaviours [88], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within unstable system dynamics.
+This often results in a lack of the fidelity and precision that hierarchical systems can enforce [63].
 Conversely, hierarchical frameworks are a direct architectural response to the lack of skill compositionality.
 By employing a “divide and conquer” strategy, they use an HL policy to sequence a library of LL, often model-based, controllers.
-This structure enforces precision and manages complex dynamics at a lower level [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
-However, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P].
+This structure enforces precision and manages complex dynamics at a lower level [60, 24, 112].
+However, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [101].
 IV-B Challenges in transferring from simulation to reality
@@
 While simulation offers a safe and efficient alternative, the ultimate goal of “zero-shot” transfer, where a policy works perfectly without any real-world fine-tuning, is rarely achieved.
-A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2021_DeepQ_DNQ_stepping_animationlonly_steppingstones, 2017_Josh_drl_animation_imitationlearning_GRAIL_TRPO_MUJOCO, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly].
+A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [84, 83, 74, 75, 105].
 Even when transfer is successful, it often comes with compromises.
-Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [2019_Yu_Phy_system_identification, 2020_Xie_firstsim2real_].
-In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [2021_siekmann_blind_DRL_stair, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2020_park_recoverstability_drl_PPO_referencebased_simulation].
+Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [124, 116].
+In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [99, 25, 81].
 This gap is caused by unavoidable differences between the virtual and physical worlds, which are especially problematic for dynamically unstable bipedal robots.
@@
   • Sensing and state estimation: A simulated robot has access to perfect, noise-free state information.
-In the real world, these states must be estimated from noisy sensors such as IMUs and joint encoders [2020_Xie_firstsim2real_, 2019_Yu_Phy_system_identification].
+In the real world, these states must be estimated from noisy sensors such as IMUs and joint encoders [116, 124].
 For a bipedal robot, precise state estimation is critical for maintaining balance.
-Simulators such as Isaac Gym [2021_Sim_Isaac], RoboCup3D [2003_Sim_Robocup], OpenAI Gym [2016_Sim_OpenAI], and MuJoCo [2012_Sim_Mujoco], detailed in Appendix B, are widely used to train policies that closely mimic real-world physical conditions.
-These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [2020_Xie_firstsim2real_, 2022_Ashish_RMA_sim2real_UCB, 2023_Rohan_sim2real_currentdecompose] have demonstrated efficient and high-performance results.
+Simulators such as Isaac Gym [93], RoboCup3D [10], OpenAI Gym [12], and MuJoCo [109], detailed in Appendix B, are widely used to train policies that closely mimic real-world physical conditions.
+These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [116, 55, 102] have demonstrated efficient and high-performance results.
 Despite these advancements, a significant gap persists between simulation and reality, exacerbated by the approximations made in simulation and the unpredictability of physical environments.
 IV-C Safety-critical locomotion
-Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [2024_tong_advancements_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot].
+Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [110, 91, 14].
 This includes ensuring the robot’s own integrity to prevent costly damage, as well as guaranteeing the safety of the surrounding environment and any humans within it.
@@
 This creates a critical barrier that separates success in controlled laboratory settings from reliable operation in the unpredictable real world.
-  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) such as joint angles and IMU data [2020_Siekmann_drl_e2e_s2r_cassie, 2021_siekmann_blind_DRL_stair, 2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], creating a major safety risk.
+  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) such as joint angles and IMU data [100, 99, 98], creating a major safety risk.
 Lacking external perception, these robots cannot anticipate obstacles, slopes, or slippery surfaces, making them purely reactive and highly prone to failure.
 Despite these significant safety drawbacks, this approach is often adopted for several reasons: omitting vision simplifies the control problem to pure motor skills and avoids the computational cost of real-time visual processing.
-Moreover, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [2023_duan_OSUDRL_heightmap_visionbased_hybrid] or path planning [2018_Tsunekawa_visualnavigation_hierarchy_DDPG-sim2real_NAO], rather than a core requirement for basic stability.
+Moreover, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [72] or path planning [69], rather than a core requirement for basic stability.
   • Lack of physical constraint satisfaction: Many DRL frameworks lack built-in mechanisms to guarantee physical constraint satisfaction.
-This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules—for example, Safe-RL on humanoids [2020_Javier_safereinforcementlearning], hybrid DRL with identified low-dimensional safety models [2022_Li_UCB_safteystability_model-based_DRL_PPO], footstep-constrained DRL policies [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation], and reactive DRL steppers operating under feasibility constraints on uneven terrain [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones].
+This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules—for example, Safe-RL on humanoids [34], hybrid DRL with identified low-dimensional safety models [64], footstep-constrained DRL policies [24], and reactive DRL steppers operating under feasibility constraints on uneven terrain [74].
 This limitation makes it difficult to prevent the robot from exceeding joint limits, applying excessive torques, or causing self-collisions, particularly when reacting to unexpected events.
@@
 V-A1 Multi-skill learning
-A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [2023_radosavovic_realworld_transformer_NNmodel, 2024_radosavovic_humanoid_locomotiontokenprediction_transformer, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
+A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [88, 89, 63].
 This research direction focuses on enabling robots to acquire, adapt, and deploy a broad and varied repertoire of motor skills, allowing them to handle unforeseen situations and operate effectively in unstructured environments.
@@
 The structured approach focuses on explicit decomposition.
-A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, an HL planner generates reference trajectories, while an LL DRL controller executes them robustly [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation, 2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller], as shown in Section III.
-Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
-A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single, compact generalist policy [huang2024diffuseloco].
+A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, an HL planner generates reference trajectories, while an LL DRL controller executes them robustly [31, 15], as shown in Section III.
+Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [83].
+A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single, compact generalist policy [46].
 V-A2 Traversing challenging environments
@@
 Validating capabilities on such terrains serves a crucial dual purpose.
-It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [2021_siekmann_blind_DRL_stair, 2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction, 2023_radosavovic_realworld_transformer_NNmodel].
-More critically, it benchmarks precision on treacherous paths such as stepping stones, which demand exact foot placement [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_junheng_steppingstones_MPCWBC, 2021_DeepQ_DNQ_stepping_animationlonly_steppingstones].
+It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [99, 111, 88].
+More critically, it benchmarks precision on treacherous paths such as stepping stones, which demand exact foot placement [25, 59, 74].
 These environments are the ultimate test of both a robot’s skill repertoire and its control fidelity.
 V-A3 Efficient learning and reward shaping
-As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [2017_DRL_survey, 2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie].
+As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [3, 98].
 Addressing this bottleneck is a crucial research frontier that calls for both more efficient algorithms and more robust reward designs.
@@
 A primary strategy is to leverage prior data rather than learning entirely from scratch.
-Leveraging prior knowledge provides strong guidance and reduces unsafe exploration by anchoring policies to feasible motion patterns [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_xie_OSUDRL_feedback_DRL_reference-based, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation].
-Curriculum learning further organises training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [2020_zhaoming_drl_steppingstones_PPOwithactorcritic_referencefree_simulation, 2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
+Leveraging prior knowledge provides strong guidance and reduces unsafe exploration by anchoring policies to feasible motion patterns [83, 115, 31].
+Curriculum learning further organises training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [117, 92, 112].
 Complementing advances in algorithms is the design of effective and robust rewards.
-Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [2017_heess_emergence_animation_PPO_comprarisonrlalgorithm_differenttarrains, 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
-Phase-aware objectives are well established for cyclic gaits such as walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], whereas reward design for non-periodic skills such as jumping is less standardised and often task specific [2023_UCB_Cassie_RL_Jumping].
-Promising directions reduce manual effort by adding higher-level guidance, including event-based terms, goal-conditioned objectives, and kinematic reference tracking [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2021_michael_velocity-basedcontrol_motioncapturedata, 2018_yang_push_recover_wholebody].
-Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand-crafted objectives with implicit ones inferred from demonstrations [2016_GAIL_imitation_learning].
+Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [39, 83].
+Phase-aware objectives are well established for cyclic gaits such as walking [98], whereas reward design for non-periodic skills such as jumping is less standardised and often task specific [62].
+Promising directions reduce manual effort by adding higher-level guidance, including event-based terms, goal-conditioned objectives, and kinematic reference tracking [24, 107, 119].
+Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand-crafted objectives with implicit ones inferred from demonstrations [41].
 Together, these directions seek to minimise skill-specific tuning and improve the transferability and reliability of learned locomotion policies.
@@
 This presents a significant opportunity: the potential to learn from vast libraries of human motion data.
-While large-scale datasets such as AMASS [2019_Naureen_AMASS] and Motion-X [2024_motion_X_datasets] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [2024_speratebody_expressive_allmotion].
+While large-scale datasets such as AMASS [71] and Motion-X [67] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [19].
 Therefore, motion retargeting emerges as a critical component to bridge this gap.
@@
 The first approach seeks to reduce the discrepancy by improving the simulation’s fidelity.
-This is often achieved through system identification (SI), where real-world robot data are used to fine-tune the simulator’s parameters to create a more accurate “digital twin” [2019_Yu_Phy_system_identification, 2023_masuda_simtoreal_identification].
-This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [2019_Yu_Phy_system_identification, 2019_simtoreal_actuatornet_quadrupedrobot].
-Other methods, such as designing specialised feedback controllers [2021_Casillo_cyberbotics_hierarchyscheme_bezierpolynominals], also contribute by making the system less sensitive to residual modelling errors.
+This is often achieved through system identification (SI), where real-world robot data are used to fine-tune the simulator’s parameters to create a more accurate “digital twin” [124, 73].
+This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [124, 48].
+Other methods, such as designing specialised feedback controllers [16], also contribute by making the system less sensitive to residual modelling errors.
 In contrast, the second philosophy accepts that simulations will always be imperfect and instead focuses on creating highly adaptive, robust policies.
 The primary method here is DR, which forces a policy to generalize by training it across a wide range of simulated physical variations.
-Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [2022_Ashish_RMA_sim2real_UCB], or via policy distillation, where a privileged “teacher” guides a “student” policy [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction] to have a knowledge of unknown information like friction .
-Additionally, techniques like adversarial motion priors [2024_Zhang_wholebody_adversarial_motion_priors, 2024_tang_humanmimic_adversarial] are used to ensure the learned behaviours are not just robust but also physically plausible.
+Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [55], or via policy distillation, where a privileged “teacher” guides a “student” policy [111] to have a knowledge of unknown information like friction .
+Additionally, techniques like adversarial motion priors [126, 104] are used to ensure the learned behaviours are not just robust but also physically plausible.
 Looking ahead, the ultimate goal remains achieving reliable zero-shot transfer, where no real-world fine-tuning is needed.
@@
 This shift from reactive to anticipatory control is essential for navigating unstructured real-world environments.
-The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_duan_OSUDRL_heightmap_visionbased_hybrid, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
+The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [25, 72, 112].
 In contrast, LiDAR is an active sensing method that generates precise 3D point clouds of the terrain.
@@
 Based on this sensory data, current research is exploring two primary pathways for processing perceptual information for control.
-The first involves creating an intermediate geometric representation, such as a height map from scanners [2023_duan_OSUDRL_heightmap_visionbased_hybrid].
+The first involves creating an intermediate geometric representation, such as a height map from scanners [72].
 This provides the policy with structured topographical data for effective footstep planning.
-The second is a more end-to-end approach, which utilises direct vision inputs such as RGB or depth images as inputs to the RL policy for real-time decision-making [2018_Tsunekawa_visualnavigation_hierarchy_DDPG-sim2real_NAO, 2022_byravan_nerf2real_basketball].
+The second is a more end-to-end approach, which utilises direct vision inputs such as RGB or depth images as inputs to the RL policy for real-time decision-making [69, 13].
 The former offers interpretability, while the latter promises more nuanced, reactive behaviours learned directly from raw perception.
@@
 While the previously discussed pathways focus on enhancing a robot’s capabilities, a parallel and equally critical frontier is ensuring that these capabilities are exercised safely and reliably.
-To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [2020_Javier_safereinforcementlearning, 2022_Li_UCB_safteystability_model-based_DRL_PPO].
+To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [34, 64].
 Soft constraints encourage desirable behaviour and penalise undesirable behaviour without forbidding it.
-They are well suited to preferences or efficiency goals, for example minimising energy use, limiting peak torques, or promoting smooth motion [2020_Javier_safereinforcementlearning].
+They are well suited to preferences or efficiency goals, for example minimising energy use, limiting peak torques, or promoting smooth motion [34].
 Hard constraints are inviolable rules that prevent catastrophic failures.
-They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2021_Casillo_cyberbotics_hierarchyscheme_bezierpolynominals].
-A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [2016_Quan_gaitlibrary].
+They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [24, 16].
+A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [79].
 In practice, a robust and trustworthy bipedal robot will likely combine both ideas.
@@
 They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for an LL policy to follow.
-This has been demonstrated effectively in legged robotics, where VLMs process raw sensory data to pass structured commands to motor controllers [2024_VLM_PC], creating a seamless link between strategic planning and physical action.
+This has been demonstrated effectively in legged robotics, where VLMs process raw sensory data to pass structured commands to motor controllers [18], creating a seamless link between strategic planning and physical action.
 Furthermore, FMs create a significant opportunity to overcome one of the most persistent bottlenecks in DRL: reward design.
 Instead of tedious manual tuning, LLMs can dynamically generate or refine reward functions based on linguistic descriptions of task success.
-Research has shown that LLMs can translate human feedback into reward adjustments [2023_kumar_wordscontrol_LLM] or even autonomously adjust rewards and control strategies to self-optimise for diverse terrains [2024_LLM_AnyBipe], drastically reducing human intervention.
+Research has shown that LLMs can translate human feedback into reward adjustments [56] or even autonomously adjust rewards and control strategies to self-optimise for diverse terrains [121], drastically reducing human intervention.
 The foremost opportunity lies in the deeper synergy between these roles.
 The integration of the HL symbolic reasoning of FMs with the LL, real-time control of DRL could create a new class of highly adaptive and flexible robots.
-As this rapidly evolving field progresses, as reviewed in [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
+As this rapidly evolving field progresses, as reviewed in [28], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
 V-B2 Loco-manipulation tasks
@@
 The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction.
-Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [2015_DAPRA_Challenge_1].
+Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [4].
 However, realising this opportunity is a significant challenge.
-Early studies, such as a ‘box transportation’ framework [2023_dao_OSUDRL_loco-manipulation_hybrid_simtoreal], often rely on inefficient, multi-policy solutions that lack visual perception.
-Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [2023_baltes_control_scooter, 2023_deepmind_haarnoja_playsoccer].
+Early studies, such as a ‘box transportation’ framework [22], often rely on inefficient, multi-policy solutions that lack visual perception.
+Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [7, 38].
 These difficulties create significant research opportunities.
 One such opportunity lies in exploring hierarchical control approaches.
-By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller].
+By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [15].
 Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution.
-Using techniques like curriculum learning and imitation from human motion-capture data [2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining, 2023_seo_deep_loco_manipulation, 2024_Zhang_wholebody_adversarial_motion_priors, 2024_speratebody_expressive_allmotion], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
+Using techniques like curriculum learning and imitation from human motion-capture data [92, 112, 63, 97, 126, 19], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
 V-B3 Insights from quadruped robots
 While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems.
-The diversity of frameworks developed for quadrupeds ranges from end-to-end, model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [quadModelBased1, quadModelBased2], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [quadDeformable].
-Furthermore, dynamic quadruped models facilitate highly adaptable policies [quadModular, humphreys2024bio], and sophisticated acrobatic motions are achieved through IL [quadImitation2].
+The diversity of frameworks developed for quadrupeds ranges from end-to-end, model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [103, 114], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [20].
+Furthermore, dynamic quadruped models facilitate highly adaptable policies [27, 47], and sophisticated acrobatic motions are achieved through IL [32].
 The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks that integrate vision-based systems.
-To date, two primary versions of such hierarchical frameworks have been developed: one where a deep-planning module is paired with model-based control [quadRLPlanning] within a deep-planning hybrid scheme, and another that combines model-based planning with LL DRL control [quadRLLL, quadmpcref] within a feedback DRL control hybrid scheme.
+To date, two primary versions of such hierarchical frameworks have been developed: one where a deep-planning module is paired with model-based control [33] within a deep-planning hybrid scheme, and another that combines model-based planning with LL DRL control [51, 52] within a feedback DRL control hybrid scheme.
 The latter has shown substantial efficacy; it employs an MPC to generate reference motions, which are then followed by an LL feedback DRL policy.
-Additionally, the Terrain-aware Motion Generation for Legged Robots module [TAMOLS] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
+Additionally, the Terrain-aware Motion Generation for Legged Robots module [50] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training.
 However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
 Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are under way.
-These include mimicking real animals through motion-capture data and IL [quadAnimal1, quadAnimal2], as well as augmenting quadrupeds with manipulation abilities.
-This is achieved either by adding a manipulator [quadManip1, quadManip2] or by using the robots’ legs [quadManip3].
-Notably, the research presented in [quadManip2] demonstrates that loco-manipulation tasks can be effectively managed using a single, unified, end-to-end framework.
+These include mimicking real animals through motion-capture data and IL [85, 122], as well as augmenting quadrupeds with manipulation abilities.
+This is achieved either by adding a manipulator [70, 30] or by using the robots’ legs [2].
+Notably, the research presented in [30] demonstrates that loco-manipulation tasks can be effectively managed using a single, unified, end-to-end framework.
 Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers, and high mechanical complexity can be attributed to this disparity between quadruped and bipedal robots.
@@
 V-C1 Bipedal foundation models
-Inspired by Robot Foundation Models (RFMs) [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture, 2023_generalpurposerobot_FMs], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion.
+Inspired by Robot Foundation Models (RFMs) [28, 44], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion.
 A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments.
 Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots.
-Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [2023_VLA_Rt-2].
+Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [128].
 The proposed BFM paradigm would operate in two stages.
@@
 Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot.
-The potential of this approach is highlighted by recent works, with frameworks like FLaRe [2024_FLaRe_RLfinetune_foundationmodel] enhancing generalisation for long-horizon tasks, MOTO [2024_MOTO_RLfinetune_foundationmodel] enabling effective offline-to-online adaptation from images, and AdA [2023_RL_finetune_RMFs] demonstrating in-context adaptation to novel environments.
+The potential of this approach is highlighted by recent works, with frameworks like FLaRe [43] enhancing generalisation for long-horizon tasks, MOTO [90] enabling effective offline-to-online adaptation from images, and AdA [8] demonstrating in-context adaptation to novel environments.
 Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
 However, realising the BFM concept for bipeds presents significant challenges.
-The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [2024_FLaRe_RLfinetune_foundationmodel, 2023_RL_finetune_RMFs].
+The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [43, 8].
 Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
@@
 Each layer processes context-specific inputs and outputs interpretable commands.
-The HL reasoning layer leverages large pre-trained models such as LLMs and VLMs [2022_DoAsICan, 2022_codeaspolicy_google] to parse commands into sub-tasks.
-For instance, Vision–Language Model Predictive Control [2024_VLM_PC] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
+The HL reasoning layer leverages large pre-trained models such as LLMs and VLMs [49, 65] to parse commands into sub-tasks.
+For instance, Vision–Language Model Predictive Control [18] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning.
 By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
-The mid-level planner selects or synthesises motions via learned motion libraries [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter] or DRL-based planners [2021_Mohammadreza_DRL_improveparameters_forplanner].
+The mid-level planner selects or synthesises motions via learned motion libraries [36, 61] or DRL-based planners [53].
 The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands.
-These include locomotion primitives like walking and climbing [2024_yangOu_LLM_hierarchical_quadrupedrobot], adaptive tracking controllers for whole-body tracking [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and imitation-based skills such as kicking and dancing [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task], by utilising RL, IL, or model-based methods.
-This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [2024_yangOu_LLM_hierarchical_quadrupedrobot].
+These include locomotion primitives like walking and climbing [80], adaptive tracking controllers for whole-body tracking [31], and imitation-based skills such as kicking and dancing [83], by utilising RL, IL, or model-based methods.
+This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [80].
 However, realising the MLAM concept would introduce challenges distinct from BFMs.
-Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [TAMOLS].
-Additionally, the computational latency incurred by HL reasoning modules like LLMs [2024_yangOu_LLM_hierarchical_quadrupedrobot] would pose limitations for tasks needing rapid reactions.
+Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [50].
+Additionally, the computational latency incurred by HL reasoning modules like LLMs [80] would pose limitations for tasks needing rapid reactions.
 VI Conclusion
@@
 DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks.
-End-to-end frameworks have demonstrated success in handling diverse locomotion skills [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], climbing stairs [2021_siekmann_blind_DRL_stair], and navigating challenging terrains such as stepping stones [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark].
+End-to-end frameworks have demonstrated success in handling diverse locomotion skills [63], climbing stairs [99], and navigating challenging terrains such as stepping stones [25].
 Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
@@
 Multi-Layer Perceptrons (MLPs), a fundamental NN structure, excel in straightforward regression tasks with lower computational resource requirements.
-A comprehensive comparison between MLPs and the memory-based NN LSTM reveals that MLPs have an advantage in convergence speed for many tasks [2023_Rohan_sim2real_currentdecompose].
-However, LSTMs, as variants of Recurrent Neural Networks (RNNs), are adept at processing time-associated data, effectively relating different states across time and modelling key physical properties vital for periodic gaits [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] and successful sim-to-real transfer in bipedal locomotion.
+A comprehensive comparison between MLPs and the memory-based NN LSTM reveals that MLPs have an advantage in convergence speed for many tasks [102].
+However, LSTMs, as variants of Recurrent Neural Networks (RNNs), are adept at processing time-associated data, effectively relating different states across time and modelling key physical properties vital for periodic gaits [98] and successful sim-to-real transfer in bipedal locomotion.
 Additionally, Convolutional Neural Networks (CNNs) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial.
@@
 Value-based methods, e.g.
-Q-learning, State–Action–Reward–State–Action (SARSA), and Deep Q-learning (DQN) [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones], excel only in discrete action spaces and often struggle with high-dimensional action spaces.
+Q-learning, State–Action–Reward–State–Action (SARSA), and Deep Q-learning (DQN) [74], excel only in discrete action spaces and often struggle with high-dimensional action spaces.
 Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy.
@@
 More advanced algorithms combine both policy-based and value-based methods.
-The Actor–Critic (AC) framework simultaneously learns both a policy (actor) and a value function (critic), combining the advantages of both approaches [2015_continous_control_deep_reinforcement_learning, 2016_guided_learning_of_Control_Graphs_physics-based_characters].
-Popular algorithms such as Trust Region Policy Optimisation (TRPO) [2015_TRPO_foundamentalpaper] and PPO, based on policy-based methods, borrow ideas from AC.
-Moreover, other novel algorithms based on the AC framework include Deep Deterministic Policy Gradient (DDPG) [2023_DDPG_e2e_openaigym_mujoco], Twin Delayed Deep Deterministic Policy Gradients (TD3) [2019_stephen_TD3_instruction], A2C (Advantage Actor–Critic), A3C (Asynchronous Advantage Actor–Critic) [2022_Jie_A3C_sim2real], and SAC (Soft Actor–Critic) [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
+The Actor–Critic (AC) framework simultaneously learns both a policy (actor) and a value function (critic), combining the advantages of both approaches [66, 68].
+Popular algorithms such as Trust Region Policy Optimisation (TRPO) [95] and PPO, based on policy-based methods, borrow ideas from AC.
+Moreover, other novel algorithms based on the AC framework include Deep Deterministic Policy Gradient (DDPG) [45], Twin Delayed Deep Deterministic Policy Gradients (TD3) [21], A2C (Advantage Actor–Critic), A3C (Asynchronous Advantage Actor–Critic) [57], and SAC (Soft Actor–Critic) [123].
 Each algorithm has its strengths for different tasks in bipedal locomotion scenarios.
 Several key factors determine their performance, such as sample efficiency, robustness and generalisation, and implementation complexity.
-A comparative analysis [2023_Omur_comparison_RLalgorithm_bipedallocomotion] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO, which attains relatively high scores.
-In [2017_Schulman_PPO_instructure], PPO demonstrates robustness and computational efficiency in complex scenarios such as bipedal locomotion, utilising fewer resources than TRPO.
-In terms of training time, PPO is much faster than SAC and DDPG [2023_Omur_comparison_RLalgorithm_bipedallocomotion].
-Moreover, many studies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2020_Siekmann_drl_e2e_s2r_cassie] have demonstrated its robustness and ease of implementation.
+A comparative analysis [5] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO, which attains relatively high scores.
+In [96], PPO demonstrates robustness and computational efficiency in complex scenarios such as bipedal locomotion, utilising fewer resources than TRPO.
+In terms of training time, PPO is much faster than SAC and DDPG [5].
+Moreover, many studies [98, 36, 100] have demonstrated its robustness and ease of implementation.
 Combined with its flexibility to integrate with various NN architectures, this has made PPO the most popular choice in the field.
-Numerous studies have shown that PPO can enable the exploration of walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], jumping [2023_UCB_Cassie_RL_Jumping], stair climbing [2021_siekmann_blind_DRL_stair], and stepping-stone traversal [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark], demonstrating its efficiency, robustness, and generalisation.
+Numerous studies have shown that PPO can enable the exploration of walking [98], jumping [62], stair climbing [99], and stepping-stone traversal [25], demonstrating its efficiency, robustness, and generalisation.
 Additionally, the DDPG algorithm integrates the Actor–Critic framework with DQN to facilitate off-policy training, further optimising sample efficiency.
-In certain scenarios, such as jumping, DDPG shows higher rewards and better learning performance than PPO [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2022_Chongben_pallel_DDPG_e23_silumation_only].
-TD3, developed from DDPG, improves upon the performance of both DDPG and SAC [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
+In certain scenarios, such as jumping, DDPG shows higher rewards and better learning performance than PPO [105, 106].
+TD3, developed from DDPG, improves upon the performance of both DDPG and SAC [123].
 SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3.
@@
 This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential.
-SAC is also known for its stability and strong performance across a wide range of tasks [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
+SAC is also known for its stability and strong performance across a wide range of tasks [123].
 While A2C offers improved efficiency and stability compared with A3C, the asynchronous update mechanism of A3C provides better exploration capability and accelerates learning.
@@
 Creating a reliable simulation environment and conducting RL training is challenging.
-The literature shows that several simulators are available, including Isaac Gym [2021_Sim_Isaac], RoboCup3D [2003_Sim_Robocup], OpenAI Gym [2016_Sim_OpenAI], MuJoCo [2012_Sim_Mujoco], Orbit [2023_Orbit_simulator], Brax [2021_Brax], and Isaac Lab [isaaclab2023].
+The literature shows that several simulators are available, including Isaac Gym [93], RoboCup3D [10], OpenAI Gym [12], MuJoCo [109], Orbit [77], Brax [29], and Isaac Lab [76].
 OpenAI developed Gym and Gymnasium to provide lightweight environments for rapid testing of RL algorithms, including simplified bipedal locomotion models.